A Complete Bibliography of Publications in the Journal of Statistical Theory and Practice

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

26 March 2025 Version 1.09

Title word cross-reference

(1, 1) [469]. (m, ni) [769]. (p^{2}, p, p^{2}, p) [257]. (q − 1) [32]. \{0, ±\frac{1}{2}, ±1, 2\} [841]. 16 [161]. 2 [111, 160]. 2^{m_{1}+m_{2}} [31]. 2^{n} [669]. 2^{n−(n−k)} [916]. 2^{n} × 2^{n} [810]. 2 ≤ m_{1}, m_{2} ≤ 4 [31]. 3 [161]. 3^{n−m} [583]. 5 [256]. $83.95 [387]. < [449, 442]. ^{b} [12]. ^{m} [265]. ^{n} [437, 810]. _{2} [115]. A [735, 673, 207, 389]. AB: BA [987]. α [21, 645, 123]. α = 6 [161]. BIBD(v, 2, 1) [361]. C_{2} [12]. χ^{2} [959]. χ^{2}(D) [369]. D [558, 142, 28, 76, 670, 598, 693, 828, 927]. E [262, 226, 209, 5]. F [903, 41]. G [726]. I [807]. K [255, 1024, 851, 876, 140, 242, 369]. L [830]. L_{q} [996]. λ = 1 [161]. M [244, 255, 198, 643, 453, 84, 260, 357]. C [617]. P(X < Y) [659]. P(X < Y < Z) [1008]. R^{d} [963]. n [260]. N ≢ 0 (mod s) [263]. p [871, 723]. ψ [458]. Q [547, 32, 883].

R [199]. R({10, 01} ∪ Ω ∗ |Ω) [31]. s^{2} [226]. s^{m} [263]. t [315, 612, 452, 254, 53, 598]. T^{2} [602]. t_{0} [865]. × [260, 29]. U [443, 971, 469, 822]. v [361]. X [449, 5]. Y [449, 5]. Z_{p}^{3} [257].

-Arcs [255]. -Bose [883]. -Bumps [851]. -Circulant [369, 242]. -Designs [256, 123, 830]. -Difference [257]. -efficient [76]. -estimation-based [198]. -estimator [244]. -Estimators [643]. -Fermi [883]. -Intersecting [255]. -level [160]. -Likelihood [996]. -optimal [226, 735, 142, 673, 369, 670, 828, 558, 28, 207, 598, 693]. -Optimality [262, 807, 209]. -Qualitatively [453]. -Record [1024]. -regression [458]. -Regular [617]. -Repeated [111, 84]. -statistic [469]. -statistics [971]. -symmetric [199]. -test [903]. -Tests [443]. -type [602, 822]. -values [723]. -Way [357].

-Wise [254].

/CRC [20]. /G/1 [12].

1 [12, 723]. 19 [856].

2 [277]. 2-Level [931]. 2007 [123]. 2D [983].

3D [983].

978-1-4398-0910-5 [387].

= [5, 449, 442].

A-ComVar [690]. A-efficient [532]. A/B [670]. Abelian [258]. Aberration [734, 533]. Absolute [711, 730, 471]. Absorbing [94]. Abundance [944]. Accelerated [623, 633, 596, 595, 586]. Accept [345]. Acceptance [149, 280]. Account [1021, 1019, 320]. Accounting [885, 891]. Accuracy [299, 995, 953, 918, 802, 535]. Achieving [911]. Action [169]. Active [807]. Activity [787]. Adaptive [534, 222, 519, 872, 1010, 422, 682, 363, 95, 929, 611, 513, 555]. Addendum [671]. Addition [769]. Additive [361]. Additivity [27]. Addressing [746]. Adjacent [213]. Adjusted [525, 16, 119, 513]. Adjustment [215, 694, 479]. Advances [925, 1022, 181, 359, 372]. Advancing [336]. Adversarial [792]. affect [856]. Affine [259, 21, 780]. After [434, 276, 822, 742]. Against [130, 131, 536]. Aggregate [215]. Aggregation [708]. Aided [162]. algebra [509]. Algebraic [261, 617]. Algorithm [878, 300, 851, 328, 40, 937, 497, 369, 882, 655, 529]. Algorithm-Based [300, 497]. Algorithmic [226, 981]. Algorithms [619, 739, 208]. Aliased [931]. Aliasing [1005]. Allocation [359, 372, 611, 513, 555]. Allocations [153]. Allowance [714]. Allowing [980]. Almost [849, 96, 1013, 587]. Almost-Pareto [96]. Aloke [741]. alpha [487]. alpha-generalized [487]. Alternative [441, 987, 50, 898, 907, 903, 504]. Alternatives [796, 798, 131]. Amicable [350]. Among [1007]. Analogue [199]. Analyses [148, 135]. Analysis [289, 261, 861, 705, 332, 301, 1020, 90, 91, 25, 756, 450, 328, 50, 902, 694, 16, 72, 843, 39, 386, 933, 1022, 914, 900, 387, 1010, 816, 266, 553, 718, 41, 198, 238, 691, 1015, 625, 633, 924, 677, 311, 800, 959, 958, 943, 481, 604, 499, 485, 600, 489, 521]. Analytic [2]. Analytical [379, 840, 71]. Analytical-Numeric [379]. Analyze [275]. Analyzing [754, 243]. Ancillarity [780]. ANCOVA [148]. and/or [665]. Anderson [988]. Annual [395]. Anonymity [368]. ANOVA [731, 524, 469]. ANOVA/random [469]. ANOVA/random-effects [469]. Anscombe [126]. answers [534]. Any [361, 182]. apparent [565]. Applicable [25]. Application [666, 1001, 887, 662, 449, 99, 622, 979, 328, 700, 83, 888, 783, 886, 183, 399, 804, 59, 890, 710, 41, 698, 813, 167, 1014, 633, 103, 114, 986, 675, 996, 121, 504, 460, 489, 521]. Applications [847, 991, 393, 983, 251, 371, 1016, 39, 239, 761, 343, 926, 255, 172, 398, 729, 812, 897, 954, 707, 943, 481, 462, 514, 487, 510, 574, 475, 54]. Applied [909]. Applying [382]. Approach [12, 116, 833, 825, 711, 987, 248, 97, 678, 674, 280, 783, 750, 62, 75, 703, 200, 17, 809, 158, 755, 36, 814, 177, 398, 728, 405, 790, 404, 73, 981, 306, 594, 1002, 548, 575, 496]. Approach-Based [405, 575]. Approaches [202, 106, 50, 472]. Approximate [654, 125, 801]. Approximately [438]. Approximation [841, 429, 857, 56, 213]. Approximations [134, 998, 626, 530]. Arbitrary [852]. ARCH [11, 151]. Arcs [255]. area [592, 489]. area-level [592].

Areas [394]. Arm [905]. Armed [889]. Array [828, 598]. Arrays [360, 827, 22, 253, 349, 734, 260, 681, 733, 252]. Arrival [157, 516]. artifacts [531]. asbestos [601]. Aspects [819, 84]. Assay [323]. Assays [830]. Assess [711]. Assessing [644, 498, 564]. Assessment [925, 387, 92, 988, 440, 565, 547, 501, 567]. Assisted [667]. Associate [870, 306]. Associated [66, 924, 548, 531, 490]. Association [952, 225, 565, 456]. assumptions [547, 541]. Asymmetric [696, 714, 1027, 468]. Asymmetrical [43]. asymmetry [585]. Asymptotic [743, 67, 374, 313, 591, 896, 131, 646, 114, 602]. Asymptotics [961]. Attention [983]. Attraction [750]. Attribute [169, 907, 278, 579, 587, 478]. Attribute-Level [278]. Attributes [660, 528, 532]. Attribution [425]. Augmented [766, 863]. Auto [171, 408]. Auto-Induced [408]. Auto-Linear [171]. Autocorrelated [914, 495]. Autologistic [325]. Automorphism [258, 115]. Autoregressive [393, 303, 42, 918, 584, 165]. Auxiliary [789, 409, 709, 632, 143, 38, 156, 271, 688, 782, 899, 907, 179, 405, 575, 579, 587]. average [535]. Averages [11]. Averaging [236, 324]. Aylmer [34, 46].

B [670]. Bagging [832]. Bagging-Enhanced [832]. Balakrishnan [589]. Balance [28]. Balanced [193, 74, 251, 27, 348, 913, 178, 900, 31, 640, 725, 213, 981, 265, 264, 517, 456]. balking [516]. Ball [963]. Bandit [889]. Bands [272]. Bandwidth [645, 868, 217, 306, 548, 472]. Base [167]. Based [152, 430, 647, 1024, 1009, 965, 861, 106, 300, 283, 650, 627, 726, 66, 674, 623, 888, 102, 229, 818, 134, 109, 1025, 971, 631, 952, 890, 995, 254, 223, 984, 896, 159, 713, 920, 845, 174, 716, 639, 642, 794, 774, 405, 362, 976, 225, 643, 655, 828, 989, 61, 919, 854, 964, 867, 421, 1018, 471, 543, 575, 497, 478, 499, 176, 198, 577, 456, 535]. Bases [22, 1013]. Basic [215, 400, 333, 536]. Basis [121]. Batch [727, 157]. Batra [123]. Baye [93]. Bayes [796, 318, 219, 134, 79, 154, 71, 658]. Bayesian [387, 116, 614, 90, 91, 245, 477, 322, 248, 719, 166, 548, 674, 902, 401, 866, 16, 72, 323, 877, 973, 809, 673, 914, 814, 1026, 473, 554, 76, 292, 119, 325, 953, 92, 177, 176, 191, 785, 910, 238, 577, 781, 790, 395, 104, 324, 1027, 306, 472]. Bechhofer [728]. Before [276, 1006, 822]. Behavior [996, 547]. Behnken [467]. Behrens [341, 185, 835]. Beichelt [20]. Bell [853, 1014]. Beloved [741]. Belt [791]. Bera [385]. Beran [629]. Beran-Inspired [629]. Bernoulli [415, 840, 164, 516]. Best [905, 178, 278]. Beta [419, 175, 877, 804, 626, 859, 625, 943, 612]. Beta-Binomial [877]. Better [728]. Between [370, 723, 241, 952, 445, 626, 352, 829, 405, 831, 558, 483]. Beyond [547]. Bias [332, 858, 1003, 786, 694, 616, 267, 966, 159, 959, 272, 576, 464]. Bias-Corrected [959]. Bias-Reduced [272]. Biased [152, 367, 930, 772, 620, 692, 989, 515]. Bibliography [190]. Big [823, 976, 821, 1018]. Bimodality [866]. Binary [298, 876, 364, 914, 660, 885, 891, 956, 693, 828, 335, 121]. BINMA [518]. Binomial [126, 429, 876, 877, 804, 507, 472]. Bioinformatics [48]. Biological [491]. Biometrical [145]. Birnbaum [549, 851, 417, 990, 510, 302]. Birth [15, 94, 3, 228, 815]. Birth-death [94]. Biswas [123]. Bivariate [289, 861, 892, 630, 619, 791, 888, 589, 342, 487, 942, 890, 710, 656, 751, 864, 737, 112, 188, 475, 972, 702, 635, 546, 494, 543, 504, 599]. Bladder [884]. Block [74, 668, 833, 825, 906, 122, 913, 207, 900, 29, 620, 994, 161, 725, 214, 558, 981]. Block-Sum [913, 906]. Blocked [436, 442, 938, 365, 572]. Blocks [32, 687, 203, 983, 358, 26, 868]. Boca [387]. Bonds [704]. Bone [961]. Book [9, 294, 371, 387, 20]. Bootstrap [700, 403, 40, 125, 942, 968, 837, 821]. Bose [883]. Both [441, 81]. Boundaries [174]. Boundary [966]. Bounded [89, 859, 13]. Bounds [1011, 840, 777, 792, 776, 93, 800, 381]. Box [467, 147, 642, 1006]. Box-Hunter [147]. Boxes [92]. Bradley [309]. Brain [983]. Breakdowns [157]. Breast [301]. Brief [45]. bulk [516]. Bumps [851]. Burr [674, 636]. Burst [111, 84]. Busy [12].

C [815]. C. [25, 10]. Cacoullos [921, 881]. Calculation [715, 499]. Calibration [150, 575, 601, 496, 984, 405, 362]. Calibration-Based [362]. Cambanis [1024]. Cancer [301, 334, 884, 110]. Canonical [959]. Cantelli [1011]. Capability [977]. Capacity [672]. Capture [219]. Cardiac [1026]. Care [401, 336]. Carl [45]. Carlo [467, 60]. CARMA [922]. Case [204, 431, 786, 750, 738, 855, 897, 61, 585]. Case-Control [61]. Cases [414]. Catalog [442, 931]. Catastrophes [50, 3, 51]. Categorical [245, 803, 220, 952, 81, 44]. Categories [1017, 437, 810]. Category [180, 243]. Cauchy [881, 812]. Causal [790, 821]. Cause [307]. Causes [1015, 919]. Cautionary [52]. cDNA [109]. Celebrating [815]. Cell [148, 980]. Censored [412, 328, 448, 997, 674, 930, 982, 636, 860, 406, 941, 648, 713, 330, 217, 929, 304, 989, 603]. Censoring [663, 726, 428, 630, 623, 912, 661, 79, 154, 732, 910, 994, 73, 611, 609, 574, 485, 577, 522]. Censorship [917]. Census [936]. Centenary [815]. Centennial [19]. Central [406, 844, 730]. centrally [467]. Cepheid [141]. Certain [194, 28]. chain [538, 467, 575]. Chains [83, 50, 783, 220]. Chamber [1026]. Change [872, 902, 72, 773, 936, 190, 742, 198, 716, 917, 774, 989]. Change-Point [190, 742, 716, 774, 72]. Change-Points [872]. Changed [742]. Changepoint [685, 331]. Changes [783, 774]. Changing [925, 686]. Channel [425]. Chapman [20]. Characteristic [881, 774]. Characteristics [301]. Characterization [220, 631, 617, 794, 935, 265]. Characterizations [865, 589]. Charles [57]. Charlier [689]. Charlier-Like [689]. Chart [837, 676]. Charts [70, 650, 477, 520]. Chatterjee [1003]. Check [669]. Chen [888, 661, 835]. Chi [550]. Chi-squared [550]. Chirp [778]. Chirp-Like [778]. Chisquared [776]. Choice [204, 431, 384, 52, 18, 422, 103, 278, 568, 528, 533, 531, 502, 532, 529]. Cholesterol [286]. Choosing [905]. Christopher [57]. Circle [779]. Circulant [369, 242]. Circular [773, 391, 826, 640, 238, 555, 473, 554]. circularity [613]. Claim [295]. Class [668, 969, 865, 834, 241, 140, 758, 732, 76, 271, 782, 907, 235, 549, 513, 806, 579, 587, 600, 605]. Classes [146, 618, 169, 855]. classical [1002]. Classification [901, 941, 977, 95, 381]. classifications [168, 225]. Clinical [318, 319, 320, 359, 372, 684, 317]. close [588]. Closed [219]. Clouds [102]. Cluster [450, 471, 201, 4, 591]. Cluster-based [471]. Clustered [957]. Clustering [248, 779, 995, 799, 802]. Cochran [597]. Codes [256, 111, 84, 664, 254]. Coefficient [652, 274, 129, 351, 340, 629, 165, 536, 544]. Coefficients [951, 413, 246, 723, 241, 665, 715]. Cohen [927]. Coherence [399]. Cointegration [174]. Collisions [164]. Column [206, 830, 939, 923]. Columns [734, 1007]. combinations [1]. Combinatorial [84, 250]. Combined [428, 18, 667, 984, 304]. Combining [494, 591, 224, 692]. Comment [318, 319, 320, 780]. Common [690, 941, 742, 562]. Commutative [122]. Compactly [659]. Comparative [219, 1010, 521]. Comparing [987, 204, 431, 767, 285, 130, 961, 829, 407]. Comparison [202, 378, 647, 106, 332, 6, 998, 660, 363, 224, 895, 284, 472, 528]. Comparisons [805, 735, 207, 610]. Competent [746]. Competing [762, 407, 919]. Complementary [803, 859]. Complete [668]. Completely [117]. Completion [790]. Complex [382, 414]. Complexity [152, 894, 415]. Component [144, 958]. Components [32, 403, 386, 144, 610]. Composite [319, 210, 844, 863]. Compound [83, 997, 886, 811, 890, 938]. Comprehensive [1022]. Computation [878, 415]. Computational [261, 840]. compute [469]. Computer [205, 162, 59, 354, 211, 928, 363, 100]. Computing [55, 291]. ComVar [690]. Conceptual [499]. Concomitants [1024, 188]. Concordant [225]. condition [492]. Conditional [757, 719, 842, 930, 653, 982, 857, 406, 664, 376, 125, 223, 672, 169, 326, 908, 934]. Conditions [750, 325]. Confidence [905, 310, 726, 341, 241, 701, 984, 716, 918, 800, 184, 272, 394, 601, 562, 500, 535]. Configurations [250, 467]. Conflict [104]. Confounding [436, 365, 583]. Conglomerability [1000, 552]. Connected [830]. Connecting [333, 822, 980]. Connectivity [357]. Conquer [655]. Consensus [507]. Considerations [450, 679]. consistency [543]. Constant [906, 129, 351, 913, 233]. Constrained [61]. Constraint [712]. Constraints [390, 90, 91, 452]. Constructed [431, 733]. Constructing [202, 266, 355, 529]. Construction [74, 360, 734, 162, 436, 259, 226, 1004, 681, 583, 242, 369, 972, 981]. Constructions [193, 253, 254, 456, 252, 725]. consulting [476]. Consumption [649].

Contaminated [247, 80]. Contingency [805, 953, 999, 897, 80, 463]. Continual [335]. Continuous [627, 987, 951, 897, 212, 506, 538, 513, 498]. Continuous-Time [627, 538]. Contour [928]. Contoured [975, 459, 608]. Contours [683]. Control [70, 650, 735, 1021, 1019, 644, 187, 207, 837, 266, 209, 676, 749, 61, 495, 477, 520, 570, 460]. Controllable [397]. Controlled [404, 539]. Controlling [889]. Controls [767]. Convergence [83]. Conversation [69, 137]. Convex [93, 484]. Convoluted [689]. Convolution [222, 199, 509]. Convolutions [402, 514]. Conway [751, 986, 960]. coordinate [529]. coordinate-exchange [529]. Copula [979, 546, 791, 888, 750, 912, 494, 754, 504, 455]. Copulae [231]. Copulas [656, 972, 466]. Correct [416]. Corrected [959]. Correcting [334]. Correction [833, 332, 934, 1021, 267, 939, 813, 891]. Corrective [169]. Correctly [647]. Correlated [839, 665, 394, 957, 588]. Correlation [652, 858, 1003, 1007, 806, 723, 241, 751, 597, 536, 544]. Correspond [1007]. Cost [680, 861, 283, 328, 916, 293, 437, 810]. Cost-Effectiveness [861, 283]. Cost-Efficient [680]. Count [703, 275, 1015, 548]. Counts [936, 980]. Coupon [704]. Covariance [811, 955, 918, 938, 556, 903, 613]. Covariate [870, 324, 513, 564]. covariate-adjusted [513]. Covariates [332, 322, 767, 8]. Coverage [569, 142]. Covering [680, 253, 349, 681]. Coverings [818]. COVID [856]. COVID-19 [856]. COVID19 [901]. Cox [332, 138, 1006, 564, 331]. Coxian [12]. Crack [887, 718]. Cracks [417]. CRC [387, 20]. credit [483]. Criteria [274]. Criterion [141, 389, 772, 198, 716, 989]. crop [489]. Cross [27, 995, 225, 488]. Cross-classifications [225]. Cross-Validation [995, 488]. Crossover [987, 697, 826, 264, 456]. Crude [678]. Cum [688]. Cum-Ratio [688]. Cumulative [1017, 413, 966, 737, 999, 407]. Cure [300, 1020, 329, 804, 762, 882, 920, 604]. Currency [231]. Current [746, 407]. Curves [829, 394]. Customer [425]. Cutpoint [330]. Cutpoints [95]. Cutpoints- [95]. CvM [131]. Cycles [366]. Cyclic [827, 120, 254, 355, 161].

Darroch [32]. Data [289, 443, 847, 965, 705, 887, 180, 412, 332, 650, 310, 441, 768, 649, 328, 428, 736, 97, 448, 754, 997, 791, 740, 803, 401, 703, 653, 16, 200, 370, 773, 843, 219, 982, 323, 877, 636, 201, 948, 132, 869, 64, 860, 645, 329, 18, 900, 942, 1010, 314, 438, 859, 890, 710, 119, 961, 176, 845, 247, 238, 729, 77, 955, 275, 1015, 71, 823, 620, 330, 44, 628, 874, 855, 407, 311, 217, 976, 790, 304, 846, 989, 326, 949, 946, 61, 185, 821, 835, 957, 964, 603, 1018, 495, 494, 528, 503, 547, 602]. data [451, 590, 571, 563, 510, 500, 104]. Data-Driven [132]. Dataset [804]. Days [69]. DBEEW [656]. Dealing [688]. Death [15, 3, 228, 94]. Decennial [936]. Deciding [370]. Decision [106, 796, 97, 96, 387, 551]. Decomposition [1017, 80, 903]. Decompositions [463]. Deconvolution [65, 738, 659]. Deep [788, 841, 983, 792, 1022, 755]. default [483]. defective [604]. Deficit [167]. Definitive [1007, 355]. Degradation [622]. Delay [755]. Delhi [123]. Delta [622]. Demimartingales [873]. Demisubmartingales [873]. Denominations [231]. Densities [402, 89, 659]. Density [12, 269, 763, 66, 379, 60, 912, 645, 814, 875, 689, 738, 620, 217]. Dependence [750, 714, 829, 838, 492, 483]. Dependency [231, 959]. Dependent [448, 630, 951, 241, 971, 295, 557]. Depth [650, 962, 845]. Depth-Based [845]. Derivable [916, 810]. derivations [903]. Derivatives [66, 222, 232, 89]. Deriving [269]. Describing [370]. Design [261, 987, 218, 450, 997, 102, 818, 948, 1025, 267, 673, 889, 824, 356, 23, 377, 389, 24, 670, 698, 813, 35, 112, 677, 628, 115, 114, 611, 555, 558, 531, 482, 522]. design-induced [531]. Designed [679, 821]. Designing [277, 205, 278]. Designs [202, 32, 193, 432, 74, 680, 108, 256, 360, 668, 819, 203, 833, 825, 870, 906, 147, 663, 204, 431, 384, 206, 263, 735, 830, 262, 367, 690, 351, 433, 358, 767, 1007, 160, 366, 120, 357, 916, 43, 162, 435, 293, 437, 142, 834, 436, 697, 259, 21, 7, 28, 913, 148, 186, 226, 1004, 207, 30, 1005, 31, 242, 29, 266, 76, 352, 354, 208, 369, 939, 923, 826, 210, 355, 442, 931, 660, 844, 211, 640, 363, 639, 745, 993, 829, 733, 669, 359, 372, 684, 161, 110, 693, 822, 828, 365, 212, 863, 265, 214, 831]. Designs [264, 467, 534, 568, 488, 528, 533, 457, 583, 992, 517, 531, 456, 493, 598, 532, 123]. designs-part [517]. Detailed [936]. Detecting [111, 970, 84, 773]. Detection [685, 869, 4, 755, 334, 742, 917, 774, 989, 578, 459]. Determinants [678, 641]. Determination [731]. Determining [403]. deterministic [486]. Development [726, 514, 451]. Developments [892, 68, 580]. deviation [471]. Dey [741]. Diabetic [888]. Diagnose [944]. Diagnosed [647]. Diagnostic [647, 298, 299, 438, 625]. Diagnostics [373, 635]. Diagonal [382, 391, 972]. Dichotomous [176, 482, 498]. Diel [950]. Difference [251, 257, 241, 254, 718, 760, 643, 394, 606]. Difference-Type [760]. Differences [711, 846]. Different [428, 166, 246, 701, 640, 639, 73, 480]. Differential [195, 748]. Differentially [936]. Diffusion [951, 557]. Diffusion-Limited [951]. Diffusions [408]. Dimension [590, 471]. Dimensional [443, 652, 288, 385, 938, 949, 835, 556, 542, 501]. Dimensions [756, 63]. Dirac [883]. direct [498].

Directional [724]. Dirichlet [449, 995, 101]. disclosure [570]. Discontinuation [364]. Discordant [225]. Discovery [1021, 1019]. Discrepancy [818]. Discrete [842, 934, 140, 814, 656, 881, 864, 808, 897, 985, 545, 546, 548, 465, 533, 599, 553, 531, 468, 532, 529]. discrete-time [465]. Discretely [875]. Discretization [897]. Discriminant [896, 489]. Discriminate [853, 416]. Discrimination [352, 114]. disease [565]. Disjoint [256]. Dismemberment [889]. Dispersion [47, 353, 14, 635]. Distance [775, 896, 875, 161]. Distance-Based [896]. Distances [843, 869, 952, 995]. distancing [856]. Distributed [762, 421]. Distribution [878, 666, 378, 705, 887, 894, 415, 892, 662, 6, 126, 429, 757, 146, 979, 720, 411, 909, 876, 276, 249, 166, 426, 619, 997, 674, 623, 888, 930, 866, 967, 382, 380, 127, 281, 338, 877, 446, 64, 661, 793, 37, 343, 439, 417, 926, 765, 990, 904, 648, 128, 438, 995, 966, 797, 884, 223, 656, 896, 177, 713, 881, 418, 910, 917, 729, 730, 167, 188, 918, 302, 71, 994, 897, 944, 646, 702, 838, 707, 1008, 73, 170, 651, 989, 182, 326, 935, 717, 1027, 996, 867, 943, 421, 614, 612, 481, 1, 480]. distribution [589, 602, 497, 487, 478, 604, 571, 474, 515, 468, 550, 606, 125]. Distribution-Free [979, 438]. Distributional [774, 547]. Distributions [297, 12, 752, 991, 851, 883, 726, 879, 194, 383, 969, 719, 175, 865, 842, 934, 379, 840, 777, 401, 370, 886, 39, 296, 696, 342, 975, 130, 140, 631, 701, 79, 732, 414, 626, 1023, 199, 656, 751, 864, 695, 716, 238, 808, 715, 801, 416, 929, 702, 985, 339, 935, 986, 523, 1, 543, 504, 514, 599, 452, 609, 484]. Divergence [770]. Diversified [231, 53]. Divide [655]. Divide-and-Conquer [655]. Divisible [993]. DNA [267]. do [856]. Does [245]. Domain [750, 200, 945, 974]. Domains [89]. dose [534, 521]. dose-response [534, 521]. Dosing [312]. Double [277, 408, 709, 696, 691, 268, 582, 614, 508]. double-ranked [508]. Double-Threshold [408]. Doubly [175, 771, 475]. Drafts [786]. drift [486]. Driven [925, 132, 922, 247, 748, 522]. Dropout [764, 264]. Dual [632, 782, 899]. Due [417]. Duration [412, 119]. during [856]. Dynamic [274, 231, 333, 36, 473, 554]. Dynamics [227].

E-Bayesian [1027]. Early [69, 334]. Easy [712]. Easy-to-Implement [712]. Ecological [189]. Economic [954]. Editor [86]. Effect [227, 367, 322, 973, 723, 1005, 992, 698, 813, 672, 846, 562]. Effective [316]. Effectiveness [861, 283]. Effects [193, 397, 987, 353, 138, 1007, 806, 151, 636, 207, 334, 640, 745, 801, 490, 850, 699, 693, 828, 946, 908, 867, 558, 568, 539, 528, 499, 572, 469, 498, 564, 567]. Efficiency [62, 885, 891, 49, 529, 593]. Efficient [432, 680, 767, 162, 452, 824, 242, 327, 42, 271, 874, 855, 907, 264, 76, 595, 560, 532]. Egypt [783]. Einstein [883]. Electroanatomical [1026]. Element [733]. Elicitation [97]. Elimination [218]. Elliptically [975]. Embedding [108, 573]. Empirical [578, 79, 154, 128, 53, 868, 728, 500, 658, 774, 643, 394, 525]. Empirical-Characteristic-Function-Based [774]. Encryption [956]. End [1006]. Endpoints [998]. Engineering [729, 20]. Engineers [371]. Enhanced [192, 832, 451]. enrollment [522]. Ensembles [807]. Entropy [415, 768, 930, 737, 976, 996, 506, 516]. Entropy-Based [976]. Envelope [695]. Environment [25, 615, 557]. Environmental [189]. Enzyme [228]. Epidemic [47, 1002]. Epidemiological [324]. Equality [290, 413, 503, 609, 613]. equally [588]. Equating [9]. Equation [398, 5, 812]. Equations [195, 748]. Equivalence [141, 7, 445, 1006, 988]. Equivalent [218]. Equivariance [834]. ergodic [492]. Ergodicity [637]. Erratum [526, 410, 372]. Error [111, 970, 84, 409, 951, 323, 775, 893, 978, 634, 642, 880, 862, 659, 717, 559, 903, 505]. Errors [849, 940, 839, 723, 771, 191, 156, 744, 488, 459, 608, 581]. Establishing [445]. ESTAR [344, 398]. Estate [345]. Estimate [382, 376, 977, 862, 492, 559, 591, 469]. Estimated [665, 494]. Estimates [1009, 763, 469]. Estimating [1002, 545, 99, 63, 982, 616, 64, 59, 710, 953, 746, 772, 363, 339, 284, 683, 568, 528, 478, 496, 482]. Estimation [915, 666, 752, 434, 894, 300, 627, 390, 11, 789, 423, 122, 66, 768, 279, 872, 411, 83, 420, 526, 166, 426, 806, 930, 340, 747, 912, 40, 871, 409, 17, 281, 877, 696, 616, 201, 444, 270, 787, 721, 820, 447, 975, 766, 857, 793, 645, 241, 631, 778, 664, 798, 154, 18, 23, 926, 732, 765, 893, 648, 327, 922, 966, 667, 797, 884, 961, 196, 875, 159, 422, 742, 42, 191, 746, 978, 928, 748, 682, 672, 634, 143, 639, 188, 625, 181, 489, 620, 330, 14, 156, 240, 440, 628, 760, 880, 179]. Estimation [217, 427, 692, 646, 1008, 184, 73, 170, 165, 651, 326, 802, 945, 946, 1027, 61, 744, 675, 306, 996, 421, 594, 538, 592, 458, 480, 548, 596, 488, 465, 576, 464, 505, 571, 198, 515, 521, 566, 541, 585, 472]. Estimations [674, 577]. Estimator [671, 970, 757, 222, 402, 173, 709, 406, 932, 629, 875, 722, 305, 71, 405, 362, 339, 244, 581, 464, 582]. Estimators [1001, 378, 849, 940, 150, 726, 967, 219, 632, 778, 758, 953, 978, 618, 753, 1014, 82, 38, 271, 688, 782, 855, 899, 907, 692, 221, 643, 284, 424, 614, 591, 575, 579, 508, 587, 502, 605]. Ethnicity [936]. Euclidean [896]. Euler [33]. Evaluating [642, 110]. Evaluation [791, 239, 681, 159, 672, 118]. even [572]. Event [1015]. Events [422, 574]. Evidence [53]. Evidential [320]. Evolutionary [227]. Exact [739, 414, 1023, 686, 569, 194, 562, 529, 125]. Exceedance [285]. Exchange [783, 529]. Excluding [213]. Exemplified [834].

Existence [906, 750, 348, 361, 725, 21]. Expanding [182]. Expansion [896]. Expansions [689, 822]. Expectation [757, 982]. Expected [345, 895]. Expected-Posterior [895]. Expectile [872, 814]. Expenditure [649]. Experiment [205, 673, 824]. Experimental [202, 389, 24, 211, 731, 35, 114]. Experimentation [26, 210]. Experiments [680, 261, 204, 431, 384, 353, 43, 810, 207, 679, 209, 660, 363, 677, 278, 534, 481, 533, 457, 531, 493, 529, 522]. Explaining [396]. Explanatory [318, 370]. Explicit [926]. Exploring [5]. Exponential [297, 638, 146, 630, 997, 623, 776, 197, 721, 860, 761, 941, 701, 797, 884, 961, 172, 177, 742, 978, 418, 910, 440, 760, 770, 848, 221, 935, 996, 599, 609, 610, 474]. Exponential-Polynomial [770]. Exponentiality [131]. Exponentially [59, 762]. Exponentiated [666, 732, 648, 882]. exposure [498]. expression [590]. Extant [284]. Extended [297, 446, 524, 577, 988, 463, 484]. Extension [622, 690, 803, 89, 417, 540, 589, 586]. Extensions [597]. Extra [95]. Extrapolation [633]. Extrema [1016]. Extremal [1023]. Extreme [289, 55, 393, 6, 423, 286, 678, 871, 857, 422, 167, 691, 311, 304, 574, 474]. Extremes [292, 484]. Extropy [1024].

Facilitate [783, 681]. Factor

[366, 816, 78, 44, 558, 568]. Factorial [261, 203, 147, 384, 735, 353, 160, 916, 43, 293, 437, 810, 436, 7, 1005, 31, 669, 693, 828]. Factorials [263, 265, 121]. Factors [397, 819, 435, 924, 850]. Fails [764]. Failure [55, 307, 661, 417, 648, 713, 917, 1015, 633, 919, 595, 586]. False [1021, 1019, 534]. Families [38, 453]. Family [1024, 991, 383, 969, 754, 197, 632, 132, 742, 882, 808, 350, 82, 848, 1012, 972, 424, 523].

Farlie [979, 466]. Fast [191]. Fatigue [417, 718]. Fatou [234]. Fault [681]. Feedback [164, 516]. Fenton [387]. Fermi [883]. fibers [601]. Fiducial [964]. Fiducial-Based [964]. Field [69]. Fields [376]. Filling [818, 1025, 287]. Filtered [139]. Filtering [781]. finance [54, 20]. Financial [662, 78]. Finding [354, 174]. Fine [113]. Finite [894, 851, 409, 447, 793, 667, 172, 746, 305, 82, 899, 907, 337, 960, 538, 579, 581, 496, 482, 582]. Fire [395]. First [303, 963, 661, 914, 356, 648, 398, 165, 499]. First-Failure [661, 648]. First-Order [303, 914, 356, 499]. Fisher [341, 34, 46, 185, 835]. Fit [6, 126, 979, 842, 934, 886, 127, 134, 775, 133, 124, 759, 135, 586, 550]. Fitting [401, 615, 342, 859, 683, 863]. Five [332, 6, 734]. Fixed [197, 731, 363, 918, 44, 646, 535, 567]. FL [387]. Fleishman [715]. Fleming [539]. Flexible [690, 950, 986, 335]. Flowgraph [182]. Focused [772]. Focusing [1021, 1019]. Fokker [727]. Foldover [436, 356]. Follow [436, 211, 1006]. Follow-Up [436, 211, 1006]. Force [276]. Forecasting [151, 399, 59, 551, 451]. Forests [977]. Formulas [379]. Formulation [952]. Four [734, 1012, 284, 981]. Four-Parameter [1012]. Fractal [983]. Fraction [422, 604]. Fractional [261, 147, 11, 384, 353, 160, 916, 293, 436, 31, 748, 121]. Fractions [617]. Frailty [8]. Framework [385, 1026, 718, 182, 483, 576]. Frameworks [743, 871]. Frank [20]. Fréchet [416]. Free [979, 916, 293, 438, 167]. Frequencies [295]. Frequency [200, 311, 974]. Frequency-Domain [974]. Frequentist [134, 908]. Frequentist-Bayes [134]. Frontier [854]. Full [437, 552]. Function [841, 269, 222, 872, 888, 379, 750, 8, 820, 797, 884, 896, 881, 795, 774, 651, 548, 489]. Functional [559, 703, 386, 779, 761, 376, 832, 812, 185, 492, 458]. functionals [462].

Functions [99, 1020, 166, 251, 777, 644, 64, 631, 772, 407, 326, 1027, 600]. Funnel [677]. Further [276, 691, 337, 835]. fused [519]. Fusion [410, 392]. Future [621, 736, 713]. Fuzzy [170, 1027, 584].

G [12, 523, 631]. GA [31]. GA-optimal [31]. Gambler [51]. Game [192, 461, 491]. Gamma [666, 434, 636, 834, 820, 682, 416, 131, 964, 297, 1]. Gaussian [455, 14]. gene [590]. General [443, 279, 806, 750, 732, 695, 82, 855, 365, 513, 583]. Generalised [249, 133]. Generalization [792, 380, 418, 1012]. Generalized [878, 991, 887, 962, 328, 997, 173, 734, 866, 834, 697, 439, 824, 128, 177, 910, 474, 143, 625, 812, 999, 782, 838, 272, 104, 212, 943, 421, 483, 581, 487, 599, 466, 574, 550, 606]. Generated [384, 656, 484]. Generating [777, 208, 395]. Generation [963, 1026]. Generative [1026]. Generators [827, 355]. Genetic [40]. Genome [39]. Genomics [48]. Geo [379]. Geo-Stable [379]. geographical [489]. Geometric [7, 56, 599, 474]. GI [12]. Giant [706]. Gini [366]. Given [435, 751, 938]. GLMM [654]. Global [724, 683]. Glued [408]. GMANOVA [836]. GMANOVA-MANOVA [836]. Gompertz [523, 967, 557, 1012, 1027]. Gompertz-G [523]. Good [287]. Goodness [979, 842, 934, 886, 127, 133, 124, 759, 135, 550]. Goodness-of-Fit [842, 934, 886, 550]. Gradual [783]. Graeco [745]. Gram [689]. Graph [227, 262, 793]. graphical [547]. Graphs [833, 825, 227, 273, 1002]. Graphs- [227]. Green [234]. grid [495]. Group [308, 641, 993, 243]. Group-wise [641]. grouped [519]. Groups [105, 312, 258, 829, 115]. Growth [615, 718, 557]. Guaranteed [905]. Guarantees [341]. Guide [314]. Gumbel [979, 904, 466]. Gumbel-Morgenstern [466]. Gupta [123].

H [1027]. H-Bayesian [1027]. Hadamard [916, 85, 810, 350]. Haenszel [597]. Half [380]. Hall [20]. Handle [874]. Hanif [657]. Happen [948]. hardcover [387]. Harrington [539]. Hauck [988]. Hazard [99, 1020, 72, 721, 971, 376, 586, 502]. Hazards [846]. Health [401, 119, 336]. Heavy [887, 401, 982, 729]. Heavy-Tailed [887, 401, 982]. Hedayat [706]. Hellinger [875]. Heredity [712]. Heston [627]. heterogeneous [610]. Heteroscedastic [767, 738, 659]. Heteroscedasticity [443, 40, 601]. Heuristically [370]. Heyde [57]. Hidden [950, 170, 591, 485]. Hierarchical [712, 791, 1010, 698, 813, 395]. Hierarchies [96]. High [443, 652, 756, 286, 678, 783, 63, 542, 385, 311, 938, 949, 835]. High-Dimensional [443, 652, 385, 949, 835, 542]. High-Frequency [311]. High-Order [783]. Higher [789, 456]. Highly [957]. Highway [236]. Hilbert [22, 232, 761]. Histone [325]. Historical [704]. Hit [963]. Hockey [784]. Homogeneity [803, 999, 243, 304, 556]. Homogeneous [305, 852, 94]. homoscedasticity [567]. horticultural [571]. Hosking [642]. Household [641]. Household-Level [641]. Hunter [147]. Hurdle [986]. Hybrid [674, 910, 686, 956, 451]. hydrology [574]. hyper [487]. hyper-Poisson [487]. Hyperbolic [689]. Hypercube [1025, 348, 1004, 208]. Hypercubes [287]. Hyperparameter [23]. Hypoexponential [935]. Hypotheses [443, 796, 319, 865, 1021, 1019, 155]. Hypothesis [315, 449, 740, 723, 853, 968, 626, 938].

I/II [335]. IASRI [123]. ICER [283]. Identifiability [944]. Identification [353, 103, 749]. Identifying [641, 476, 850].

II [705, 887, 726, 623, 696, 1025, 609, 713, 686, 577, 994, 929, 335]. III [658]. Illustration [884]. Illustrations [961, 571]. Image [1022]. Imagery [909]. Images [109, 755]. Immediately [276]. Immigration [3]. Impact [47, 665]. Implement [712]. implementation [499]. Implementing [1010]. Importance [144]. Important [836]. Imposition [90]. Imprecise [449, 99, 96, 98, 101]. Imprecision [87, 104]. Improved [146, 709, 270, 508, 505, 990, 984, 78, 14, 688, 221, 424, 675, 579, 566, 634]. Improvement [800]. Improving [368, 253, 818, 953, 691]. Imputation [789, 323, 201, 107, 224, 628, 760, 874]. imputing [605]. INARMA [635]. Incidence [407]. Inclusionwise [117]. Income [196]. Incomplete [833, 825, 301, 913, 30, 900, 29, 266]. Incomplete-Block [833, 825]. Incorrect [1021, 1019, 541]. Incremental [283, 818]. Independence [307, 414, 938]. independent [453, 1, 610]. Index [724, 423, 366, 871, 616, 231, 857, 722, 576]. Indexes [953, 635]. India [123]. Indices [53]. Indirect [222, 498]. Indirectly [314]. Individual [698, 813, 477]. Individualized [312]. Individuals [647]. Induced [408, 373, 531]. Inequalities [873, 761, 794]. Inequality [1011, 452]. Infections [47]. Inference [621, 622, 411, 298, 428, 400, 700, 736, 448, 630, 171, 299, 866, 195, 291, 246, 661, 624, 325, 785, 686, 658, 224, 770, 994, 929, 790, 643, 104, 821, 919, 573, 569, 494, 486]. Inferences [623, 380, 107, 801]. Inferential [518]. inferiority [176]. Infested [71]. Inflated [703, 275, 1014, 964, 475]. Influence [849, 862, 802]. Influential [970]. Information [152, 430, 409, 52, 462, 632, 250, 141, 708, 681, 984, 772, 716, 14, 38, 156, 271, 688, 907, 179, 655, 989, 530]. Information-Based [655]. Informative [135, 611]. Inhomogeneous [820]. Initial [227]. Injuries [986]. innovations [518]. Input [397]. Insights [1011]. Inspired [629]. Insurance [886, 898, 890, 898]. Integer [303, 606]. Integer-Valued [303, 606]. Integral [398]. Integrals [75]. Integrated [911, 812]. Integration [239, 288]. Interacting [665]. Interaction [397, 142, 558, 568]. Interactions [660, 528]. Intercept [771]. Interest [276]. Interface [831]. interior [457]. interlaboratory [601]. Interpolation [67, 651]. Interpolator [831]. Interpretability [95]. Intersecting [255]. intersection [540]. Intersections [13]. Interval [663, 146, 806, 860, 241, 918, 465, 562]. Interval-Censored [860]. Intervals [726, 879, 94, 984, 394, 964, 535]. Intervened [343]. Intra [806]. Intra-class [806]. Intraclass [241]. Introducing [969, 624]. Introduction [58, 124]. Invariance [834, 780]. Invariant [269, 556, 576]. Inverse [638, 928, 416, 14, 994, 405, 707]. Inverted [720, 732, 648]. Investigating [116]. Investigation [792, 1022, 746]. Involves [205]. Involving [167, 195, 575]. Iraq [551]. irreducible [538]. ISBN [387]. Ising [230]. Isolation [742]. Isomorphism [669]. Isotones [130]. Issue [921, 411, 250, 815]. Iterations [37]. IV [916].

J [137]. Jack [45]. Jackknifed [244]. jackknifing [464]. Jagdish [347]. Jarque [385]. Jeffreys [564]. Jeffreys-type [564]. John [137]. Johnson [897]. Joint [419, 276, 109, 436, 329, 145, 41, 275, 161]. Jointly [929]. Journal [1029]. Journey [425]. Judgments [98].

Kaplan [464, 339]. Kernel [763, 150, 857, 645, 761, 966, 875, 639, 642, 217, 306, 492, 548, 472]. Kernel-Based [639, 642]. kernels [490]. Key [368]. Kiefer [45]. Known [434, 894]. Kolmogorov [700, 381]. Kotz [662]. Kullback [52]. Kumar [634]. Kumaraswamy [726, 504, 631, 302]. Kumaraswamy- [726]. Kumaraswamy-G [631]. Kumaraswamy-type [504]. Kurdistan [551]. Kurtosis [313].

L [454]. Lack [134, 775, 885, 891, 944]. Lack-of-Fit [134, 775]. Lagged [954]. Lam [671]. Laplace [1, 134, 468]. Large [654, 727, 1023, 223, 852, 556, 607]. large-dimensional [556]. large-scale [607]. Lasso [1010, 519, 685, 872]. late [539]. Latent [973, 288, 78, 473, 554]. Latin [1025, 1004, 208, 287, 745, 725, 981]. Lattice [12, 30]. Law [223]. Laws [1023, 852]. layer [1002]. Layout [767, 62, 44]. Learners [764]. Learning [788, 792, 807, 1022, 755]. Least [274, 627, 1021, 1019, 59, 471]. Least-Squares [274, 627]. Left [630, 458]. left-truncation [458]. Leibler [52]. Length [152, 930, 989, 515]. Length-Biased [152, 930, 989, 515]. Leone [705]. Leone-Power [705]. Less [894, 898]. Letter [86]. Level [680, 367, 358, 353, 734, 1007, 436, 1005, 242, 641, 369, 442, 931, 617, 278, 822, 365, 214, 568, 592, 160, 120, 528, 162, 186]. Levels [286, 341, 532]. leverage [1018]. Levy [748, 922, 486]. Library [267]. Life [741, 865, 623, 777, 57, 237, 10, 933, 137, 45, 34, 46, 657, 706, 596]. Lifetime [705, 16, 119, 73, 553, 510]. Lifetimes [762, 882]. Like [778, 689]. Likelihood [878, 300, 412, 318, 319, 423, 194, 309, 877, 696, 787, 820, 414, 626, 624, 327, 868, 320, 643, 165, 945, 61, 317, 394, 996, 614, 578, 525, 503, 500, 577, 564]. Likelihood-Based [643]. Likelihoodist [320]. Limit [406]. Limited [951]. Lindley [752, 720, 864, 702, 850, 985]. Line [830, 262]. Linear [443, 1001, 11, 970, 872, 248, 171, 173, 806, 927, 839, 973, 293, 834, 697, 824, 23, 159, 665, 771, 305, 917, 625, 988, 337, 272, 212, 118, 269, 559, 1, 903, 195, 820, 570, 452, 512, 567]. Linex [571]. Link [335, 591]. link-tracing [591]. Linking [9]. Literature [848]. Liu [1001, 970]. liver [565]. Ljung [642]. Local [11, 962, 197, 1025, 559]. Localization [681]. Locally [772, 693, 828]. Locating [252]. Location [556, 650, 619, 576, 132, 941, 646, 643]. Location-invariant [556, 576]. Location-Scale [132]. Locations [961]. Lockhart [137]. Log [139, 803, 636, 820, 837, 926, 53, 808, 633, 651, 867, 481, 592]. Log-Burr [636]. Log-Log [803]. Log-Logistic [837, 926, 867, 481]. Log-periodogram [139]. Log-Symmetric [808]. log-transformed [592]. Logarithmic [940, 420, 798, 418, 526]. Logistic [878, 901, 127, 333, 837, 926, 641, 772, 691, 633, 61, 867, 481]. Logit [99, 530]. Lomax [705, 246, 610]. Long [139, 525]. long-memory [525]. Longitudinal [329, 326, 563]. Look [110]. Loss [872, 166, 708, 1027]. Losses [890]. Lot [277]. Low [833, 825, 818]. Lower [93, 365, 583]. Luminosity [141].

M [237]. M. [237]. Machines [381]. Magic [913]. Mahalanobis [869]. Main [367, 672, 568, 528, 572]. Main-Effect [367]. Making [106, 551]. management [504]. MANOVA [836]. Mantel [597]. Mapping [1026]. Maps [395]. Marginal [415, 309, 803, 971, 999, 606]. Marginals [751, 702, 494]. margins [545]. Market [315, 675]. Markov [538, 467, 83, 50, 783, 1028, 950, 465, 220, 233, 914, 852]. Markov-chain [467]. Markovian [775]. Marrow [961]. Marshall [892, 383, 630, 619]. Martin [387]. Martingale [718, 73]. Masking [653]. mass [553]. Matched [112]. Matching [348]. Mathematical [383]. Matrices [916, 85, 810, 254, 955, 350, 852, 556, 613].

Matrix [175, 975, 790, 947, 959, 958, 530]. Matrix-Variate [975]. Matter [245]. Matters [944]. Maximal [269, 873]. Maximin [352, 208]. Maximization [60]. Maximizing [345]. Maximum [878, 614, 423, 435, 877, 696, 787, 820, 348, 327, 731, 165, 945, 61, 996, 577, 516]. Maximum-Likelihood [878]. Maxwell [986, 960, 879, 166, 480, 751, 515, 707]. McDonald [446, 577]. MCMC [674]. MDS [354]. Mean [430, 789, 419, 969, 777, 747, 871, 63, 709, 197, 632, 447, 975, 893, 1010, 797, 746, 978, 618, 398, 597, 634, 955, 781, 82, 38, 156, 271, 628, 760, 880, 907, 179, 676, 862, 960, 744, 551, 579, 581, 587, 505, 460, 582]. Mean-of-Order [871]. Mean-Parameterized [960]. Mean-Variance [969, 781]. Means [148, 178, 945]. Measure [285, 952, 708, 995, 885, 891, 14, 225, 506, 593, 509]. Measurement [849, 940, 970, 323, 893, 978, 634, 156, 880, 862, 744, 581, 505]. Measurements [640, 477]. Measures [1024, 726, 232, 196, 144, 542, 493]. Measuring [724, 314, 243]. Mechanisms [73]. Median [283, 594]. Median-Based [283]. Mediation [699]. Medical [909, 286, 1022]. Medicine [698, 813]. Meier [464, 339]. Melanoma [804]. membership [158]. Memoriam [57]. Memory [921, 139, 881, 525]. Memoryless [904]. Method [390, 622, 685, 281, 386, 857, 266, 710, 868, 143, 181, 760, 699, 749, 585]. Methodologies [239, 990]. Methodology [374, 444]. Methods [915, 666, 1009, 894, 332, 789, 9, 149, 1019, 401, 747, 912, 246, 259, 58, 942, 968, 624, 713, 176, 124, 440, 628, 874, 976, 949, 335, 467, 480, 518, 452, 595, 1021]. Metric [952]. metrics [54]. Michael [137]. Microarray [109, 207]. Microbiome [847]. Microdata [749]. Middle [448]. Middle-Censored [448]. Minima [727]. Minimal [117, 843, 640]. Minimal-Variance [843].

Minimally [931]. Minimization [488]. Minimizing [951]. Minimum [734, 1007, 916, 293, 437, 810, 533, 775, 701, 765, 884, 961, 875, 731, 416, 404, 365, 919, 583]. Mining [729, 986]. minus [598]. Misclassification [633]. Misclassified [332]. Mises [125, 866, 779]. Missing [432, 412, 441, 201, 948, 955, 275, 628, 874, 855, 946, 503, 602, 500, 605]. Mixed [680, 970, 806, 163, 973, 377, 242, 369, 731, 78, 1015, 801, 733, 118, 908, 546, 992, 530]. Mixed-Level [680, 242, 369]. Mixed-State [78]. Mixing [645]. Mixture [32, 752, 894, 851, 858, 969, 719, 197, 342, 793, 457, 172, 882, 733, 427, 717, 563]. Mixtures [779, 801, 960, 484]. ML [743]. MLE [847]. mobility [856]. modal [492]. Mode [406, 639]. Model [32, 432, 1001, 1017, 274, 393, 415, 332, 627, 269, 99, 622, 1020, 836, 150, 979, 222, 138, 623, 783, 309, 60, 803, 703, 615, 382, 927, 168, 636, 973, 52, 230, 787, 244, 192, 860, 308, 933, 775, 761, 778, 804, 334, 898, 417, 853, 732, 282, 859, 325, 641, 667, 172, 159, 785, 762, 882, 920, 718, 398, 716, 682, 275, 686, 625, 956, 1014, 1006, 823, 999, 954, 850, 182, 635, 324, 114, 101, 212, 118, 80, 863, 957, 919, 331, 867, 545, 558, 592, 458, 557, 593, 570, 478, 553, 563, 510, 500, 595, 577]. model [560, 566, 541, 564, 530]. Model-Assisted [667]. Model-Inference [325]. Model-Robust [823]. Modeling [152, 847, 887, 419, 113, 322, 719, 286, 483, 151, 200, 8, 109, 231, 288, 329, 422, 954, 324, 986, 946, 551, 473, 554, 485]. Modelling [769, 950, 759]. Models [300, 310, 215, 663, 970, 638, 858, 122, 419, 872, 383, 322, 754, 351, 433, 791, 950, 806, 27, 163, 911, 195, 183, 839, 342, 834, 936, 697, 155, 447, 333, 36, 914, 853, 824, 23, 1026, 327, 76, 352, 890, 119, 961, 145, 42, 771, 698, 813, 344, 772, 211, 81, 731, 78, 642, 885, 891, 714, 801, 633, 812, 944, 395, 272, 165, 212, 717, 61, 744, 908, 854, 121, 549, 578, 546, 903, 511, 525, 593, 451, 463, 586, 475, 491, 584, 567]. Moderate [1021, 1019]. Modern [124, 35]. Modes [773]. Modification [386, 337]. Modifications [325]. Modified [915, 652, 720, 876, 385, 439, 571, 716, 838, 989, 596, 536]. MOGN [909]. Moment [789, 777, 128, 794, 544]. Moments [281, 904, 730]. Monitoring [676]. Monograph [85, 123]. Monotone [955, 795, 503, 602, 564]. Monte [467, 60]. monthly [551]. Morgenstern [466]. Morgernstern [979]. Mortality [399]. Moves [618]. mRNA [49]. MRPE [884]. MSE [826]. MSE-Optimal [826]. MTD [110]. Multi [25, 162, 889, 961, 785, 999, 781, 243, 425, 214, 831]. Multi-Armed [889]. Multi-category [243]. Multi-Environment [25]. Multi-level [162]. Multi-period [781]. Multi-stage [961]. Multi-Step [785, 831]. Multi-Touch [425]. Multi-way [999]. Multicollinearity [248, 244]. Multicomponent [732]. Multicriteria [189]. Multilevel [924, 101]. Multinomial [180, 765, 507]. Multiple [290, 796, 872, 312, 323, 107, 933, 417, 928, 1015, 774, 224, 184, 683, 588, 570, 512, 490]. Multiplicative [775, 642, 880]. multiplicity [479]. Multisample [414]. multistage [594]. Multivariate [297, 432, 215, 113, 719, 783, 379, 840, 843, 239, 869, 385, 966, 771, 845, 56, 78, 918, 240, 311, 780, 985, 974, 957, 854, 540, 545, 470, 548, 503, 452]. Multiway [819]. Mutant [227]. Mutually [256, 1013, 255].

N [347]. Nakagami [729]. Nath [19]. natural [498]. navigation [1029]. NBA [786]. NBRUL [865]. NBRUL- [865]. Near [194, 998, 141, 414, 287, 153]. Near-Exact [414, 194]. Near-Orthogonal [287]. Nearly [687, 913]. needed [595]. Negative [126, 804, 1000, 961, 980, 768, 620]. Negatively [66]. Neighbor [74, 745]. Neil [387]. Nelson [187]. Nephropathy [888].

Nested [831, 903]. nested-error [903]. Nesting [27, 163]. Network [314, 670, 1002]. Networks [841, 983, 792, 809, 1026, 100, 387]. Neural [841, 983, 792, 787, 809, 1026, 1002]. Neuroblastoma [427]. Neurophysiological [710]. Newton [878]. Neyman [114]. Nikodým [232]. Noise [109, 1008, 974]. Noises [738]. Nominal [952]. Non [849, 796, 763, 768, 67, 160, 195, 871, 21, 7, 893, 844, 176, 746, 738, 620, 82, 240, 688, 782, 899, 852, 744, 634]. Non-equivalence [7]. Non-existence [21]. Non-Homogeneous [852]. Non-inferiority [176]. Non-linear [195]. Non-negative [768, 620]. Non-parametric [67]. Non-regular [871]. Non-Response [82, 688, 782, 899, 849, 893, 746, 240, 744, 634]. Non-rotatable [844]. Non-standard [763, 738]. Non-symmetric [796]. Nonadditive [88]. Nonbinary [517]. Noncentral [175]. Noncentralities [373]. Nonexistence [906]. Nonlinear [151, 76, 78, 801, 908, 451, 500]. Nonnormality [443, 536, 544]. Nonparametric [647, 621, 70, 180, 650, 987, 150, 298, 428, 105, 149, 400, 700, 736, 322, 129, 299, 280, 930, 747, 465, 17, 937, 64, 900, 377, 624, 554, 667, 764, 845, 748, 188, 759, 407, 944, 73, 284, 326, 573, 477, 494, 458, 473, 502, 490]. Nonparanormal [793]. Nonrandomized [279, 382]. Nonregular [147]. Nonresponse [694, 107, 440, 505]. Nonstationary [678, 2, 235, 518]. Normal [297, 290, 99, 969, 370, 380, 281, 338, 239, 413, 973, 130, 439, 178, 896, 771, 56, 716, 918, 801, 646, 80, 854, 867, 481, 503, 514, 452, 571]. Normal-Based [854]. Normality [374, 313, 130, 385, 780, 567]. Normalized [916, 810]. Norman [387]. Note [262, 60, 927, 52, 391, 37, 761, 853, 356, 210, 287, 634, 470, 903, 466, 584]. Novel [940, 285, 176]. NSGA [1025]. NSGA-II [1025]. Null [723]. Number [47, 1021, 1019, 403, 773, 435, 37, 511, 572]. Numbers [647, 1023, 852, 532]. Numeric [379]. Numerical [653, 998].

Objective [155]. Observation [432, 522]. Observational [336, 821]. Observations [627, 970, 63, 1008, 511]. Observed [8, 314, 875, 850, 919, 554]. Obtain [800]. Occasion [746, 440]. Occasions [316]. Odd [867, 481, 510]. Odds [1017, 176, 651]. Odds-Symmetry [1017]. Odisha [641]. Offenders [103]. Offer [345]. Oil [678, 856]. Olkin [892, 383, 630, 619]. Olympic [289]. Omission [665]. Oncology [658]. One [752, 603, 767, 1021, 1019, 62, 233, 731, 44, 553, 607]. One- [603, 731]. One-Parameter [752]. one-sided [607]. One-Way [767, 62, 44]. Online [425]. Open [168, 296]. Operations [368, 121]. Opposite [972]. Optical [254]. Optimal [32, 668, 825, 277, 390, 663, 384, 206, 263, 568, 735, 367, 351, 433, 997, 374, 767, 345, 1011, 65, 142, 948, 834, 697, 673, 645, 369, 26, 826, 660, 670, 698, 174, 112, 829, 482, 330, 899, 359, 372, 684, 404, 655, 828, 212, 522, 214, 675, 558, 1018, 28, 226, 207, 598, 529, 693, 833, 354, 813, 31]. Optimality [819, 789, 262, 67, 807, 37, 209, 923, 305, 252, 517, 572, 939]. Optimization [102, 1025, 677]. Optimizing [990]. Optimum [870, 615, 153]. Option [183]. Optional [447, 893, 157, 282, 885, 891, 744]. Options [83]. Order [769, 432, 621, 789, 189, 303, 876, 358, 783, 871, 616, 444, 140, 914, 356, 961, 442, 660, 785, 188, 918, 295, 165, 365, 863, 265, 576, 583, 499, 456, 469]. Order-Restricted [785]. Order-Statistic [295]. Ordered [1017, 310, 312, 941, 44]. Orderings [736, 971]. Ordinal [301, 299, 803, 869]. Ordinary [195]. Ornstein [787]. Orthogonal [32, 360, 819, 827, 22, 122, 358, 734, 1004, 254, 208, 287, 931, 733, 828, 863, 121, 598]. Orthogonality [1025]. Orthogonalized [118]. Orthogonally [442]. Other [819, 906, 950, 273, 89]. Our [741, 236]. Out-of-Sample [392, 410]. Outcome [946, 335, 611, 555, 498]. Outcomes [364]. Outlier [459, 71, 578]. Outliers [1001, 244, 869, 177]. outlying [608]. Output [397]. Overdispersed [703, 877]. Overlapping [246]. Overview [108, 120, 229]. OxyELITE [565].

P [123]. Pabitra [123]. package [537]. Page [441]. Paillier [956]. Pair [754, 676]. Pair-copula [754]. Paired [660, 528, 992]. Pairs [174, 112, 225]. Pairwise [711, 309, 361]. Pakistan [924]. palindromic [463]. Panel [742, 1015, 790, 530]. Paradigm [318, 319, 320, 317]. Parallel [769, 55, 830, 291]. Parameter [752, 650, 390, 146, 166, 138, 674, 967, 616, 444, 778, 990, 223, 422, 14, 1012, 646, 1027, 421, 480, 609, 486]. Parameterization [672]. Parameterized [960]. Parameters [878, 434, 841, 409, 281, 701, 59, 771, 643, 170, 1002, 596, 496, 500, 577, 502]. Parametric [573, 622, 303, 401, 403, 933, 667, 136, 67, 973, 4]. PARDs [711]. Paretian [315]. Pareto [757, 249, 619, 96, 696, 1023, 713, 574, 996, 421]. Pareto-Type [757, 1023, 713]. Parsad [123]. Parsimonious [1017]. part [517]. Partial [189, 528, 436, 710, 103, 564]. Partial-profile [528]. Partially [301, 623, 913, 31, 136, 993, 919, 456]. Particles [113]. Partition [98, 184]. Partitioned [962, 927]. Passage [398]. Path [12, 273, 81]. Paths [769]. Patient [1026]. Patients [312, 901, 884]. Pattern [427, 455]. Pattern-Mixture-Type [427]. Patterned [798]. Patterns [924, 316]. PBIB [870]. PCA [848]. PCR [228, 758]. PCR-Ridge [758]. PDHS [924]. Pearson [723, 114]. Penalized [954]. Penalty [270, 1010]. perceived [476]. Percentiles [837, 363]. Perceptions [236]. Perfect [1013]. Performance [691]. Period [12, 141, 781]. Periodic [67, 950, 151]. Periodical [168, 296]. Periodogram [778, 139]. Periodogram-Type [778]. Periods [640]. Permitting [441]. Permutation [739, 540]. Permutation/Randomization [739]. Permutations [617, 822]. Persons [936]. Perspective [727, 554]. Perturbation [390]. Phase [893, 658, 82, 628, 362, 335, 575]. phases [500]. Physical [211]. Piecewise [860]. Pilot [824]. Placement [227]. Plan [277, 910, 268, 169, 482]. Planck [727]. Plans [687, 819, 367, 800, 213, 572]. Player [192, 461]. Pless [256]. Plot [29, 266, 30]. plots [547]. Plug [402]. Plug-In [402]. PMC [758]. Point [1020, 902, 190, 710, 797, 884, 961, 258, 742, 198, 716, 917, 774, 646, 989, 72]. Point-Regular [258]. Points [872]. Poisson [666, 614, 342, 155, 518, 343, 487, 478, 853, 751, 864, 730, 216, 475, 985, 635, 986, 960, 606]. Polygonal [161]. Polynomial [432, 11, 15, 352, 24, 770]. Polytopes [366]. Pooling [267]. Population [789, 245, 925, 911, 340, 409, 219, 709, 632, 721, 447, 667, 746, 978, 618, 305, 634, 82, 271, 628, 760, 899, 907, 862, 802, 945, 424, 591, 579, 581, 587, 505, 496, 482, 582]. Populations [290, 434, 245, 168, 296, 246, 941, 178, 896, 337, 304, 503]. PORT [444]. Portfolio [390, 781, 675, 483]. Portmanteau [642]. Positive [423]. Possible [531]. Possibly [692, 541]. Post [749]. Post-randomization [749]. Posterior [654, 224, 895]. Potential [911, 102, 966]. Poverty [641]. Power [705, 805, 318, 860, 679, 128, 710, 223, 275, 131, 895]. pp [387]. Practical [1011, 445, 314, 585]. Practice [9, 87]. Pre [174]. Pre-set [174]. Precedence [105].

Precision [698, 813]. Predicting [784]. Prediction [113, 322, 791, 1028, 410, 392, 648, 922, 713, 698, 813, 574, 78, 603, 488]. Predictions [839, 285]. Predictive [647, 621, 70, 180, 298, 428, 494, 105, 149, 400, 700, 736, 299, 280, 624]. Predictor [665, 81, 272]. Predictors [248]. Preface [921, 411, 561, 346, 388, 815, 527]. Preliminary [726, 582]. Preparations [830]. Prescriptive [97]. Presence [152, 940, 630, 177, 978, 634, 82, 156, 240, 899, 179, 693, 828, 558, 596, 581, 505, 604, 541, 564]. Present [893]. Press [387]. prevention [539]. Previsions [93]. Price [678, 345, 856]. Prices [649]. Pricing [183]. Prime [804, 121]. Priming [98]. Principal [403, 386, 958]. principle [516]. Prior [415, 638, 681, 14, 104, 479, 564]. Prior-data [104]. Priorities [925]. Prioritization [189]. Priority [18]. Priors [431, 155, 895, 507]. Privacy [885, 891, 593]. Private [936]. Pro [565]. Probabilistic [158, 285, 100]. Probabilities [88, 239, 413, 766, 977, 98, 326]. Probability [47, 449, 368, 94, 379, 1021, 1019, 371, 187, 285, 333, 18, 765, 92, 56, 77, 812, 54, 852, 284, 569, 501]. Problem [72, 357, 341, 695, 928, 198, 185]. Problems [108, 280, 190, 795, 835, 100, 573]. Procedure [397, 341, 853, 178, 746, 728, 633]. Procedures [886, 631, 795, 988]. Process [449, 66, 303, 963, 228, 820, 128, 710, 922, 223, 748, 295, 676, 495, 518, 460, 486, 606]. Processes [11, 15, 94, 171, 139, 3, 75, 1016, 761, 233, 333, 875, 191, 211, 2, 20, 235, 635]. Processing [909]. Product [199, 978, 82, 156, 362, 221, 575, 544]. Product-Convolution [199]. Production [49, 489]. Prof. [657]. Professor [815, 921, 741]. Profile [756, 528]. Progressive [726, 623, 661, 732, 994, 929, 609, 574, 577]. Progressively [997, 648, 603]. Project [943]. Prolongation [116]. Proof [234]. Properties [752, 991, 383, 426, 480, 888, 380, 338, 439, 487, 287, 737, 515, 707, 725, 114, 612, 591]. Property [904]. Proportion [279]. proportional [586]. Proposed [377]. Propriety [325]. Protein [49]. Pseudo [318, 484]. Pseudo-convex [484]. Pseudo-Likelihood [318]. PSO [451]. Publication [123]. Pulled [704]. Pulled-to-Par [704]. Purchase [425]. Pure [1007]. Purely [797, 884, 571]. Purging [699]. Pyramid [273].

QAL [16, 339]. Qin [835]. QML [743]. QT [116]. Quadratic [32, 1007, 832]. Quadruple [258]. Qualitative [673, 100, 558]. Qualitatively [453]. Quality [119, 268, 800, 16]. quantal [534]. Quantile [915, 685, 420, 526, 937, 809, 664, 798, 814, 1010, 519, 600]. Quantiles [292, 609]. Quantitative [673, 77, 558, 593]. Quantities [649]. Quartile [676]. Quasi [108, 254]. Quasi-Difference [254]. Quasi-Residual [108]. Question [13]. Question-sets [13]. Questions [956]. Queue [164, 157, 516]. Queues [12].

R [10, 815, 731, 537]. Race [936]. Radar [755]. Radhakrishna [25]. Radon [232]. Rainfall [200]. Rajender [123]. Ramifications [5]. Random [274, 218, 963, 138, 615, 776, 1016, 951, 616, 636, 948, 933, 207, 376, 79, 154, 157, 890, 977, 698, 813, 917, 38, 899, 897, 947, 170, 165, 989, 424, 946, 867, 506, 1002, 558, 458, 511, 557, 579, 462, 581, 512, 509, 584, 522]. Random-Effects [867, 469]. Randomised [1009, 965]. Randomization [739, 669, 749]. Randomized [450, 364, 447, 308, 893, 282, 77, 956, 658, 862, 539, 478, 482, 560]. Randomly [982, 406]. Rank [858, 420, 526, 435, 869, 900, 958, 326, 470, 605]. Rank-Tracking [326]. Ranked [915, 430, 671, 932, 978, 112, 167, 676, 919, 594, 454, 508]. Ranking [18, 691]. Rankings [711]. Rao [25, 237, 10, 815]. Raphson [878]. rare [478]. Rate [300, 1020, 1021, 1019, 72, 951, 721, 820, 971, 804, 762, 920, 917, 538]. Rate-Based [920]. Rater [711]. Rates [15, 783, 233, 483]. Rating [18]. Ratio [594, 283, 789, 194, 60, 447, 414, 626, 624, 176, 978, 181, 82, 156, 240, 688, 179, 221, 575, 503, 508, 605]. Ratio-Type [789, 508]. Rational [433]. Ratios [413, 507]. Raton [387]. Rayleigh [603, 648, 702, 550]. Re [17, 898, 168]. Re-classifications [168]. Re-Estimation [17]. Reactions [951]. Real [345, 414]. Real-Estate [345]. Reality [766]. Rearrangement [144]. Reassessment [335]. Recapture [219]. Reclassifications [296]. recognition [455]. Reconciling [95]. Record [1024]. Records [631]. Recovery [55, 483]. Rectangles [913, 725, 981]. Recurrent [1015]. Recursive [253, 239, 857, 645, 288, 59, 875, 81, 492]. Redistribute [328]. Redistribute-to-the-Right [328]. Reduced [272, 576]. reduced-bias [576]. Reducing [595]. Reduction [1003, 616, 109, 288, 966, 470, 471, 590]. Reductions [230]. Reference [277]. Reflection [45]. Regime [408]. Regimes [427]. Region [551]. Regions [310, 701, 268, 601, 500]. Regression [769, 1001, 274, 671, 861, 269, 1020, 150, 419, 67, 872, 248, 129, 351, 433, 173, 138, 139, 694, 340, 747, 703, 901, 927, 709, 839, 636, 270, 937, 244, 809, 158, 373, 64, 761, 804, 814, 1010, 859, 119, 641, 667, 722, 665, 718, 41, 772, 198, 682, 832, 216, 691, 1015, 714, 829, 1014, 759, 688, 405, 954, 850, 895, 284, 717, 61, 331, 867, 492, 578, 559, 1018, 470, 546, 520, 519, 458, 548, 471, 570, 932, 563, 510, 582]. Regression-Based [861]. Regression-Type [405]. Regressions [438, 988, 960, 490]. Regret [889]. Regular [262, 750, 160, 616, 258, 617, 993, 365, 214, 871, 583]. Regularization [949]. Regularize [764]. Reject [345]. Rejection [1021, 1019]. Rejoinder [321]. Related [819, 906, 444]. Relation [141]. Relationship [405]. Relationships [723]. Relative [995, 176, 559]. relative-error [559]. Releasing [749]. Relevance [425]. Reliability [769, 106, 726, 631, 732, 172, 737, 101, 465, 600]. Reliability-Based [106]. Remainder [512]. remaining [574]. Reminiscence [19]. Remission [884]. Removal [769]. Renewal [1016, 172]. Repairable [1028]. Repairing [844]. Reparametrization [36]. Repeated [111, 84, 766, 640, 1008, 542, 493]. repeated-measures [493]. Replicated [353]. Replicates [121]. Replication [833, 825, 889]. Replications [668]. Reported [649]. Representation [15, 235]. Representations [24]. Reproducibility [1009, 965, 400, 700, 736, 624]. Research [145, 336, 110]. Residual [193, 108, 836, 777, 737, 640, 625, 603]. Residuals [118]. Resistant [810]. Resolution [687, 147, 916, 435]. Resolution-Rank [435]. Resolutions [31]. Resolvable [349, 259, 29, 993, 21]. respondent [593]. Response [432, 1009, 965, 374, 703, 370, 409, 447, 308, 893, 282, 81, 77, 956, 82, 688, 782, 899, 862, 683, 849, 534, 513, 478, 746, 500, 482, 521, 240, 560, 744, 634]. response-adaptive [513]. Responses [987, 673, 914, 814, 275, 693, 828, 513, 546]. Restricted [747, 785, 272]. Restrictions [669, 856]. Result [83]. Results [906, 25, 276, 334, 646]. Retrial [164]. Return [396]. Returns [662, 53, 704]. Reverse [873]. Review [9, 294, 371, 85, 186, 190, 387, 145, 123, 848, 684, 20, 707, 580, 501]. Revisit [339]. Revisited [1011, 132, 816, 753]. Revisiting [423, 13, 320]. reward [538]. Richard [137].

Richards [615]. Ridge [173, 244, 758]. Right [328, 428, 930, 982, 713]. Right-Censored [982, 713]. Right-Censoring [428]. Rigorous [749]. Risk [116, 410, 392, 92, 884, 961, 176, 704, 395, 850, 749, 943, 504, 387]. Risks [762, 407, 919]. Ristic [589]. Ritov [69]. Robust [202, 432, 1001, 90, 91, 245, 757, 102, 901, 948, 158, 464, 837, 926, 990, 823, 770, 103, 536, 585]. Robustness [621, 902, 95, 958, 544]. ROC [394]. role [479]. Ronald [34, 46]. Root [344, 165]. rotatable [844]. Rotation [316]. Row [711, 206, 830, 939, 923]. Roy [19]. RRT [593, 885, 891, 744]. Ruin [276, 51]. Rule [796]. Rules [941, 224]. Run [442, 598]. Runs [263, 931, 167, 572]. Rural [641]. RV [652].

S [123]. Saddlepoint [998]. Samad [706]. Samarendra [19]. Sample [700, 930, 17, 382, 413, 392, 377, 223, 422, 112, 988, 416, 44, 38, 156, 874, 907, 646, 185, 835, 540, 603, 497, 992, 595, 535, 469, 410]. Samples [218, 654, 941, 701, 713, 574]. Sampling [915, 152, 430, 671, 277, 218, 150, 279, 149, 736, 786, 709, 201, 246, 391, 932, 893, 746, 978, 910, 153, 305, 753, 832, 77, 143, 167, 639, 268, 181, 82, 240, 271, 316, 440, 628, 337, 362, 676, 404, 213, 424, 919, 580, 594, 591, 575, 454, 579, 581, 508, 587, 452, 512, 496, 595, 582, 605]. Satpati [123]. Saturated [827, 366, 356]. Saunders [549, 851, 417, 990, 510, 302]. Scalar [1027]. Scale [146, 619, 444, 132, 801, 607]. Scales [941, 845]. Scaling [9]. Scan [375]. Scattering [755]. Schedule [832]. Scheme [368, 674, 910, 82, 454]. Schemes [428, 456]. Schwarz [141]. Science [20]. Sciences [286, 250]. Scientists [371]. Score [197, 285, 771, 167, 607]. Scores [288, 597, 1018]. Scoring [784]. Scramble [880]. Screening [1007, 267, 355]. Seafood [949]. Search [360, 273, 1025, 369, 13]. Search-Based [1025]. Searches [354]. Seasonal [215, 551]. Seat [791]. Secant [689]. Second [432, 358, 616, 444, 157, 961, 442, 660, 265]. Second-Order [432, 358, 616, 444, 961, 442, 660, 265]. Secondary [47]. Section [972]. Seeking [608]. Segmentation [983]. Segregated [163]. Selected [340, 721]. Selecting [178]. Selection [434, 180, 712, 248, 173, 645, 853, 868, 198, 216, 268, 795, 823, 781, 416, 95, 655, 324, 306, 1018, 548, 472]. Selective [41]. Self [543, 811]. Self-consistency-based [543]. Selling [345]. Semi [269, 1028, 973, 4, 922, 725, 981, 100, 465]. Semi-Latin [725, 981]. Semi-Lévy-Driven [922]. Semi-linear [269]. Semi-Markov [1028, 465]. Semi-parametric [973, 4]. Semi-Qualitative [100]. Semiparametric [520, 62, 69, 58, 1015, 68, 972, 717, 563]. semirecursive [502]. Sensitive [279, 382, 893, 618, 956, 880, 179, 862, 744, 478]. sensitivity [499]. Separating [117]. Separations [803]. seqtest [537]. Sequence [876]. Sequences [818]. Sequential [280, 664, 797, 884, 210, 742, 13, 363, 493, 800, 571, 536, 607, 537]. Series [769, 215, 703, 151, 200, 342, 406, 868, 772, 928, 78, 275, 71, 947, 551, 525, 518, 451, 554, 610, 606]. Server [164]. Service [157]. Set [915, 430, 671, 767, 158, 932, 710, 978, 112, 167, 676, 919, 594, 454, 508, 174, 605]. Set-membership [158]. Sets [96, 257, 13]. Setting [938, 476]. Settings [966]. Setup [448]. Several [397, 613]. Severities [295]. Shape [434, 90, 91, 967, 990, 223, 14, 421, 484]. shape-extended [484]. Shared [368, 850]. Sharp [727]. Sharpe [413]. Shoulders [706]. Shrinkage [967, 975, 1014, 895]. SIC [141]. Sided [128, 943, 607]. Sieve [327]. Sign [112]. Signal [187]. signals [607]. Signed [303]. significance [582].

Silverman [386]. Similar [811]. Similarity [196]. Simple [218, 1003, 75, 686, 988, 699, 424, 579]. Simplex [37]. Simplicial [843]. Simplified [404]. Simulated [961]. Simulation [1001, 332, 711, 654, 951, 230, 633, 224, 704, 944]. Simulator [683]. Simulators [928]. Simultaneous [849, 248, 460, 198, 955, 272, 613, 567]. Simultaneously [1021, 1019, 893]. Single [164, 362]. Single- [362]. Singled [192, 461]. Singular [175]. Sir [741]. Situation [763]. Size [687, 245, 219, 17, 810, 723, 233, 731, 988, 416, 44, 161, 725, 591, 562, 595, 981]. Sizes [241, 442, 640, 846]. Skew [297, 370, 281, 338, 973, 130, 439, 771, 716, 801, 646, 838, 854, 867, 612, 514, 606]. Skew-Normal [297, 973, 771, 801, 646]. Skewed [662, 549]. Skip [277]. Skip-Lot [277]. SkSP [277]. SkSP-2 [277]. Slash [380, 838]. Sliced [1004]. Slope [771]. Small [687, 592, 358, 1021, 1019, 413, 210, 931]. Small-area [592]. Small-Sample [413]. smaller [489]. Smirnov [700, 381]. Smooth [249, 132, 42, 247, 133, 135]. Smoothed [386, 942, 968, 643]. Smoothers [667]. smoothing [502]. Smoothness [91]. Social [314, 641, 856]. Software [817]. Solution [374, 540]. Solutions [993]. Solving [928]. Some [193, 1001, 300, 906, 25, 206, 84, 619, 408, 777, 504, 338, 547, 155, 998, 439, 939, 923, 223, 618, 737, 68, 38, 855, 899, 646, 511, 583]. Song [313]. Sources [740]. Space [818, 1025, 232, 36, 287, 473, 554]. Space-Filling [818, 1025, 287]. Spaces [761]. Spacings [818]. Sparse [847, 1013, 848, 470, 607]. Spatial [113, 171, 936, 761, 375, 924]. Spatial-temporal [113]. Spearman [958]. Special [921, 411, 250, 815, 170]. Speciated [113]. Species [944]. Specific [1026]. Specified [390, 568]. Spectral [708, 235]. Spectrum [710]. Speed [311, 236]. Spinelli [137]. Spline [67]. Split [30, 29, 266]. Split-Block [29]. Split-Plot [29, 266, 30]. Spread [227, 1002]. Spring [28]. Square [724, 1017, 30, 745, 80, 463]. Squared [869, 550]. Squares [274, 627, 59]. SRGD [259]. Srivastava [347]. Stable [315, 638, 963, 379, 296, 191, 484, 486]. Stage [107, 178, 282, 728, 611, 201, 478, 496, 571, 961, 143]. standard [763, 738]. Standardization [712]. Standardized [352]. Standing [706]. star [459, 608]. star-contoured [459, 608]. starts [512]. State [94, 366, 78, 473, 554]. Stated [384]. States [119, 473, 554]. Statistic [375, 414, 626, 295, 602, 469]. Statistical [965, 397, 106, 495, 879, 87, 448, 171, 195, 200, 565, 291, 109, 570, 817, 196, 336, 795, 455, 486, 929, 541, 135, 707, 182, 964, 511, 514, 992, 460, 476]. Statistician [788]. Statistics [621, 261, 743, 189, 883, 194, 420, 526, 313, 48, 88, 1023, 33, 642, 188, 68, 131, 959, 462, 501, 971, 497, 371]. Status [407]. STD [115]. Steiner [258]. Step [878, 663, 623, 933, 785, 920, 955, 686, 503, 602, 831]. Step-Stress [785, 920]. Stephens [137]. Stillbirth [924]. Stochastic [883, 75, 183, 951, 857, 664, 610, 875, 882, 718, 748, 854, 483, 20]. Stock [315, 662, 231, 53, 675]. straightforward [604]. Strategies [961, 682]. Strategy [345, 174, 482]. Stratified [709, 153, 639, 38, 581, 508, 478, 560]. Strength [732]. Stress [663, 623, 933, 732, 785, 920, 686]. Stress-Strength [732]. Strong [712, 637]. Structural [215]. Structure [227, 122, 983, 120, 399, 811, 1005, 356, 314, 918, 938, 903, 588]. student [452, 315, 53]. student- [452, 53]. Studentized [118]. Study [204, 431, 303, 786, 40, 219, 141, 897, 405, 835, 996, 521, 472]. Stuttering [343]. Sub [932]. Sub-regression [932]. Subcritical [627]. Subdata [823, 655, 1018]. Subject [228, 168, 296, 323, 264, 452]. subjects [522]. Subquantiles [307]. Subsampling [688, 976]. subsequent [582]. Subset [173, 216]. subsets [608]. Success [284]. Successive [746, 618, 240, 316, 440]. sufficiency [570]. Sum [711, 913, 890, 906]. Summary [1005]. sums [1]. Supersaturated [263, 120, 43, 162, 435, 186, 226, 1005, 242, 354, 369]. Supplementary [187, 607]. Supplemented [266]. Support [936, 998, 381]. Supported [659]. Surface [432, 374, 683]. Surplus [276]. Survey [218, 150, 64, 993, 658, 582]. Surveying [956]. Surveys [382, 156, 874, 907]. Survival [747, 329, 438, 884, 485, 633, 339, 846, 331, 611, 464, 604]. SVR [451]. Swimming [289]. Switches [205]. Switching [408]. Sylvester [916, 810]. Symmetric [991, 203, 963, 7, 808, 115, 938, 467, 796, 199, 486]. Symmetrical [830]. Symmetry [724, 1017, 256, 811, 167, 80, 463]. System [769, 839, 157, 144, 101]. Systematic [437, 391, 271, 337, 580, 587, 512]. Systems [397, 117, 1028, 644, 258, 897, 852, 465, 610].

T. [881]. Tables [724, 1017, 301, 805, 953, 999, 897, 80, 980, 463]. Tabu [369]. Taguchi [733]. Tail [757, 750, 776, 982, 616, 766, 629, 722, 838, 576, 547]. Tailed [887, 401, 982]. Tails [729]. Tampered [933]. target [521]. Teaches [788]. Technique [893, 174, 77, 956, 688]. Technological [925]. temperature [551]. Temporal [708, 714, 113]. Ternary [193]. Terry [309]. Test [449, 743, 441, 735, 979, 726, 9, 194, 700, 736, 767, 623, 313, 187, 127, 933, 775, 853, 377, 414, 247, 112, 167, 642, 774, 407, 165, 114, 835, 974, 903, 596, 539, 503, 613, 460, 536, 582, 607, 544]. Testimation [141]. Testing [680, 290, 397, 412, 796, 245, 105, 129, 307, 865, 740, 1021, 1019, 155, 308, 631, 817, 79, 968, 609, 414, 626, 128, 670, 136, 677, 780, 427, 938, 184, 331, 540, 574, 586, 588]. Tests [443, 652, 647, 965, 6, 126, 269, 298, 400, 249, 299, 842, 934, 403, 739, 197, 134, 413, 132, 130, 385, 971, 798, 125, 624, 438, 223, 771, 176, 845, 344, 133, 597, 955, 1006, 988, 135, 304, 556, 543, 992, 41, 550, 607, 537]. Their [1024, 911, 89, 266, 255]. Them [416]. Theophilos [921]. Theorem [406, 318, 199, 234, 93]. Theorems [758]. Theoretic [430]. Theoretical [711, 792]. Theory [1001, 481, 411, 87, 286, 172, 753, 181, 491]. Theory- [181]. Thinking [336]. Third [794, 863, 576]. Third-Order [863, 576]. Three [878, 202, 870, 279, 358, 674, 734, 1007, 75, 437, 810, 445, 282, 442, 731, 640, 955, 169, 461]. Three-Level [358, 734, 1007, 442]. Three-Parameter [674]. three-player [461]. Three-Stage [282]. Three-Step [878, 955]. Three-way [731]. Threshold [408, 42]. Thresholding [912, 620]. Tied [441]. Tilted [877]. Time [627, 94, 307, 138, 747, 703, 151, 200, 951, 342, 406, 36, 59, 327, 119, 398, 868, 772, 928, 784, 78, 686, 633, 71, 954, 947, 425, 538, 551, 525, 465, 557, 518, 451, 554, 574, 595, 586, 516, 606]. Time-Dependent [951, 557]. Time-homogeneous [94]. Time-Lagged [954]. Time-Series [200, 342, 551, 525, 451, 606]. Time-Series-Valued [928]. Time-Transformation [327]. Time-Varying [59, 138]. Time-Weighted [425]. Times [396, 884, 713]. Tool [949]. Tools [625, 547]. Topp [705]. Torus [799]. Total [647, 286]. Totals [802]. Touch [425]. Traceability [949]. tracing [591]. Tracking [326]. Trading [174]. Training [727]. Trait [973]. Transfer [644]. Transformation [327, 1006, 421]. transformed [592]. Transient [50]. Transition [15, 342, 233, 852, 182]. Transplant [961]. Transversal [115]. Transversals [260]. Trapezoidal [39]. Treatment [193, 987, 207, 334, 209, 103, 427, 946, 555].

Treatments [767, 445, 266]. Trees [977, 95]. Trend [916, 293, 810, 748]. Trend-Free [916]. Trends [375]. Trials [318, 319, 25, 450, 364, 658, 320, 359, 372, 684, 317, 335, 539]. Triangular [906, 39, 21, 536, 537]. Trivariate [335]. Trivial [160]. True [723]. Truncated [666, 430, 514, 452, 607]. Truncation [616, 458]. Trust [885, 891]. Tsallis [737]. Tumor [983]. Turning [1020]. Two [752, 647, 397, 70, 830, 367, 253, 105, 700, 480, 353, 366, 120, 709, 413, 201, 246, 436, 192, 107, 186, 241, 233, 971, 941, 1005, 178, 377, 893, 626, 361, 128, 352, 26, 961, 24, 746, 695, 731, 728, 143, 82, 316, 440, 628, 688, 362, 725, 643, 365, 185, 214, 394, 835, 943, 540, 603, 1002, 611, 568, 528, 575, 503, 602, 497, 478, 609, 992, 496, 571, 500, 456, 469]. Two- [731]. Two-Envelope [695]. Two-Factor [366, 568]. two-layer [1002]. Two-Level [367, 353, 436, 1005, 365, 214, 120, 186, 568, 528]. Two-Occasion [746, 440]. Two-parameter [480, 609]. Two-Phase [893, 628, 362, 575]. Two-Sample [700, 377, 185, 835, 540, 603, 497, 992, 469]. Two-Sided [128, 943]. Two-Stage [107, 178, 728, 201, 143, 611, 478, 496, 571]. two-step [503, 602]. Two-Valued [397]. Type [668, 277, 789, 940, 757, 709, 932, 385, 629, 778, 609, 1023, 713, 978, 760, 405, 427, 221, 80, 477, 504, 575, 602, 508, 605, 822, 564, 705, 887, 726, 623, 696, 723, 713, 910, 686, 994, 1012, 929, 577]. Type-I [910]. Type-II [887, 726, 713, 994, 929, 577].

Uhlenbeck [787]. Unbiased [849, 279, 1013, 692, 587, 482]. Uncertain [740]. Uncertainty [390, 212]. unconstrained [518]. Understanding [786]. Undesirable [218]. Undirected [793]. Unequal [434, 668, 368, 178, 77, 992, 532]. Unified [833, 825, 674, 281, 338, 885, 891, 593]. Uniform [65, 79, 225, 863]. Uniformly [264]. Unimodal [1020, 199, 717]. Union [160, 540]. union-intersection [540]. Unit [344, 994, 165]. Units [213]. Univariate [290, 215, 599, 890, 966, 984, 306]. university [476]. Unknown [1008]. Unobserved [850]. Unordering [753]. unrelated [478]. unreliable [516]. Updated [684]. Upon [818]. Use [791, 48, 632, 436, 130, 182, 324]. useful [600]. User [390]. User-Specified [390]. Using [12, 652, 847, 833, 825, 711, 851, 269, 789, 970, 649, 969, 307, 173, 674, 313, 777, 901, 281, 342, 246, 1025, 231, 779, 64, 447, 288, 778, 798, 681, 893, 1026, 314, 354, 995, 641, 953, 667, 875, 355, 13, 978, 398, 238, 143, 639, 691, 885, 891, 38, 156, 271, 628, 688, 733, 760, 782, 770, 880, 899, 907, 944, 972, 676, 469, 135, 862, 1008, 959, 395, 850, 643, 651, 846, 635, 744, 594, 215, 40, 579, 508, 587, 451, 571, 144, 463, 167, 595, 188, 455, 82, 516, 564, 472]. Utility [96, 229]. Utility-Based [229].

V [123, 687]. Vacation [157]. Validation [995, 488]. Validity [952]. Valuation [229]. Value [393, 6, 423, 286, 678, 871, 857, 952, 704, 311, 304, 476]. Valued [397, 303, 63, 232, 928, 606]. Values [1024, 55, 649, 723, 605]. VARFIMA [191]. variability [601]. Variable [301, 789, 712, 370, 933, 241, 893, 198, 782, 880, 899, 179, 217, 405, 862, 506, 587]. Variables [776, 709, 971, 952, 81, 143, 688, 897, 243, 272, 170, 214, 744, 421, 575, 536, 584]. Variance [849, 743, 419, 690, 969, 40, 843, 201, 889, 781, 38, 899, 404, 213, 424, 908, 488, 499, 992, 517, 571, 460, 482, 469]. variance-balanced [517]. variance-based [499]. Variances [290]. Variate [963, 175, 975]. Variates [890]. Variation [769, 129, 351, 950, 750, 616, 413]. Variations [619]. Varied [454]. Variety [25]. Various [777]. Varying [270, 59, 138].

Vector [642, 381, 802]. Vectors [652, 955]. Versatile [957]. Version [882]. Versus [315, 735, 1010, 209]. vertically [535]. Via [935, 872, 912, 504, 570, 928, 620, 683]. Viewpoint [261]. Vision [100]. Volatility [183]. Volume [701].

Waiting [516]. walk [462]. Walks [951]. Waller [32]. Warner [482]. Wavelet [66, 222, 402, 912, 200, 109, 975, 620]. Wavelet-Based [109]. Way [767, 62, 357, 44, 731, 999]. Weak [637]. Webpage [677]. Weibull [378, 545, 426, 596, 623, 446, 636, 629, 785, 882, 625, 586, 577, 994, 1012, 131, 929, 566, 304, 989]. Weibull-Type [629]. Weighing [367, 28]. Weighted [430, 11, 426, 281, 59, 850, 846, 425, 535]. Weights [390]. Well [894]. Well- [894]. which [160]. White [974]. width [535]. Wildfires [396]. Wildlife [925]. Wind [311]. Windfalls [51]. Wintner [199]. Wise [254, 641]. Within [829, 243, 483]. Without [431, 946, 500]. Work [741, 57, 237, 10, 137, 45, 657, 706, 34, 46]. Workload [59]. World [347, 231, 53]. Worst [278]. Wrapped [238, 1]. Wrong [334].

XII [674, 636].

Ya’acov [69]. Yantram [457]. Young [103]. Yu [671].

Zag [187]. Zero [666, 652, 251, 703, 119, 275, 1014, 704, 964, 553]. Zero-Coupon [704]. Zero-Difference [251]. Zero-Inflated [703, 275, 1014, 964]. Zero-Truncated [666]. Zig [187]. Zig-Zag [187].

References

Coelho:2007:WGD

[1] Carlos A. Coelho. The wrapped Gamma distribution and wrapped sums and linear combinations of independent Gamma and Laplace distributions. Journal of Statistical Theory and Practice, 1(1):1–29, March 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                              //link.springer.com/article/10.
                                              1080/15598608.2007.10411821.

Swift:2007:ANP

[2] Randall J. Swift. Analytic nonstationary processes. Journal of Statistical Theory and Practice, 1 (1):31–38, March 2007. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                              //link.springer.com/article/10.
                                              1080/15598608.2007.10411822.

Gani:2007:DBD

[3] J. Gani and R. J. Swift. Death and birth–death and immigration processes with catastrophes. Journal of Statistical Theory and Practice, 1(1):39–48, March 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                              //link.springer.com/article/10.
                                              1080/15598608.2007.10411823.

Kedem:2007:SPC

[4] Benjamin Kedem and Shihua Wen. Semi-parametric cluster detection. Journal of Statistical Theory and Practice, 1(1):49–72, March 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                              //link.springer.com/article/10.
                                              1080/15598608.2007.10411824.

Rao:2007:ERE

[5] M. M. Rao. Exploring ramifications of the equation E ( Y — X ) = X. Journal of Statistical Theory and Practice, 1(1):73–88, March 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2007.10411825.

Best:2007:CFT

[6] D. J. Best, J. C. W. Rayner, and O. Thas. Comparison of five tests of fit for the extreme value distribution. Journal of Statistical Theory and Practice, 1(1):89–99, March 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2007.10411826.

Katsaounis:2007:GEN

[7] T. I. Katsaounis, C. A. Dingus, and A. M. Dean. On the geometric equivalence and non-equivalence of symmetric factorial designs. Journal of Statistical Theory and Practice, 1(1):101–115, March 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2007.10411827.

Govindarajulu:2007:MFF

[8] Usha S. Govindarajulu, Mark E. Glickman, and Ralph B. D’agostino. Modeling frailty as a function of observed covariates. Journal of Statistical Theory and Practice, 1(1): 117–135, March 2007. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2007.10411828.

Chiu:2007:BRT

[9] Chris Chiu, Ivy Li, and Peggy Carr. Book review: Test equating, scaling, and linking: Methods and practice. Journal of Statistical Theory and Practice, 1(1):137–139, March 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2007.10411829.

Gupta:2007:LWC

[10] Sat Gupta. Life and work of C. R. Rao. Journal of Statistical Theory and Practice, 1(1):141–146, March 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2007.10411830.

Beran:2007:WAL

[11] Jan Beran and Yuanhua Feng. Weighted averages and local polynomial estimation for fractional linear ARCH processes. Journal of Statistical Theory and Practice, 1 (2):149–166, June 2007. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2007.10411831.

Agarwal:2007:LPA

[12] Manju Agarwal, Kanwar Sen, and Bidisha Borkakaty. Lattice path approach for busy period density of GI^{b} /G/1 queues using C_{2} Coxian distributions. Journal of Statistical Theory and Practice, 1(2):167–198, June 2007. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2007.10411832.

Palvolgyi:2007:RSS

[13] Dömötör Pálvölgyi. Revisiting sequential search using question-sets with bounded intersections. Journal of Statistical Theory and Practice, 1(2):199–204, June 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2007.10411833.

Singh:2007:IEI

[14] Housila P. Singh and Sheetal Pandit. Improved estimation of inverse Gaussian shape parameter and measure of dispersion with prior information. Journal of Statistical Theory and Practice, 1(2):205–226, June 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2007.10411834.

Bewernick:2007:RBD

[15] Robert L. Bewernick, Jeremy D. Dewar, Eunice Gray, Nancy Y. Rodriguez, and Randall J. Swift. On the representation of birth–death processes with polynomial transition rates. Journal of Statistical Theory and Practice, 1(2):227–231, June 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2007.10411835.

Ghosh:2007:BAQ

[16] Sujit K. Ghosh and Pabak Mukhopadhyay. Bayesian analysis of Quality Adjusted Lifetime (QAL) data. Journal of Statistical Theory and Practice, 1(2):233–251, June 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2007.10411836.

Govindarajulu:2007:SSR

[17] Z. Govindarajulu. Sample size reestimation: Nonparametric approach. Journal of Statistical Theory and Practice, 1(2):253–264, June 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2007.10411837.

Lipovetsky:2007:PCP

[18] Stan Lipovetsky. Priority and choice probability estimation by ranking, rating and combined data. Journal of Statistical Theory and Practice, 1(2):265–278, June 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2007.10411838.

Sen:2007:SNR

[19] Pranab Kumar Sen. Samarendra Nath Roy: Reminiscence at the centennial. Journal of Statistical Theory and Practice, 1(2):279–283, June 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2007.10411839.

Swift:2007:BRS

[20] Randall Swift. Book review: Stochastic Processes in Science, Engineering and Finance by Frank Beichelt — Chapman & Hall /CRC (2006). Journal of Statistical Theory and Practice, 1(2):285–287, June 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2007.10411840.

Kageyama:2007:NEA

[21] Sanpei Kageyama. On non-existence of affine α resolvable triangular designs. Journal of Statistical Theory and Practice, 1(3–4):291–298, September 2007. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2007.10411841.

Carlini:2007:HBO

[22] Enrico Carlini and Giovanni Pistone. Hilbert bases for orthogonal arrays. Journal of Statistical Theory and Practice, 1(3–4):299–309, September 2007. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2007.10411842.

Liu:2007:DHE

[23] Qing Liu, Angela M. Dean, and Greg M. Allenby. Design for hyperparameter estimation in linear models. Journal of Statistical Theory and Practice, 1(3–4):311–328, September 2007. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2007.10411843.

Notari:2007:TPR

[24] Roberto Notari, Eva Riccomagno, and Maria-Piera Rogantin. Two polynomial representations of experimental design.

Journal of Statistical Theory and Practice, 1(3–4):329–346, September 2007. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2007.10411844.

Calinski:2007:SRC

[25] Tadeusz Caliński. On some results of C. Radhakrishna Rao applicable to the analysis of multienvironment variety trials. Journal of Statistical Theory and Practice, 1 (3–4):347–356, September 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2007.10411845.

Morgan:2007:OET

[26] J. P. Morgan and Bo Jin. Optimal experimentation in two blocks. Journal of Statistical Theory and Practice, 1 (3–4):357–375, September 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2007.10411846.

Ferreira:2007:CAB

[27] S. S. Ferreira, D. Ferreira, and J. T. Mexia. Cross additivity in balanced cross nesting models. Journal of Statistical Theory and Practice, 1(3– 4):377–392, September 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2007.10411847.

Katulska:2007:COS

[28] Krystyna Katulska and Katarzyna Przybył. On certain D-optimal spring balance weighing designs. Journal of Statistical Theory and Practice, 1 (3–4):393–404, September 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2007.10411848.

Mejza:2007:RIS

[29] I. Mejza and K. Ambrozy. Resolvable incomplete split-plot × splitblock designs. Journal of Statistical Theory and Practice, 1(3–4): 405–416, September 2007. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2007.10411849.

Kuriki:2007:SLD

[30] Shinji Kuriki and Kiyoaki Nakajima. Square lattice designs in incomplete split-plot designs. Journal of Statistical Theory and Practice, 1(3– 4):417–426, September 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2007.10411850.

Lu:2007:GOP

[31] Shujie Lu, Eiji Taniguchi, Yoshifumi Hyodo, and Masahide Kuwada. GA-optimal partially balanced fractional 2^{m_{1}+m_{2}} factorial designs of resolutions R({10, 01} ∪ Ω ∗ |Ω) with 2 ≤ m_{1}, m_{2} ≤ 4. Journal of Statistical Theory and Practice, 1(3– 4):427–464, September 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2007.10411851.

Aggarwal:2007:ODO

[32] M. L. Aggarwal, Poonam Singh, and L. Y. Chan. Optimal designs in (q − 1) orthogonal blocks for Darroch and Waller’s quadratic mixture model in q components. Journal of Statistical Theory and Practice, 1(3– 4):465–477, September 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2007.10411852.

Morgenthaler:2007:ES

[33] S. Morgenthaler and P. R. Parthasarathy. Euler on statistics. Journal of Statistical Theory and Practice, 1(3– 4):479–487, September 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2007.10411853.

Rao:2007:LWR

[34] C. R. Rao. Life and work of Ronald Aylmer Fisher. Journal of Statistical Theory and Practice, 1(3– 4):489–499, September 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2007.10411854.

Ryan:2007:MED

[35] Thomas P. Ryan and J. P. Morgan. Modern experimental design. Journal of Statistical Theory and Practice, 1 (3–4):501–506, September 2007. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2007.10411855.

Lee:2008:RAD

[36] Hyeyoung Lee and Sujit K. Ghosh. A reparametrization approach for dynamic space–time models. Journal of Statistical Theory and Practice, 2(1):1–14, March 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411856.

Kim:2008:NDN

[37] S. Kim and J. Switkes. A note on the distribution of the number of simplex iterations to optimality. Journal of Statistical Theory and Practice, 2(1):15–19, March 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411857.

Singh:2008:SFE

[38] Housila P. Singh and Gajendra K. Vishwakarma. Some families of estimators of variance of stratified random sample mean using auxiliary information. Journal of Statistical Theory and Practice, 2(1):21–43, March 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411858.

Goncalves:2008:TTD

[39] Luzia Gonçalves and Maria Antónia Amaral-Turkman. Triangular and trapezoidal distributions: Applications in the genome analysis. Journal of Statistical Theory and Practice, 2(1):45–54, March 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2008.10411859.

Ghosh:2008:BSV

[40] Himadri Ghosh, M. A. Iquebal, and Prajneshu. A bootstrap study of variance estimation under heteroscedasticity using genetic algorithm. Journal of Statistical Theory and Practice, 2(1):55–69, March 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2008.10411860.

Pereira:2008:AST

[41] D. G. Pereira and J. T. Mexia. Application of selective F tests in joint regression analysis. Journal of Statistical Theory and Practice, 2(1):71–81, March 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2008.10411861.

Nur:2008:EES

[42] D. Nur, G. M. Nair, and N. D. Yatawara. Efficient estimation in smooth threshold autoregressive(1) models. Journal of Statistical Theory and Practice, 2(1):83–94, March 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2008.10411862.

Gupta:2008:SDA

[43] V. K. Gupta, Rajender Parsad, Lal Mohan Bhar, and Basudev Kole. Supersaturated designs for asymmetrical factorial experiments. Journal of Statistical Theory and Practice, 2(1):95–108, March 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411863.

Simeckova:2008:SSO

[44] Marie Simecková and Dieter Rasch. Sample size for the one-way layout with one fixed factor for ordered categorical data. Journal of Statistical Theory and Practice, 2(1):109–123, March 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411864.

Notz:2008:BRL

[45] William Notz. A brief reflection on the life and work of Jack Carl Kiefer. Journal of Statistical Theory and Practice, 2(1):125–130, March 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411865.

Rao:2008:LWR

[46] C. R. Rao. Life and work of Ronald Aylmer Fisher. Journal of Statistical Theory and Practice, 2 (1):131–141, March 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411866.

Barnes:2008:IDN

[47] Belinda Barnes and Niels G. Becker. The impact of dispersion in the number of secondary infections on the probability of an epidemic. Journal of Statistical Theory and Practice, 2(2):145–158, June 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411867;   http:
//link.springer.com/content/pdf/
10.1080/15598608.2008.10411867.
pdf.

Ewens:2008:USG

[48] Warren J. Ewens. On the use of statistics in genomics and bioinformatics. Journal of Statistical Theory and Practice, 2(2):159–172, June 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411868.

Suchard:2008:EPP

[49] Marc A. Suchard, Kenneth Lange, and Janet S. Sinsheimer. Efficiency of protein production from mRNA. Journal of Statistical Theory and Practice, 2(2):173–182, June 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411869.

Economou:2008:AAT

[50] Antonis Economou and Demetris Fakinos. Alternative approaches for the transient analysis of Markov chains with catastrophes. Journal of Statistical Theory and Practice, 2(2):183–197, June 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411870.

Hunter:2008:GRC

[51] B. Hunter, A. C. Krinik, C. Nguyen, J. M. Switkes, and H. F. Von Bremen. Gambler’s ruin with catastrophes and windfalls. Journal of Statistical Theory and Practice, 2(2):199–219, June 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411871.

Heyde:2008:CNM

[52] C. C. Heyde and K. All. A cautionary note on model choice and the Kullback–Leibler information. Journal of Statistical Theory and Practice, 2(2):221–232, June 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411872.

Platen:2008:EES

[53] Eckhard Platen and Renata Rendek. Empirical evidence on Student-t log — returns of diversified world stock indices. Journal of Statistical Theory and Practice, 2(2):233–251, June 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411873.

Stoyanov:2008:PMA

[54] S. V. Stoyanov, S. T. Rachev, and F. J. Fabozzi. Probability metrics with applications in finance. Journal of Statistical Theory and Practice, 2(2):253–277, June 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411874.

Andersen:2008:PCF

[55] Lars Nørvang Andersen and Søren Asmussen. Parallel computing, failure recovery, and extreme values. Journal of Statistical Theory and Practice, 2(2):279–292, June 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2008.10411875.

Penrose:2008:MNA

[56] Mathew D. Penrose and Andrew R. Wade. Multivariate normal approximation in geometric probability. Journal of Statistical Theory and Practice, 2(2):293–326, June 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2008.10411876.

Gani:2008:LWC

[57] Joe Gani. Life and work of Christopher Charles Heyde: In memoriam. Journal of Statistical Theory and Practice, 2(2):327–329, June 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2008.10411877.

Kedem:2008:ISM

[58] Benjamin Kedem and Guanhua Lu. Introduction to semiparametric methods. Journal of Statistical Theory and Practice, 2(3):333–337, September 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2008.10411878.

Li:2008:EWR

[59] Ta-Hsin Li. On exponentially weighted recursive least squares for estimating time-varying parameters and its application to computer workload forecasting. Journal of Statistical Theory and Practice, 2(3): 339–354, September 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411879.

Fokianos:2008:NMC

[60] Konstantinos Fokianos and Jing Qin. A note on Monte Carlo maximization by the density ratio model. Journal of Statistical Theory and Practice, 2 (3):355–367, September 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411880.

Zhang:2008:CML

[61] Biao Zhang. Constrained maximum likelihood estimation under logistic regression models based on case-control data. Journal of Statistical Theory and Practice, 2(3): 369–383, September 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411881.

Gagnon:2008:ESA

[62] Richard Gagnon, Benjamin Kedem, and Ying Qi. On the efficiency of a semiparametric approach to the one-way layout. Journal of Statistical Theory and Practice, 2(3): 385–406, September 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2008.10411882.

Greenshtein:2008:EMH

[63] Eitan Greenshtein, Junyong Park, and Ya’acov Ritov. Estimating the mean of high valued observations in high dimensions. Journal of Statistical Theory and Practice, 2(3): 407–418, September 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2008.10411883.

Johnson:2008:EDF

[64] Alicia A. Johnson, F. Jay Breidt, and Jean D. Opsomer. Estimating distribution functions from survey data using nonparametric regression. Journal of Statistical Theory and Practice, 2(3):419–431, September 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2008.10411884.

Feuerverger:2008:OUD

[65] Andrey Feuerverger, Peter T. Kim, and Jiayang Sun. On optimal uniform deconvolution. Journal of Statistical Theory and Practice, 2(3): 433–451, September 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2008.10411885.

Chaubey:2008:WBE

[66] Yogendra P. Chaubey, Hassan Doosti, and B. L. S. Prakasa Rao. Wavelet based estimation of the derivatives of a density for a negatively associated process. Journal of Statistical Theory and Practice, 2(3): 453–463, September 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411886.

Cho:2008:AOP

[67] Jaerin Cho and Boris Levit. Asymptotic optimality of periodic spline interpolation in non-parametric regression. Journal of Statistical Theory and Practice, 2(3):465–474, September 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411887.

Schick:2008:SDS

[68] Anton Schick and Wolfgang Wefelmeyer. Some developments in semiparametric statistics. Journal of Statistical Theory and Practice, 2 (3):475–491, September 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411888.

Kedem:2008:EDS

[69] Benjamin Kedem. From the early days of the semiparametric field: a conversation with ya’acov ritov. Journal of Statistical Theory and Practice, 2(3):493–496, September 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411889.

Arts:2008:TNP

[70] Gerda R. J. Arts and Frank P. A. Coolen. Two nonparametric predictive control charts. Journal of Statistical Theory and Practice, 2 (4):499–512, December 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2008.10411890.

Shangodoyin:2008:ABE

[71] D. K. Shangodoyin and R. Arnab. Analytical Bayes estimator and distribution for outlier infested time series data. Journal of Statistical Theory and Practice, 2(4):513– 521, December 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2008.10411891.

Ghosh:2008:BAC

[72] Sujit K. Ghosh and Nader Ebrahimi. Bayesian analysis of change-point hazard rate problem. Journal of Statistical Theory and Practice, 2(4): 523–533, December 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2008.10411892.

Wang:2008:MAN

[73] Jiantian Wang. Martingale approach to nonparametric estimation of lifetime distribution under different censoring mechanisms. Journal of Statistical Theory and Practice, 2(4): 535–549, December 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411893.

Ahmed:2008:CNB

[74] R. Ahmed and M. Akhtar. Construction of neighbor balanced block designs. Journal of Statistical Theory and Practice, 2(4):551–558, December 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic).

URL http://link.springer.com/
article/10.1080/15598608.2008.10411894.

Gani:2008:SAI

[75] J. Gani and R. J. Swift. A simple approach to the integrals under three stochastic processes. Journal of Statistical Theory and Practice, 2 (4):559–568, December 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411895.

Melas:2008:EBD

[76] Viatcheslav B. Melas and Yury M. Staroselskiy. D-efficient Bayesian designs for a class of nonlinear models. Journal of Statistical Theory and Practice, 2(4):569–587, December 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411896.

Saha:2008:RRT

[77] Amitava Saha. A randomized response technique for quantitative data under unequal probability sampling. Journal of Statistical Theory and Practice, 2(4):589–596, December 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411897.

Saidane:2008:INM

[78] Mohamed Saidane and Christian Lavergne. Improved nonlinear multivariate financial time series prediction with mixed-state latent factor models. Journal of Statistical Theory and Practice, 2(4):597–632, December 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411898.

Liang:2008:EBT

[79] TaChen Liang. Empirical Bayes testing for uniform distributions with random censoring. Journal of Statistical Theory and Practice, 2(4): 633–649, December 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411899.

Yamamoto:2008:CNT

[80] Kouji Yamamoto, Hiroyuki Kurakami, Toshiya Iwashita, and Sadao Tomizawa. Contaminated normal type symmetry model and decomposition of symmetry for square contingency tables. Journal of Statistical Theory and Practice, 2(4): 651–661, December 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411900.

Rao:2008:RPM

[81] P. V. Rao, Haihong Li, and Jeffrey Roth. Recursive path models when both predictor and response variables are categorical. Journal of Statistical Theory and Practice, 2 (4):663–676, December 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411901.

Singh:2008:GFE

[82] Housila P. Singh and Sunil Kumar. A general family of estimators of finite population ratio, product and mean using two phase sampling scheme in the presence of non-response. Journal of Statistical Theory and Practice, 2(4):677–692, December 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411902.

DAmico:2008:CRE

[83] Guglielmo D’Amico. A convergence result in the estimation of Markov chains with application to compound options. Journal of Statistical Theory and Practice, 2(4):693–705, December 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411903.

Dass:2008:SCA

[84] Bal Kishan Dass, Poonam Garg, and Mauro Zannetti. Some combinatorial aspects of m-repeated burst error detecting codes. Journal of Statistical Theory and Practice, 2 (4):707–711, December 2008. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2008.10411904.

Gupta:2008:MRH

[85] V. K. Gupta, A. Dhandapani, R. Parsad, and Warren F. Kuhfeld. Monograph review: Hadamard matrices. Journal of Statistical Theory and Practice, 2(4):713– 714, December 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2008.10411905.

Ryan:2008:LE

[86] Thomas P. Ryan. Letter to the Editor. Journal of Statistical Theory and Practice, 2(4):715–716, December 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2008.10411906.

Coolen-Schrijner:2009:IST

[87] Pauline Coolen-Schrijner, Frank P. A. Coolen, Matthias C. M. Troffaes, and Thomas Augustin. Imprecision in statistical theory and practice. Journal of Statistical Theory and Practice, 3(1):1–9, March 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2009.10411907.

Hampel:2009:NPS

[88] Frank Hampel. Nonadditive probabilities in statistics. Journal of Statistical Theory and Practice, 3 (1):11–23, March 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2009.10411908.

Kozine:2009:BDT

[89] I. Kozine and V. Krymsky. Bounded densities and their derivatives: Extension to other domains. Journal of Statistical Theory and Practice, 3(1):25–38, March 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411909.

Bose:2009:ISC

[90] Sudip Bose. On the imposition of shape constraints in a robust Bayesian analysis. Journal of Statistical Theory and Practice, 3(1):39–55, March 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411910.

Bose:2009:SCS

[91] Sudip Bose. On smoothness constraints with shape constraints in a robust Bayesian analysis. Journal of Statistical Theory and Practice, 3(1):57–67, March 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411911.

Montgomery:2009:BPB

[92] Victoria J. Montgomery, Frank P. A. Coolen, and Andy D. M. Hart. Bayesian probability boxes in risk assessment. Journal of Statistical Theory and Practice, 3(1):69–83, March 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411912.

Pelessoni:2009:BTB

[93] Renato Pelessoni and Paolo Vicig. Bayes’ Theorem bounds for convex lower previsions. Journal of Statistical Theory and Practice, 3 (1):85–101, March 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411913.

Crossman:2009:THB

[94] Richard J. Crossman, Pauline Coolen-Schrijner, and Frank P. A. Coolen. Time-homogeneous birth-death processes with probability intervals and absorbing state. Journal of Statistical Theory and Practice, 3(1): 103–118, March 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411914.

Strobl:2009:ASE

[95] Carolin Strobl and Thomas Augustin. Adaptive selection of extra cutpointstowards reconciling robustness and interpretability in classification trees. Journal of Statistical Theory and Practice, 3(1):119–135, March 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411915.

Farrow:2009:APD

[96] M. Farrow and M. Goldstein. Almost-Pareto decision sets in imprecise utility hierarchies. Journal of Statistical Theory and Practice, 3(1): 137–155, March 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411916.

Danielson:2009:PAE

[97] Mats Danielson, Love Ekenberg, and Ari Riabacke. A prescriptive approach to elicitation of decision data. Journal of Statistical Theory and Practice, 3(1):157–168, March 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411917.

Smithson:2009:PPJ

[98] Michael Smithson and Carl Segale. Partition priming in judgments of imprecise probabilities. Journal of Statistical Theory and Practice, 3 (1):169–181, March 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411918.

Bickis:2009:ILN

[99] Mikelis Bickis. The imprecise logit– normal model and its application to estimating hazard functions. Journal of Statistical Theory and Practice, 3(1):183–195, March 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411919;   http:
//link.springer.com/content/pdf/
10.1080/15598608.2009.10411919.
pdf.

deCampos:2009:SQP

[100] Cassio P. de Campos, Lei Zhang, Yan Tong, and Qiang Ji. Semiqualitative probabilistic networks in computer vision problems. Journal of Statistical Theory and Practice, 3(1):197–210, March 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2009.10411920.

Wilson:2009:IDM

[101] Alyson G. Wilson, Aparna V. Huzurbazar, and Kari Sentz. The imprecise Dirichlet model for multilevel system reliability. Journal of Statistical Theory and Practice, 3 (1):211–223, March 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2009.10411921.

Fuchs:2009:PBC

[102] Martin Fuchs and Arnold Neumaier. Potential based clouds in robust design optimization. Journal of Statistical Theory and Practice, 3(1): 225–238, March 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2009.10411922.

Stoye:2009:PIR

[103] Jörg Stoye. Partial identification and robust treatment choice: an application to Young offenders. Journal of Statistical Theory and Practice, 3(1):239–254, March 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2009.10411923.

Walter:2009:IPD

[104] Gero Walter and Thomas Augustin. Imprecision and prior-data conflict in generalized Bayesian inference. Journal of Statistical Theory and Practice, 3(1):255–271, March 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411924.

Coolen-Schrijner:2009:NPP

[105] Pauline Coolen-Schrijner, Tahani A. Maturi, and Frank P. A. Coolen. Nonparametric predictive precedence testing for two groups. Journal of Statistical Theory and Practice, 3 (1):273–287, March 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411925.

Aughenbaugh:2009:RBD

[106] J. M. Aughenbaugh and J. W. Herrmann. Reliability-based decision making: a comparison of statistical approaches. Journal of Statistical Theory and Practice, 3(1): 289–303, March 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411926.

Kinney:2009:ITS

[107] S. K. Kinney and J. P. Reiter. Inferences for two-stage multiple imputation for nonresponse. Journal of Statistical Theory and Practice, 3(2):307–318, June 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411927.

Alraqad:2009:OEP

[108] T. A. Alraqad and M. S. Shrikhande. An overview of embedding problems of quasi-residual designs. Journal of Statistical Theory and Practice, 3(2):319–347, June 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411928.

Howlader:2009:WBN

[109] T. Howlader and Y. P. Chaubey. Wavelet-based noise reduction by joint statistical modeling of cDNA microarray images. Journal of Statistical Theory and Practice, 3(2): 349–370, June 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411929.

Tremmel:2009:NLE

[110] L. T. Tremmel. A new look at evaluating MTD designs in cancer research. Journal of Statistical Theory and Practice, 3(2):371–379, June 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411930.

Berardi:2009:RBE

[111] Luigia Berardi, Bal Kishan Dass, and Rashmi Verma. On 2-repeated burst error detecting codes. Journal of Statistical Theory and Practice, 3(2):381–391, June 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411931.

Samawi:2009:OBR

[112] Hani M. Samawi and Mavis Pararai. An optimal bivariate ranked set sample design for the matched pairs sign test. Journal of Statistical Theory and Practice, 3(2):393–406, June 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411932.

Choi:2009:MST

[113] Jungsoon Choi, Brian J. Reich, Montserrat Fuentes, and Jerry M. Davis. Multivariate spatial-temporal modeling and prediction of speciated fine particles. Journal of Statistical Theory and Practice, 3 (2):407–418, June 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411933.

Wiens:2009:APN

[114] Douglas P. Wiens. Asymptotic properties of a Neyman–Pearson test for model discrimination, with an application to experimental design. Journal of Statistical Theory and Practice, 3(2):419–427, June 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411934.

Suetake:2009:AGS

[115] C. Suetake. Automorphism groups of a symmetric transversal design STD_{2} [12;6]. Journal of Statistical Theory and Practice, 3(2):429–443, June 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2009.10411935.

Anand:2009:BAI

[116] Suraj P. Anand and Sujit K. Ghosh. A Bayesian approach for investigating the risk of QT prolongation. Journal of Statistical Theory and Practice, 3(2):445–454, June 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2009.10411936.

Balazs:2009:IMC

[117] Patkós Balázs, Tichler Krisztián, and Wiener Gábor. Inclusionwise minimal completely separating systems. Journal of Statistical Theory and Practice, 3(2):455–462, June 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2009.10411937.

Wulff:2009:EML

[118] Shaun S. Wulff. Evaluation of the mixed linear model with orthogonalized and Studentized residuals. Journal of Statistical Theory and Practice, 3(2):463–476, June 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2009.10411938.

Mishra:2009:BRM

[119] Kaushal K. Mishra and Sujit K. Ghosh. Bayesian regression models for the quality adjusted lifetime data with zero time duration health states. Journal of Statistical Theory and Practice, 3(2):477–487, June 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411939.

Georgiou:2009:OTL

[120] Stelios D. Georgiou, Danel Draguljić, and Angela M. Dean. An overview of two-level supersaturated designs with cyclic structure. Journal of Statistical Theory and Practice, 3(2):489–504, June 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411940.

deJesus:2009:BOO

[121] Vera de Jesus, Mexia Tiago João, and Paulo Canas Rodrigues. Binary operations on orthogonal models, application to prime basis factorials and fractional replicates. Journal of Statistical Theory and Practice, 3(2):505–521, June 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411941.

Carvalho:2009:EMC

[122] Francisco P. Carvalho, Mexia T. João, and Manuela M. Oliveira. Estimation in models with commutative orthogonal block structure. Journal of Statistical Theory and Practice, 3(2):523–533, June 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411942.

Singh:2009:MRD

[123] Murari Singh. Monograph review: “αdesigns” by Rajender Parsad, V. K. Gupta, P. K. Batra, S. K. Satpati and Pabitra Biswas, IASRI Publication, New Delhi, India, 2007. Journal of Statistical Theory and Practice, 3(2):535–536, June 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2009.10411943.

Rayner:2009:IMG

[124] J. C. W. Rayner, O. Thas, and D. J. Best. Introduction to modern goodness of fit methods. Journal of Statistical Theory and Practice, 3(3):537–541, September 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2009.10411944.

Lockhart:2009:ECT

[125] Richard A. Lockhart, Federico O’Reilly, and Michael Stephens. Exact Conditional Tests and Approximate Bootstrap Tests for the von Mises Distribution. Journal of Statistical Theory and Practice, 3(3):543– 554, September 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2009.10411945.

Best:2009:ATF

[126] D. J. Best, J. C. W. Rayner, and O. Thas. Anscombe’s tests of fit for the negative binomial distribution. Journal of Statistical Theory and Practice, 3(3):555–565, September 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411946.

Gulati:2009:NGF

[127] Sneh Gulati and Samuel Shapiro. A new goodness of fit test for the logistic distribution. Journal of Statistical Theory and Practice, 3(3): 567–576, September 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411947.

Meintanis:2009:EMP

[128] Simos G. Meintanis and George Iliopoulos. The empirical moment process in testing for the generalized twosided power distribution. Journal of Statistical Theory and Practice, 3 (3):577–586, September 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411948.

Dette:2009:TCC

[129] Holger Dette and Gabriele Wieczorek. Testing for a constant coefficient of variation in nonparametric regression. Journal of Statistical Theory and Practice, 3(3):587–612, September 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411949.

Ketterer:2009:UIC

[130] F. Ketterer, B. Klar, and N. Henze. The use of isotones for comparing tests of normality against skew normal distributions. Journal of Statistical Theory and Practice, 3(3): 613–626, September 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411950.

Stephens:2009:APC

[131] Michael A. Stephens. Asymptotic power of CvM statistics for exponentiality against Weibull and gamma alternatives. Journal of Statistical Theory and Practice, 3(3):627– 643, September 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411951.

Janic:2009:DDS

[132] A. Janic and T. Ledwina. Datadriven smooth tests for a locationscale family revisited. Journal of Statistical Theory and Practice, 3 (3):645–664, September 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411952.

Rayner:2009:GST

[133] J. C. W. Rayner, D. J. Best, and O. Thas. Generalised smooth tests of goodness of fit. Journal of Statistical Theory and Practice, 3(3): 665–679, September 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411953.

Hart:2009:FBL

[134] Jeffrey D. Hart. Frequentist-Bayes lack-of-fit tests based on Laplace approximations. Journal of Statistical Theory and Practice, 3(3):681– 704, September 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411954.

Thas:2009:ISA

[135] O. Thas, J. C. W. Rayner, D. J. Best, and B. De Boeck. Informative statistical analyses using smooth goodness of fit tests. Journal of Statistical Theory and Practice, 3 (3):705–733, September 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411955.

Rayner:2009:PPT

[136] J. C. W. Rayner and A. M. Carolan. Partially parametric testing. Journal of Statistical Theory and Practice, 3 (3):735–750, September 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411956.

Lockhart:2009:LWM

[137] Richard A. Lockhart and John J. Spinelli. The life and work of Michael A. Stephens: a conversation with Richard A. Lockhart and John J. Spinelli. Journal of Statistical Theory and Practice, 3(3): 751–762, September 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2009.10411957.

Dupuy:2009:REC

[138] Jean-François Dupuy. On the random effects Cox model with timevarying regression parameter. Journal of Statistical Theory and Practice, 3(4):763–776, December 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2009.10411958.

Feng:2009:FLP

[139] Yuanhua Feng and Jan Beran. Filtered log-periodogram regression of long memory processes. Journal of Statistical Theory and Practice, 3 (4):777–793, December 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2009.10411959.

Kumar:2009:CDD

[140] C. Satheesh Kumar. A class of discrete distributions of order k. Journal of Statistical Theory and Practice, 3(4):795–803, December 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2009.10411960.

Kanbur:2009:NET

[141] S. M. Kanbur, A. Nanthakumar, and C. Ngeow. On the near equivalence of the testimation and Schwarz information criterion (SIC) to study Cepheid period–luminosity relation. Journal of Statistical Theory and Practice, 3 (4):805–815, December 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411961.

Hoskins:2009:ODI

[142] Dean S. Hoskins, Charles J. Colbourn, and Douglas C. Montgomery. D-optimal designs with interaction coverage. Journal of Statistical Theory and Practice, 3(4): 817–830, December 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411962.

Sahoo:2009:GME

[143] L. N. Sahoo, Nirupama Mahapatra, and S. C. Senapati. A generalized method of estimation for two-stage sampling using two auxiliary variables. Journal of Statistical Theory and Practice, 3(4):831–839, December 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411963.

Roychowdhury:2009:RCS

[144] Soma Roychowdhury and Debasis Bhattacharya. Rearrangement of components in a system using component importance measures. Journal of Statistical Theory and Practice, 3 (4):841–854, December 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411964.

Neuhaus:2009:RJM

[145] A. Neuhaus, T. Augustin, C. Heumann, and D. Daumer. A review on joint models in biometrical research. Journal of Statistical Theory and Practice, 3(4):855–868, December 2009. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2009.10411965.

Chander:2009:ICS

[146] Vankim Chander and Housila P. Singh. Improved classes of scale parameter towards an interval of exponential distribution. Journal of Statistical Theory and Practice, 3(4): 869–877, December 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2009.10411966.

Beder:2009:BHR

[147] Jay H. Beder and Wiebke S. Diestelkamp. Box-Hunter resolution in nonregular fractional factorial designs. Journal of Statistical Theory and Practice, 3(4):879–889, December 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2009.10411967.

Kirk:2009:CMA

[148] Roger E. Kirk and B. Nebiyou Bekele. Cell means analyses for ANCOVA designs. Journal of Statistical Theory and Practice, 3(4): 891–906, December 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411968.

Coolen:2009:NPM

[149] Frank P. A. Coolen and Mohamed A. Elsaeiti. Nonparametric predictive methods for acceptance sampling. Journal of Statistical Theory and Practice, 3(4):907–921, December 2009. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2009.10411969.

Cadigan:2010:KRE

[150] N. G. Cadigan and J. Chen. Kernel regression estimators for nonparametric model calibration in survey sampling. Journal of Statistical Theory and Practice, 4(1):1–25, March 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10411970.

Ghosh:2010:NTS

[151] Himadri Ghosh, Ranjit Kumar Paul, and Prajneshu. Nonlinear time series modeling and forecasting for periodic and ARCH effects. Journal of Statistical Theory and Practice, 4(1):27–44, March 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10411971.

Akman:2010:ICB

[152] Olcay Akman. Information complexity based modeling in the presence of length-biased sampling. Journal of Statistical Theory and Practice, 4(1):45–55, March 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2010.10411972.

Rao:2010:NOA

[153] T. J. Rao. Near optimum allocations in stratified sampling. Journal of Statistical Theory and Practice, 4(1):57–69, March 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2010.10411973.

Liang:2010:EBE

[154] TaChen Liang. Empirical Bayes estimation with random censoring. Journal of Statistical Theory and Practice, 4(1):71–83, March 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2010.10411974.

Jiang:2010:OPT

[155] Dongming Jiang and S. Sivaganesan. On objective priors for testing hypotheses about some Poisson models. Journal of Statistical Theory and Practice, 4(1):85–110, March 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2010.10411975.

Singh:2010:EMR

[156] Housila P. Singh and Namrata Karpe. Estimation of mean, ratio and product using auxiliary information in the presence of measurement errors in sample surveys. Journal of Statistical Theory and Practice, 4(1): 111–136, March 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10411976.

Maraghi:2010:BAV

[157] Farzana A. Maraghi, Kailash C. Madan, and Ken Darby-Dowman. Batch arrival vacation queue with second optional service and random system breakdowns. Journal of Statistical Theory and Practice, 4(1): 137–153, March 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10411977.

Jaulin:2010:PSM

[158] Luc Jaulin. Probabilistic setmembership approach for robust regression. Journal of Statistical Theory and Practice, 4(1): 155–167, March 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10411978.

Neath:2010:LME

[159] Andrew A. Neath and Joseph E. Cavanaugh. Linear model evaluation based on estimation of model bias. Journal of Statistical Theory and Practice, 4(1):169–179, March 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10411979.

Fontana:2010:LFF

[160] Roberto Fontana and Giovanni Pistone. 2-level fractional factorial designs which are the union of non trivial regular designs. Journal of Statistical Theory and Practice, 4 (1):181–202, March 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10411980.

Tahir:2010:CPD

[161] M. H. Tahir, Ijaz Iqbal, Munir Akhtar, and Javid Shabbir. Cyclic polygonal designs with block size 3 and λ = 1 for joint distance α = 6 to 16. Journal of Statistical Theory and Practice, 4(2):203–220, June 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10411981.

Gupta:2010:CAC

[162] V. K. Gupta, Poonam Singh, Basudev Kole, and Rajender Parsad. Computer aided construction of efficient multilevel supersaturated designs. Journal of Statistical Theory and Practice, 4(2):221–231, June 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10411982.

Ferreira:2010:NSM

[163] S. S. Ferreira, D. Ferreira, C. Nunes, and J. T. Mexia. Nesting segregated mixed models. Journal of Statistical Theory and Practice, 4 (2):233–242, June 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10411983.

Kumar:2010:SSR

[164] B. Krishna Kumar, R. Rukmani, V. Thangaraj, and Udo R. Krieger. A single server retrial queue with Bernoulli feedback and collisions. Journal of Statistical Theory and Practice, 4(2):243–260, June 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10411984.

Wang:2010:MLE

[165] Dazhe Wang, Sujit K. Ghosh, and Sastry G. Pantula. Maximum likelihood estimation and unit root test for first order random coefficient AutoRegressive models. Journal of Statistical Theory and Practice, 4 (2):261–278, June 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10411985.

Dey:2010:BEP

[166] Sanku Dey and Sudhansu S. Maiti. Bayesian estimation of the parameter of Maxwell distribution under different loss functions. Journal of Statistical Theory and Practice, 4 (2):279–287, June 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10411986.

Samawi:2010:DFR

[167] Hani M. Samawi, Robert Vogel, and Christopher K. Senkowski. On distribution–free runs test for symmetry using extreme ranked set sampling with an application involving base deficit score. Journal of Statistical Theory and Practice, 4 (2):289–301, June 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2010.10411987.

Guerreiro:2010:MOP

[168] G. R. Guerreiro, J. T. Mexia, and M. F. Miguens. A model for open populations subject to periodical Re-classifications. Journal of Statistical Theory and Practice, 4(2):303–321, June 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2010.10411988.

Shankar:2010:CCA

[169] G. Shankar and P. K. Chourasia. Conditional corrective action plan for three attribute classes. Journal of Statistical Theory and Practice, 4(2):323–336, June 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2010.10411989.

Wang:2010:EHP

[170] Dabuxilatu Wang. Estimation of the hidden parameters for the distribution of special fuzzy random variables. Journal of Statistical Theory and Practice, 4(2):337–344, June 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2010.10411990.

Dimitriou-Fakalou:2010:SIS

[171] Chrysoula Dimitriou-Fakalou. Statistical inference for spatial autolinear processes. Journal of Statistical Theory and Practice, 4(3): 345–365, September 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10411991.

Nair:2010:FME

[172] Maya T. Nair and Saleem.E Abdul. Finite mixture of exponential model and its applications to renewal and reliability theory. Journal of Statistical Theory and Practice, 4 (3):367–373, September 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10411992.

Dorugade:2010:SSL

[173] A. V. Dorugade and D. N. Kashid. Subset selection in linear regression using generalized ridge estimator. Journal of Statistical Theory and Practice, 4(3):375–389, September 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10411993.

Puspaningrum:2010:FOP

[174] Heni Puspaningrum, Yan-Xia Lin, and Chandra M. Gulati. Finding the optimal pre-set boundaries for pairs trading strategy based on cointegration technique. Journal of Statistical Theory and Practice, 4 (3):391–419, September 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10411994.

Diaz-Garcia:2010:DNS

[175] José A. Díaz-García and Ramón Gutiérrez-Jáimez. Doubly noncentral singular matrix variate beta distributions. Journal of Statistical Theory and Practice, 4(3):421– 431, September 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2010.10411995.

Osman:2010:NBM

[176] Muhtarjan Osman and Sujit K. Ghosh. Novel Bayesian methods for noninferiority tests based on relative risk and odds ratio for dichotomous data. Journal of Statistical Theory and Practice, 4(3):433–452, September 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2010.10411996.

Nasiri:2010:BAG

[177] Parviz Nasiri and Hassan Pazira. Bayesian approach on the generalized exponential distribution in the presence of outliers. Journal of Statistical Theory and Practice, 4 (3):453–475, September 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2010.10411997.

Liu:2010:BTS

[178] Chengan Liu and Zakkula Govindarajulu. A balanced two-stage procedure for selecting the best of normal populations with unequal means. Journal of Statistical Theory and Practice, 4(3):477–494, September 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10411998.

Sousa:2010:REM

[179] Rita Sousa, Javid Shabbir, Pedro Corte Real, and Sat Gupta. Ratio estimation of the mean of a sensitive variable in the presence of auxiliary information. Journal of Statistical Theory and Practice, 4(3):495–507, September 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10411999.

Baker:2010:NPC

[180] Rebecca M. Baker and Frank P. A. Coolen. Nonparametric predictive category selection for multinomial data. Journal of Statistical Theory and Practice, 4(3):509–526, September 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10412000.

Shabbir:2010:AST

[181] Javid Shabbir. Advances in sampling theory- ratio method of estimation. Journal of Statistical Theory and Practice, 4(3):527, September 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10412001;   http:
//link.springer.com/content/pdf/
10.1080/15598608.2010.10412001.
pdf.

Warr:2010:ESF

[182] Richard L. Warr and Aparna V. Huzurbazar. Expanding the statistical flowgraph model framework to use any transition distribution. Journal of Statistical Theory and Practice, 4(4):529–539, December 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2010.10412002.

Gong:2010:SVM

[183] H. Gong, A. Thavaneswaran, and J. Singh. Stochastic volatility models with application in option pricing. Journal of Statistical Theory and Practice, 4(4):541–557, December 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2010.10412003.

Voss:2010:TCE

[184] D. T. Voss. Testing is confidence estimation: Partition multiple testing. Journal of Statistical Theory and Practice, 4(4):559–569, December 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2010.10412004.

Zhang:2010:TSB

[185] Jin-Ting Zhang, Xuehua Liang, and Shengning Xiao. On the twosample Behrens–Fisher problem for functional data. Journal of Statistical Theory and Practice, 4(4): 571–587, December 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10412005.

Kole:2010:TLS

[186] Basudev Kole, Jyoti Gangwani, V. K. Gupta, and Rajender Parsad. Twolevel supersaturated designs: a review. Journal of Statistical Theory and Practice, 4(4):589–608, December 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10412006.

Griffiths:2010:PCS

[187] David Griffiths, Martin Bunder, Chandra Gulati, and Takeo Onizawa. The probability of an out of control signal from Nelson’s supplementary zig-zag test. Journal of Statistical Theory and Practice, 4(4): 609–615, December 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10412007.

Sankaran:2010:NEB

[188] P. G. Sankaran, Ashis SenGupta, and V. N. Sreeja. Nonparametric estimation of bivariate distribution using concomitants of order statistics. Journal of Statistical Theory and Practice, 4(4):617–630, December 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10412008.

Bruggemann:2010:MPP

[189] Rainer Bruggemann and Ganapati P. Patil. Multicriteria prioritization and partial order in environmental and ecological statistics. Journal of Statistical Theory and Practice, 4 (4):631–641, December 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10412009.

Lee:2010:CPP

[190] Tze-San Lee. Change-point problems: Bibliography and review. Journal of Statistical Theory and Practice, 4(4):643–662, December 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10412010.

Pai:2010:FBE

[191] Jeffrey S. Pai and Nalini Ravishanker. Fast Bayesian estimation for VARFIMA processes with stable errors. Journal of Statistical Theory and Practice, 4(4):663– 677, December 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10412011.

Johanis:2010:EMT

[192] Michal Johanis and Jan Rychtár. An enhanced model of a two player singled out game. Journal of Statistical Theory and Practice, 4(4): 679–697, December 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10412012.

Aggarwal:2010:SCB

[193] Manohar Aggarwal and Mithilesh Kumar Jha. Some constructions of balanced ternary residual treatment effects designs. Journal of Statistical Theory and Practice, 4(4): 699–710, December 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10412013.

Coelho:2010:NED

[194] Carlos A. Coelho, Barry C. Arnold, and Filipe J. Marques. Near-exact distributions for certain likelihood ratio test statistics. Journal of Statistical Theory and Practice, 4(4): 711–725, December 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10412014.

Ghosh:2010:SIN

[195] Sujit K. Ghosh and Lovely Goyal. Statistical inference for non-linear models involving ordinary differential equations. Journal of Statistical Theory and Practice, 4(4): 727–742, December 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10412015.

Mulekar:2010:ESM

[196] Madhuri S. Mulekar and Takeshi Fukasawa. Estimation of statistical measures of income similarity. Journal of Statistical Theory and Practice, 4 (4):743–755, December 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10412016.

Gupta:2010:LST

[197] A. K. Gupta and Yanhong Wu. Local score tests in mixture exponential family with fixed mean. Journal of Statistical Theory and Practice, 4(4):757–771, December 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2010.10412017.

Rao:2010:EBC

[198] C. R. Rao, Y. Wu, and X. Shi. An Mestimation-based criterion for simultaneous change point analysis and variable selection in a regression problem. Journal of Statistical Theory and Practice, 4(4):773–801, December 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2010.10412018.

Mudholkar:2010:PCS

[199] Govind S. Mudholkar and Hongyue Wang. Product-convolution of Rsymmetric unimodal distributions: an analogue of Wintner’s Theorem. Journal of Statistical Theory and Practice, 4(4):803–811, December 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2010.10412019.

Ghosh:2010:WFD

[200] Himadri Ghosh, R. K. Paul, and Prajneshu. Wavelet frequency domain approach for statistical modeling of rainfall time-series data. Journal of Statistical Theory and Practice, 4(4):813–825, December 2010. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10412020.

Haziza:2010:VET

[201] David Haziza and J. N. K. Rao. Variance estimation in two-stage cluster sampling under imputation for missing data. Journal of Statistical Theory and Practice, 4(4):827– 844, December 2010. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2010.10412021.

Agboto:2011:CTA

[202] Vincent Agboto, William Li, and Christopher Nachtsheim. A comparison of three approaches for constructing robust experimental designs. Journal of Statistical Theory and Practice, 5(1):1–11, March 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10412046.

Bailey:2011:SFD

[203] R. A. Bailey. Symmetric factorial designs in blocks. Journal of Statistical Theory and Practice, 5 (1):13–24, March 2011. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10412047.

Burgess:2011:CDC

[204] Leonie Burgess, Deborah J. Street, and Nada Wasi. Comparing designs for choice experiments: a case study. Journal of Statistical Theory and Practice, 5(1):25–46, March 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10412048.

Challenor:2011:DCE

[205] Peter Challenor. Designing a computer experiment that involves switches. Journal of Statistical Theory and Practice, 5(1):47–57, March 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10412049.

Chai:2011:SOR

[206] Feng-Shun Chai and Ching-Shui Cheng. Some optimal row–column designs. Journal of Statistical Theory and Practice, 5(1):59–67, March 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10412050.

Kunert:2011:ODT

[207] J. Kunert and S. Mersmann. Aoptimal designs for treatment — control comparisons in microarray experiments with random block effects. Journal of Statistical Theory and Practice, 5(1):69–79, March 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10412051.

Moon:2011:AGM

[208] Hyejung Moon, Angela Dean, and Thomas Santner. Algorithms for generating maximin Latin hypercube and orthogonal designs. Journal of Statistical Theory and Practice, 5(1):81–98, March 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10412052.

Morgan:2011:OTV

[209] J. P. Morgan and Xiaowei Wang. Eoptimality in treatment versus control experiments. Journal of Statistical Theory and Practice, 5(1): 99–107, March 2011. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10412053.

Nguyen:2011:NSC

[210] Nam-Ky Nguyen and Dennis K. J. Lin. A note on small composite designs for sequential experimentation. Journal of Statistical Theory and Practice, 5(1):109–117, March 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10412054.

Ranjan:2011:FED

[211] Pritam Ranjan, Wilson Lu, Derek Bingham, Shane Reese, Brian J. Williams, Chuan-Chih Chou, Forrest Doss, Michael Grosskopf, and James Paul Holloway. Follow-up experimental designs for computer models and physical processes. Journal of Statistical Theory and Practice, 5(1):119–136, March 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10412055.

Woods:2011:COD

[212] D. C. Woods and S. M. Lewis. Continuous optimal designs for generalized linear models under model uncertainty. Journal of Statistical Theory and Practice, 5(1):137–145, March 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10412056.

Wright:2011:VAU

[213] James H. Wright and John Stufken. Variance approximation under balanced sampling plans excluding adjacent units. Journal of Statistical Theory and Practice, 5(1): 147–160, March 2011. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10412057.

Zhang:2011:OTL

[214] Runchu Zhang, Peng Li, and Jialin Wei. Optimal two-level regular designs with multi block variables. Journal of Statistical Theory and Practice, 5(1):161–178, March 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10412058.

Birrell:2011:SAA

[215] Carole L. Birrell, David G. Steel, and Yan-Xia Lin. Seasonal adjustment of an aggregate series using univariate and multivariate basic structural models. Journal of Statistical Theory and Practice, 5(2):179–205, June 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10412023.

Sakate:2011:SSP

[216] D. M. Sakate, D. N. Kashid, and D. T. Shirke. Subset selection in Poisson regression. Journal of Statistical Theory and Practice, 5(2):207–219, June 2011. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10412024.

Suaray:2011:VBK

[217] Kagba Suaray. Variable bandwidth kernel density estimation for censored data. Journal of Statistical Theory and Practice, 5(2):221–229, June 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10412025.

Bisht:2011:SRS

[218] Ila Pant Bisht and Raj Kishor Bisht. Simple random sampling equivalent survey design for the elimination of undesirable samples. Journal of Statistical Theory and Practice, 5 (2):231–239, June 2011. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10412026.

Gosky:2011:CSB

[219] R. M. Gosky and S. K. Ghosh. A comparative study of Bayes estimators of closed population size from capture–recapture data. Journal of Statistical Theory and Practice, 5(2):241–260, June 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10412027.

Hosseini:2011:CCM

[220] Reza Hosseini, Nhu Le, and Jim Zidek. A characterization of categorical Markov chains. Journal of Statistical Theory and Practice, 5 (2):261–284, June 2011. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10412028.

Upadhyaya:2011:IRP

[221] Lakshmi N. Upadhyaya, Housila P. Singh, S. Chatterjee, and Rohini Yadav. Improved ratio and product exponential type estimators. Journal of Statistical Theory and Practice, 5(2):285–302, June 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10412029.

Chesneau:2011:AWE

[222] Christophe Chesneau. Adaptive wavelet estimator for a function and its derivatives in an indirect convolution model. Journal of Statistical Theory and Practice, 5(2):303–326, June 2011. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10412030.

Muralidharan:2011:SLS

[223] K. Muralidharan and Kuang-Chao Chang. Some large sample tests based on conditional distribution for the shape parameter in power law process.

Journal of Statistical Theory and Practice, 5(2):327–334, June 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10412031.

Si:2011:CPS

[224] Yajuan Si and Jerome P. Reiter. A comparison of posterior simulation and inference by combining rules for multiple imputation. Journal of Statistical Theory and Practice, 5(2):335–347, June 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10412032.

Tahata:2011:MUA

[225] Kouji Tahata and Sadao Tomizawa. Measure for uniform association based on concordant and discordant pairs for cross-classifications. Journal of Statistical Theory and Practice, 5 (2):349–356, June 2011. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10412033.

Koukouvinos:2011:ACO

[226] C. Koukouvinos, K. Mylona, and D. E. Simos. An algorithmic construction of E ( s^{2} )-optimal supersaturated designs. Journal of Statistical Theory and Practice, 5(2):357–367, June 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10412034.

Broom:2011:EDG

[227] M. Broom, J. Rychtár, and B. T. Stadler. Evolutionary dynamics on graphs- the effect of graph structure and initial placement on mutant spread. Journal of Statistical Theory and Practice, 5(3):369– 381, September 2011. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10412035.

Gani:2011:PBP

[228] J. Gani and R. J. Swift. A PCR birth process subject to an enzyme death process. Journal of Statistical Theory and Practice, 5(3): 383–396, September 2011. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10412036.

German:2011:OUB

[229] David German. Overview of utilitybased valuation. Journal of Statistical Theory and Practice, 5(3): 397–411, September 2011. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10412037.

Huber:2011:SRI

[230] Mark Huber. Simulation reductions for the Ising model. Journal of Statistical Theory and Practice, 5 (3):413–424, September 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10412038.

Ignatieva:2011:UDC

[231] Katja Ignatieva, Eckhard Platen, and Renata Rendek. Using dynamic copulae for modeling dependency in currency denominations of a diversified world stock index. Journal of Statistical Theory and Practice, 5 (3):425–452, September 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10412039.

Kakihara:2011:RND

[232] Y. Kakihara. Radon–Nikodým derivatives of Hilbert space valued measures. Journal of Statistical Theory and Practice, 5(3):453–473, September 2011. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10412040.

Krinik:2011:MPC

[233] A. C. Krinik and K. Shun. Markov processes with constant transition rates of size one & two. Journal of Statistical Theory and Practice, 5 (3):475–495, September 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10412041.

ONeill:2011:GPF

[234] Michael D. O’Neill. A green proof of Fatou’s Theorem. Journal of Statistical Theory and Practice, 5 (3):497–513, September 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10412042.

Swift:2011:SRC

[235] Randall J. Swift. A spectral representation of a class of nonstationary processes. Journal of Statistical Theory and Practice, 5(3): 515–523, September 2011. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10412043.

Switkes:2011:AOP

[236] Jennifer Switkes. Averaging our perceptions of highway speed. Journal of Statistical Theory and Practice, 5 (3):525–535, September 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10412044.

Green:2011:LWM

[237] Michael L. Green, Alan C. Krinik, and Randall J. Swift. The life and work of M. M. Rao. Journal of Statistical Theory and Practice, 5 (3):537–545, September 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10412045.

Ravindran:2011:BAC

[238] Palanikumar Ravindran and Sujit K. Ghosh. Bayesian analysis of circular data using wrapped distributions. Journal of Statistical Theory and Practice, 5(4):547–561, December 2011. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10483731.

Hayter:2011:RIM

[239] A. J. Hayter. Recursive integration methodologies with applications to the evaluation of multivariate normal probabilities. Journal of Statistical Theory and Practice, 5(4): 563–589, December 2011. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10483732.

Singh:2011:MRE

[240] Housila P. Singh, Sunil Kumar, and Sandeep Bhougal. Multivariate ratio estimation in presence of nonresponse in successive sampling. Journal of Statistical Theory and Practice, 5(4):591–611, December 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10483733.

Kowalik:2011:CIE

[241] Danuta Kowalik, Yun-Hee Choi, and G. Y. Zou. Confidence interval estimation for a difference between two dependent intraclass correlation coefficients with variable class sizes. Journal of Statistical Theory and Practice, 5(4):613–625, December 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10483734.

Mandal:2011:CEM

[242] B. N. Mandal, V. K. Gupta, and Rajender Parsad. Construction of efficient mixed-level k-circulant supersaturated designs. Journal of Statistical Theory and Practice, 5 (4):627–648, December 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10483735.

Steel:2011:MAW

[243] David G. Steel and Mark D. Tranmer. Measuring and analyzing the within group homogeneity of multicategory variables. Journal of Statistical Theory and Practice, 5(4): 649–658, December 2011. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10483736.

Jadhav:2011:JRE

[244] N. H. Jadhav and D. N. Kashid. A jackknifed ridge M-estimator for regression model with multicollinearity and outliers. Journal of Statistical Theory and Practice, 5(4): 659–673, December 2011. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10483737.

Bose:2011:WPS

[245] Sudip Bose and Mark Bauder. When population size does not matter: Robust Bayesian testing for categorical populations. Journal of Statistical Theory and Practice, 5(4): 675–682, December 2011. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10483738.

Helu:2011:IOC

[246] Amal Helu and Hani Samawi. On inference of overlapping coefficients in two Lomax populations using different sampling methods. Journal of Statistical Theory and Practice, 5 (4):683–696, December 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10483739.

Pommeret:2011:DDS

[247] D. Pommeret. Data driven smooth test for contaminated data. Journal of Statistical Theory and Practice, 5(4):697–714, December 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10483740.

Curtis:2011:BAM

[248] S. McKay Curtis and Sujit K. Ghosh. A Bayesian approach to multicollinearity and the simultaneous selection and clustering of predictors in linear regression. Journal of Statistical Theory and Practice, 5(4): 715–735, December 2011. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2011.10483741.

DeBoeck:2011:GST

[249] B. De Boeck, O. Thas, J. C. W. Rayner, and D. J. Best. Generalised smooth tests for the generalised Pareto distribution. Journal of Statistical Theory and Practice, 5 (4):737–750, December 2011. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2011.10483742.

Jimbo:2012:SIC

[250] Masakazu Jimbo and Sanpei Kageyama. Special issue on “Combinatorial configurations and information sciences”. Journal of Statistical Theory and Practice, 6(1):1–2, March 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2012.647477;   http:
                                        //link.springer.com/content/pdf/
                                        10.1080/15598608.2012.647477.pdf.

Ding:2012:ZDB

[251] Cunsheng Ding and Yin Tan. Zerodifference balanced functions with applications. Journal of Statistical Theory and Practice, 6(1):3–19, March 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2012.647479.

Tang:2012:OCL

[252] Yu Tang, Charles J. Colbourn, and Jianxing Yin. Optimality and constructions of locating arrays. Journal of Statistical Theory and Practice, 6(1):20–29, March 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2012.647484.

Colbourn:2012:ITR

[253] Charles J. Colbourn and Junling Zhou. Improving two recursive constructions for covering arrays. Journal of Statistical Theory and Practice, 6(1):30–47, March 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.647489.

Momihara:2012:COO

[254] Koji Momihara. Constructions of optical orthogonal codes based on cyclic t-wise quasi-difference matrices. Journal of Statistical Theory and Practice, 6(1):48–68, March 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.647495.

Miyamoto:2012:MIA

[255] N. Miyamoto and S. Shinohara. Mutually M-intersecting K-arcs and their applications. Journal of Statistical Theory and Practice, 6(1): 69–77, March 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.647511.

Angata:2012:MDD

[256] Masayuki Angata and Keisuke Shiromoto. Mutually disjoint 5-designs from Pless symmetry codes. Journal of Statistical Theory and Practice, 6(1):78–87, March 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.647525.

Hiramine:2012:DS

[257] Yutaka Hiramine. On (p^{2}, p, p^{2}, p)- difference sets in Z_{p}^{3}. Journal of Statistical Theory and Practice, 6 (1):88–96, March 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2012.647527.

Munemasa:2012:SQS

[258] Akihiro Munemasa and Masanori Sawa. Steiner quadruple systems with point-regular Abelian automorphism groups. Journal of Statistical Theory and Practice, 6(1):97–128, March 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2012.647536.

Kadowaki:2012:NCM

[259] Satoru Kadowaki and Sanpei Kageyama. New construction methods of affine resolvable SRGD designs. Journal of Statistical Theory and Practice, 6(1):129–138, March 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2012.647551.

Fu:2012:TA

[260] Hung-Lin Fu and Chang-Chun Lee. Transversals in m × n arrays. Journal of Statistical Theory and Practice, 6(1):139–146, March 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2012.647554.

Aoki:2012:DAF

[261] Satoshi Aoki and Akimichi Takemura. Design and analysis of fractional factorial experiments from the viewpoint of computational algebraic statistics.

Journal of Statistical Theory and Practice, 6(1):147–161, March 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.647556.

Cheng:2012:NOR

[262] Ching-Shui Cheng. A note on the Eoptimality of regular line graph designs. Journal of Statistical Theory and Practice, 6(1):162–168, March 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.647558.

Chai:2012:OSD

[263] Feng Shun Chai, Kashinath Chatterjee, Ashish Das, and Chand Midha. Optimal supersaturated designs for s^{m} factorials in N ≢ 0 (mod s) runs. Journal of Statistical Theory and Practice, 6(1):169–177, March 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.647577.

Zhao:2012:UBC

[264] Shi Zhao and Dibyen Majumdar. On uniformly balanced crossover designs efficient under subject dropout. Journal of Statistical Theory and Practice, 6(1):178–189, March 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.647579.

Yumiba:2012:CBS

[265] Hiromu Yumiba, Yoshifumi Hyodo, and Masahide Kuwada. Characterization of balanced second-order designs for 3^{m} factorials. Journal of Statistical Theory and Practice, 6(1):190–203, March 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.647582.

Mejza:2012:MCI

[266] Iwona Mejza, Stanisław Mejza, and Shinji Kuriki. A method of constructing incomplete split-plot designs supplemented by control treatments and their analysis. Journal of Statistical Theory and Practice, 6(1):204–219, March 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.647583.

Kanamori:2012:PDB

[267] Takafumi Kanamori, Hiroaki Uehara, and Masakazu Jimbo. Pooling design and bias correction in DNA library screening. Journal of Statistical Theory and Practice, 6(1): 220–238, March 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.647585.

Sangeetha:2012:SDS

[268] V. Sangeetha and K. K. Suresh. Selection of the double sampling plan through quality regions. Journal of Statistical Theory and Practice, 6 (2):239–250, June 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.673866.

Bhowmik:2012:DTS

[269] Jahar L. Bhowmik and Maxwell L. King. Deriving tests of the semilinear regression model using the density function of a maximal invariant. Journal of Statistical Theory and Practice, 6(2):251–259, June 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.673871.

Hoque:2012:IER

[270] Zahirul Hoque and Shahadut Hossain. Improved estimation in regression with varying penalty. Journal of Statistical Theory and Practice, 6(2):260–273, June 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.673878.

Singh:2012:ECE

[271] Housila P. Singh and Ramkrishna S. Solanki. An efficient class of estimators for the population mean using auxiliary information in systematic sampling. Journal of Statistical Theory and Practice, 6(2):274–285, June 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.673881.

Wagler:2012:BRS

[272] Amy Wagler and Melinda McCann. Bias-reduced simultaneous confidence bands on generalized linear models with restricted predictor variables. Journal of Statistical Theory and Practice, 6(2):286–302, June 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2012.673882.

Gerbner:2012:PSP

[273] Dániel Gerbner and Balázs Keszegh. Path search in the pyramid and in other graphs. Journal of Statistical Theory and Practice, 6(2):303–314, June 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2012.673885.

Araveeporn:2012:LSC

[274] Autcha Araveeporn. The least-squares criteria of the random coefficient dynamic regression model. Journal of Statistical Theory and Practice, 6(2):315–333, June 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2012.673891.

Samani:2012:ZIP

[275] E. Bahrami Samani, M. Ganjali, and Y. Amirian. Zero-inflated power series joint model to analyze count data with missing responses. Journal of Statistical Theory and Practice, 6(2):334–343, June 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2012.673892.

Das:2012:FRJ

[276] Kumer Pial Das and William Ted Mahavier. Further results for the joint distribution of the surplus immediately before and after ruin under force of interest. Journal of Statistical Theory and Practice, 6(2):344–353, June 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.673895.

Balamurali:2012:ODS

[277] S. Balamurali and J. Subramani. Optimal designing of skip-lot sampling plan of type SkSP-2 with double sampling plan as the reference plan. Journal of Statistical Theory and Practice, 6(2):354–362, June 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.673897.

Street:2012:DAL

[278] Deborah J. Street and Stephanie A. Knox. Designing for attribute-level best–worst choice experiments. Journal of Statistical Theory and Practice, 6(2):363–375, June 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.673900.

Chaudhuri:2012:UES

[279] Arijit Chaudhuri. Unbiased estimation of a sensitive proportion in general sampling by three nonrandomized. Journal of Statistical Theory and Practice, 6(2):376–381, June 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.673904.

Elsaeiti:2012:NPA

[280] Mohamed A. Elsaeiti and Frank P. A. Coolen. A nonparametric predictive approach to sequential acceptance problems. Journal of Statistical Theory and Practice, 6(3): 383–401, September 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.695930.

Gupta:2012:EPU

[281] Arjun K. Gupta and Mohammad A. Aziz. Estimation of parameters of the unified skew normal distribution using the method of weighted moments. Journal of Statistical Theory and Practice, 6(3):402–416, September 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.697841.

Mehta:2012:TSO

[282] Samridhi Mehta, B. K. Dass, Javid Shabbir, and Sat Gupta. A three-stage optional randomized response model. Journal of Statistical Theory and Practice, 6(3):417–427, September 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.695558.

Bang:2012:MBI

[283] Heejung Bang and Hongwei Zhao. Median-based incremental cost-effectiveness ratio (ICER). Journal of Statistical Theory and Practice, 6(3):428– 442, September 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.695571.

Wilcox:2012:NRW

[284] Rand R. Wilcox. Nonparametric regression when estimating the probability of success: a comparison of four extant estimators. Journal of Statistical Theory and Practice, 6 (3):443–451, September 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.695639.

Juutilainen:2012:EPS

[285] I. Juutilainen, S. Tamminen, and J. Röning. Exceedance probability score: a novel measure for comparing probabilistic predictions. Journal of Statistical Theory and Practice, 6 (3):452–467, September 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.695663.

DeZeaBermudez:2012:EVT

[286] P. De Zea Bermudez and Zilda Mendes. Extreme value theory in medical sciences: Modeling total high cholesterol levels. Journal of Statistical Theory and Practice, 6(3): 468–491, September 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.695673.

Nguyen:2012:NNO

[287] Nam-Ky Nguyen and Dennis K. J. Lin. A note on near-orthogonal Latin hypercubes with good spacefilling properties. Journal of Statistical Theory and Practice, 6(3): 492–500, September 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2012.695700.

Kiatsupaibul:2012:DRL

[288] S. Kiatsupaibul and A. J. Hayter. Dimensional reduction for latent scores modeling using recursive integration. Journal of Statistical Theory and Practice, 6(3):501–509, September 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2012.695701.

Adam:2012:BEA

[289] M. B. Adam and J. A. Tawn. Bivariate extreme analysis of Olympic swimming data. Journal of Statistical Theory and Practice, 6(3): 510–523, September 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2012.695702.

Allingham:2012:TEV

[290] David Allingham and J. C. W. Rayner. Testing equality of variances for multiple univariate normal populations. Journal of Statistical Theory and Practice, 6(3):524–535, September 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2012.695703.

Guo:2012:PSC

[291] Guangbao Guo. Parallel statistical computing for statistical inference. Journal of Statistical Theory and Practice, 6(3):536–565, September 2012.

CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.695705.

Miladinovic:2012:BQE

[292] Branko Miladinovic and Chris P. Tsokos. Bayesian quantiles of extremes. Journal of Statistical Theory and Practice, 6(3):566–579, September 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.698206.

Hilow:2012:MCL

[293] Hisham Hilow. Minimum cost linear trend free fractional factorial designs. Journal of Statistical Theory and Practice, 6(3):580–589, September 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.698209.

Coelho:2012:BR

[294] Carlos A. Coelho and Abel M. Rodrigues. Book review. Journal of Statistical Theory and Practice, 6 (3):590–595, September 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.695708.

Sendova:2012:OSC

[295] Kristina P. Sendova and Ricardas Zitikis. The order-statistic claim process with dependent claim frequencies and severities. Journal of Statistical Theory and Practice, 6(4): 597–620, December 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.719736.

Guerreiro:2012:SDO

[296] Gracinda Rita Guerreiro, João Tiago Mexia, and Maria de Fátima Miguens. Stable distributions for open populations subject to periodical reclassifications. Journal of Statistical Theory and Practice, 6(4):621– 635, December 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.719741.

Adcock:2012:MES

[297] C. J. Adcock and K. Shutes. On the multivariate extended skew-normal, normal–exponential, and normal– Gamma distributions. Journal of Statistical Theory and Practice, 6 (4):636–664, December 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.719799.

Coolen-Maturi:2012:NPI

[298] Tahani Coolen-Maturi, Pauline Coolen-Schrijner, and Frank P. A. Coolen. Nonparametric predictive inference for binary diagnostic tests. Journal of Statistical Theory and Practice, 6 (4):665–680, December 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.719800.

Elkhafifi:2012:NPI

[299] Faiza F. Elkhafifi and Frank P. A. Coolen. Nonparametric predictive inference for accuracy of ordinal diagnostic tests. Journal of Statistical Theory and Practice, 6(4): 681–697, December 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.719802.

Balakrishnan:2012:EAB

[300] N. Balakrishnan and S. Pal. EM algorithm-based likelihood estimation for some cure rate models. Journal of Statistical Theory and Practice, 6 (4):698–724, December 2012. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.719803.

Bekele:2012:API

[301] B. Nebiyou Bekele, Luis E. Nieto-Barajas, and Mark F. Munsell. Analysis of partially incomplete tables of breast cancer characteristics with an ordinal variable. Journal of Statistical Theory and Practice, 6(4): 725–744, December 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.719805.

Saulo:2012:KBS

[302] Helton Saulo, Jeremias Leão, and Marcelo Bourguignon. The Kumaraswamy Birnbaum–Saunders distribution. Journal of Statistical Theory and Practice, 6(4):745– 759, December 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.719814.

Chesneau:2012:PSF

[303] Christophe Chesneau and Maher Kachour. A parametric study for the first-order signed integer-valued autoregressive process. Journal of Statistical Theory and Practice, 6(4): 760–782, December 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2012.719816.

Thiagarajah:2012:CTH

[304] K. Thiagarajah. Combined tests for the homogeneity of Weibull (or extreme value) populations with censored data. Journal of Statistical Theory and Practice, 6(4):783–792, December 2012. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2012.719817.

Rao:2013:HLE

[305] T. J. Rao. Homogeneous linear estimator and optimality in finite population sampling. Journal of Statistical Theory and Practice, 7(1):1–7, March 2013. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.756263.

Zougab:2013:BAB

[306] Nabil Zougab, Smail Adjabi, and Célestin C. Kokonendji. A Bayesian approach to bandwidth selection in univariate associate kernel estimation. Journal of Statistical Theory and Practice, 7(1):8–23, March 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.756286.

Dewan:2013:TIF

[307] Isha Dewan, P. G. Sankaran, and P. Anisha. On testing independence of failure time and cause of failure using subquantiles. Journal of Statistical Theory and Practice, 7(1):24–32, March 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.756296.

Kim:2013:RRG

[308] Jong-Min Kim and Tae-Young Heo. Randomized response group testing model. Journal of Statistical Theory and Practice, 7(1):33–48, March 2013. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.756324.

Feddag:2013:PML

[309] M.-L. Feddag. Pairwise marginal likelihood for the Bradley–Terry model. Journal of Statistical Theory and Practice, 7(1):49–58, March 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.756336.

Bedbur:2013:CRM

[310] S. Bedbur, J. M. Lennartz, and U. Kamps. Confidence regions in models of ordered data. Journal of Statistical Theory and Practice, 7(1):59–72, March 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.756340.

Steinkohl:2013:EVA

[311] Christina Steinkohl, Richard A. Davis, and Claudia Klüppelberg. Extreme value analysis of multivariate highfrequency wind speed data. Journal of Statistical Theory and Practice, 7(1):73–94, March 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.756346.

Das:2013:IDM

[312] I. Das, S. Mukhopadhyay, and H. Xu. Individualized dosing for multiple ordered groups of patients. Journal of Statistical Theory and Practice, 7(1):95–106, March 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.756348.

Enomoto:2013:ANT

[313] Rie Enomoto, Naoya Okamoto, and Takashi Seo. On the asymptotic normality of test statistics using Song’s kurtosis. Journal of Statistical Theory and Practice, 7(1): 107–119, March 2013. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.756351.

McCormick:2013:PGM

[314] Tyler H. McCormick, Amal Moussa, Johannes Ruf, Thomas A. DiPrete, Andrew Gelman, Julien Teitler, and Tian Zheng. A practical guide to measuring social structure using indirectly observed network data. Journal of Statistical Theory and Practice, 7(1):120–132, March 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.756360.

Alparslan:2013:SPV

[315] Artun Alparslan, Anthony Tessitore, and Nilufer Usmen. Stable Paretian versus Student’s t stock market hypothesis. Journal of Statistical Theory and Practice, 7(1): 133–145, March 2013. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.756321.

Singh:2013:ERP

[316] G. N. Singh and F. Homa. Effective rotation patterns in successive sampling over two occasions. Journal of Statistical Theory and Practice, 7(1):146–155, March 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.755484.

Zhang:2013:LPC

[317] Zhiwei Zhang and Bo Zhang. A likelihood paradigm for clinical trials. Journal of Statistical Theory and Practice, 7(2):157–177, June 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.771545.  See

comments [318, 319, 320].

Bickel:2013:PLE

[318] David R. Bickel. Pseudo-likelihood, explanatory power, and Bayes’s Theorem [comment on “A Likelihood Paradigm for Clinical Trials”]. Journal of Statistical Theory and Practice, 7(2):178–182, June 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.771546.  See

[317].

Blume:2013:LCH

[319] Jeffrey D. Blume. Likelihood and composite hypotheses [comment on “A Likelihood Paradigm for Clinical Trials”]. Journal of Statistical Theory and Practice, 7(2):183–186, June 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.771548.  See

[317].

Spanos:2013:RLE

[320] Aris Spanos. Revisiting the likelihoodist evidential account [comment on “A Likelihood Paradigm for Clinical Trials”]. Journal of Statistical Theory and Practice, 7(2):187–195, June 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.771549.  See

[317].

Zhang:2013:R

[321] Zhiwei Zhang and Bo Zhang. Rejoinder. Journal of Statistical Theory and Practice, 7(2):196–203, June 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.771552.

Cruz-Marcelo:2013:EPW

[322] Alejandro Cruz-Marcelo, Gary L. Rosner, Peter Müller, and Clinton F. Stewart. Effect on prediction when modeling covariates in Bayesian nonparametric models. Journal of Statistical Theory and Practice, 7(2):204–218, June 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.772811.

Guo:2013:BMI

[323] Ying Guo and Roderick J. Little. Bayesian multiple imputation for assay data subject to measurement error. Journal of Statistical Theory and Practice, 7(2):219–232, June 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.772018.

Whitney:2013:UBM

[324] Melissa Whitney, Louise Ryan, and Jens Walkowiak. On the use of Bayesian model averaging for covariate selection in epidemiological modeling. Journal of Statistical Theory and Practice, 7(2):233–247, June 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.772037.

Mitra:2013:PCB

[325] Riten Mitra, Peter Müller, and Yuan Ji. Propriety conditions for the Bayesian autologistic model-inference for histone modifications. Journal of Statistical Theory and Practice, 7(2):248–258, June 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.772838.

Wu:2013:NEC

[326] Colin O. Wu and Xin Tian. Nonparametric estimation of conditional distribution functions and rank-tracking probabilities with longitudinal data. Journal of Statistical Theory and Practice, 7(2):259–284, June 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.771598.

McLain:2013:ESM

[327] Alexander C. McLain and Sujit K. Ghosh. Efficient sieve maximum likelihood estimation of time-transformation models. Journal of Statistical Theory and Practice, 7(2):285–303, June 2013. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.772835.

Chen:2013:GRR

[328] Shuai Chen and Hongwei Zhao. Generalized redistribute-to-the-right algorithm: Application to the analysis of censored cost data. Journal of Statistical Theory and Practice, 7(2):304–323, June 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.771556.

Kim:2013:JML

[329] Sehee Kim, Donglin Zeng, Yi Li, and Donna Spiegelman. Joint modeling of longitudinal and cure–survival data. Journal of Statistical Theory and Practice, 7(2):324–344, June 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.772036.

Sima:2013:OCE

[330] Camelia S. Sima and Mithat Gönen. Optimal cutpoint estimation with censored data. Journal of Statistical Theory and Practice, 7(2): 345–359, June 2013. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.772022.

Zucker:2013:TCC

[331] David M. Zucker, Sarit Agami, and Donna Spiegelman. Testing for a changepoint in the Cox survival regression model. Journal of Statistical Theory and Practice, 7 (2):360–380, June 2013. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.772030.

Bang:2013:BCM

[332] Heejung Bang, Ya-Lin Chiu, Jay S. Kaufman, Mehul D. Patel, Gerardo Heiss, and Kathryn M. Rose. Bias correction methods for misclassified covariates in the Cox model: Comparison of five correction methods by simulation and data analysis. Journal of Statistical Theory and Practice, 7(2):381–400, June 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.772830.

Lauer:2013:CLP

[333] Jeremy A. Lauer and Sander Greenland. Connecting logistic probability models with basic dynamic processes. Journal of Statistical Theory and Practice, 7(2):401–420, June 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.771599.

Lee:2013:CRW

[334] Shih-Yuan Lee and Alex Tsodikov. Correcting the results of the wrong model: Treatment effects under early detection of cancer. Journal of Statistical Theory and Practice, 7 (2):421–441, June 2013. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.772033.

Zhong:2013:FLC

[335] Wei Zhong, Bradley P. Carlin, and Joseph S. Koopmeiners. Flexible link continual reassessment methods for trivariate binary outcome phase I/II trials. Journal of Statistical Theory and Practice, 7(2):442–455, June 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.772815.

Obenchain:2013:AST

[336] R. L. Obenchain and S. S. Young. Advancing statistical thinking in observational health care research. Journal of Statistical Theory and Practice, 7(2):456–469, June 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.772821.

Subramani:2013:FML

[337] J. Subramani. A further modification on linear systematic sampling for finite populations. Journal of Statistical Theory and Practice, 7 (3):471–479, September 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.790231.

Gupta:2013:SPU

[338] Arjun K. Gupta, Mohammad A. Aziz, and Wei Ning. On some properties of the unified skew normal distribution. Journal of Statistical Theory and Practice, 7(3):480–495, September 2013. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.790233.

Wang:2013:RKM

[339] J. Wang and Y. Li. Revisit Kaplan– Meier estimator in estimating QAL survival distributions. Journal of Statistical Theory and Practice, 7 (3):496–504, September 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.790235.

Gangopadhyay:2013:ERC

[340] Aditi Kar Gangopadhyay, Praveen Kulshreshtha, and Monika Verma. Estimation of regression coefficient of a selected population. Journal of Statistical Theory and Practice, 7 (3):505–514, September 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.790237.

Hayter:2013:NPB

[341] A. J. Hayter. A new procedure for the Behrens–Fisher problem that guarantees confidence levels. Journal of Statistical Theory and Practice, 7 (3):515–536, September 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.790239.

Hassan:2013:FPT

[342] M. Y. Hassan and M. Y. El-Bassiouni. Fitting Poisson time-series models using bivariate mixture transition distributions. Journal of Statistical Theory and Practice, 7(3):537– 543, September 2013. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.790241.

Kumar:2013:ISP

[343] C. Satheesh Kumar and D. S. Shibu. On intervened stuttering Poisson distribution and its applications. Journal of Statistical Theory and Practice, 7(3):544–557, September 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.790242.

Puspaningrum:2013:URT

[344] Heni Puspaningrum, Yan-Xia Lin, and Chandra Gulati. Unit root tests for ESTAR models. Journal of Statistical Theory and Practice, 7 (3):558–595, September 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.719812.

Egozcue:2013:OSM

[345] Martín Egozcue, Luis Fuentes García, and Ričardas Zitikis. An optimal strategy for maximizing the expected real-estate selling price: Accept or reject an offer? Journal of Statistical Theory and Practice, 7 (3):596–609, September 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2012.737297.

Gupta:2013:P

[346] Sudhir Gupta and Rahul Mukerjee. Preface. Journal of Statistical Theory and Practice, 7(4):611, December 2013. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.828976;   http:
//link.springer.com/content/pdf/
10.1080/15598608.2013.828976.pdf.

Ghosh:2013:WJS

[347] Subir Ghosh, Sudhir Gupta, and Rahul Mukerjee. The world of Jagdish N. Srivastava. Journal of Statistical Theory and Practice, 7 (4):612–616, December 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.828958.

Katona:2013:EMB

[348] Gyula O. H. Katona and Krisztián Tichler. Existence of a maximum balanced matching in the hypercube. Journal of Statistical Theory and Practice, 7(4):617–629, December 2013. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.781454.

Colbourn:2013:RCA

[349] Charles J. Colbourn. Resolvable covering arrays. Journal of Statistical Theory and Practice, 7(4): 630–649, December 2013. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.781461.

Seberry:2013:NFA

[350] Jennifer Seberry. A new family of amicable Hadamard matrices. Journal of Statistical Theory and Practice, 7(4):650–657, December 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.781469.

Dette:2013:ODR

[351] Holger Dette and Werner G. Müller. Optimal designs for regression models with a constant coefficient of variation. Journal of Statistical Theory and Practice, 7(4):658–673, December 2013. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.781833.

Melas:2013:SMD

[352] Viatcheslav B. Melas. On standardized maximin designs for discrimination between two polynomial models. Journal of Statistical Theory and Practice, 7(4):674–686, December 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.781841.

Dingus:2013:IDE

[353] Cheryl Dingus, Bruce Ankenman, Angela Dean, and Fangfang Sun. Identification of dispersion effects in replicated two-level fractional factorial experiments. Journal of Statistical Theory and Practice, 7(4): 687–702, December 2013. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.781851.

Miller:2013:FMO

[354] Arden Miller and Boxin Tang. Finding MDS–Optimal supersaturated designs using computer searches. Journal of Statistical Theory and Practice, 7(4):703–712, December 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.781888.

Nguyen:2013:CDS

[355] Nam-Ky Nguyen and Stella Stylianou. Constructing definitive screening designs using cyclic generators. Journal of Statistical Theory and Practice, 7(4):713–724, December 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.781891.

Lin:2013:NSF

[356] Dennis K. J. Lin and Lili Xiao. A note on saturated first-order design with foldover structure. Journal of Statistical Theory and Practice, 7 (4):725–731, December 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.782184.

Godolphin:2013:CPW

[357] J. D. Godolphin. On the connectivity problem for m-way designs. Journal of Statistical Theory and Practice, 7(4):732–744, December 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.782193.

Dey:2013:STL

[358] Aloke Dey and Basudev Kole. Small three-level second-order designs with orthogonal blocks. Journal of Statistical Theory and Practice, 7(4): 745–752, December 2013. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.783671.

Sverdlov:2013:RAO

[359] Oleksandr Sverdlov and William F. Rosenberger. On recent advances in optimal allocation designs in clinical trials. Journal of Statistical Theory and Practice, 7(4):753– 773, December 2013. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.783726. See er-

ratum [372].

Angelopoulos:2013:CSD

[360] P. Angelopoulos, K. Chatterjee, and C. Koukouvinos. Construction of search designs from orthogonal arrays. Journal of Statistical Theory and Practice, 7(4):774–782, December 2013. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.783730.

Matsubara:2013:ETP

[361] K. Matsubara and S. Kageyama. The existence of two pairwise additive BIBD(v, 2, 1) for any v. Journal of Statistical Theory and Practice, 7 (4):783–790, December 2013. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.783742.

Sud:2014:CBP

[362] U. C. Sud, Hukum Chandra, and V. K. Gupta. Calibration-based product estimator in single- and two-phase sampling. Journal of Statistical Theory and Practice, 8(1):1–11, March 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.840487.

Roy:2014:EPC

[363] Soma Roy and William I. Notz. Estimating percentiles in computer experiments: a comparison of sequential– adaptive designs and fixed designs. Journal of Statistical Theory and Practice, 8(1):12–29, March 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.840491.

Fedorov:2014:RDT

[364] Valerii V. Fedorov and Tao Liu. Randomized discontinuation trials with binary outcomes. Journal of Statistical Theory and Practice, 8 (1):30–45, March 2014. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.840492.

Wei:2014:BTL

[365] Jialin Wei, Peng Li, and Runchu Zhang. Blocked two-level regular designs with general minimum lower order confounding. Journal of Statistical Theory and Practice, 8(1):46–65, March 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.840517.

Fontana:2014:TFS

[366] Roberto Fontana, Fabio Rapallo, and Maria-Piera Rogantin. Two-factor saturated designs: Cycles, Gini index, and state polytopes. Journal of Statistical Theory and Practice, 8(1):66–82, March 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.840518.

Cheng:2014:OBW

[367] Ching-Shui Cheng. Optimal biased weighing designs and twolevel main-effect plans. Journal of Statistical Theory and Practice, 8 (1):83–99, March 2014. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.840520.

Bose:2014:UPS

[368] Mausumi Bose and Rahul Mukerjee. An unequal probability scheme for improving anonymity in shared key operations. Journal of Statistical Theory and Practice, 8(1):100–112, March 2014. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.840521.

Morales:2014:TSA

[369] Luis B. Morales and Sudhir Gupta. A tabu search algorithm for the construction of χ^{2}(D)-optimal mixed-level kcirculant supersaturated designs. Journal of Statistical Theory and Practice, 8(1):113–125, March 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.840522.

Ghosh:2014:HDB

[370] Subir Ghosh and Debarshi Dey. Heuristically deciding between normal and skew normal distributions for describing the data on a response variable and an explanatory variable. Journal of Statistical Theory and Practice, 8(1):126–137, March 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.823581.

Ghosh:2014:BRS

[371] Subir Ghosh. Book review: Statistics and Probability with Applications for Engineers and Scientists. Journal of Statistical Theory and Practice, 8(1):138–139, March 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2014.846718.

Sverdlov:2014:ERA

[372] Oleksandr Sverdlov and William F. Rosenberger. Erratum to: On recent advances in optimal allocation designs in clinical trials. Journal of Statistical Theory and Practice, 8(1):140, March 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2014.871906;   http:
                                        //link.springer.com/content/pdf/
                                        10.1080/15598608.2014.871906.pdf.

See [359].

Jensen:2014:NIR

[373] D. R. Jensen and D. E. Ramirez. Noncentralities induced in regression diagnostics. Journal of Statistical Theory and Practice, 8(2):141–165, June 2014. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2014.847758.

Diaz-Garcia:2014:ANO

[374] José A. Díaz-García, José E. Rodríguez, and Rogelio Ramos-Quiroga. Asymptotic normality of the optimal solution in response surface methodology. Journal of Statistical Theory and Practice, 8(2):166–175, June 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.847763.

Luo:2014:SSS

[375] Jun Luo. A spatial scan statistic on trends. Journal of Statistical Theory and Practice, 8(2):176–191, June 2014. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.847765.

Laksaci:2014:CHE

[376] Ali Laksaci and Boubaker Mechab. Conditional hazard estimate for functional random fields. Journal of Statistical Theory and Practice, 8 (2):192–220, June 2014. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.847766.

Magel:2014:PNT

[377] Rhonda C. Magel and Ran Fu. Proposed nonparametric test for the mixed two-sample design. Journal of Statistical Theory and Practice, 8(2):221–237, June 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.847768.

Akram:2014:CEW

[378] Muhammad Akram and Aziz Hayat. Comparison of estimators of the Weibull distribution. Journal of Statistical Theory and Practice, 8 (2):238–259, June 2014. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.847771.

Fallahgoul:2014:ANF

[379] Hassan Fallahgoul, S. M. Hashemiparast, Frank J. Fabozzi, and L. Klebanov. Analytical-numeric formulas for the probability density function of multivariate stable and geo-stable distributions. Journal of Statistical Theory and Practice, 8(2):260–282, June 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.780989.

Gui:2014:GSH

[380] Wenhao Gui. A generalization of the slash half normal distribution: Properties and inferences. Journal of Statistical Theory and Practice, 8(2):283–296, June 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.785733.

Utkin:2014:CSV

[381] Lev V. Utkin and Frank P. A. Coolen. Classification with support vector machines and Kolmogorov–Smirnov bounds. Journal of Statistical Theory and Practice, 8(2):297–318, June 2014. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.788985.

Groenitz:2014:AND

[382] Heiko Groenitz. Applying the nonrandomized diagonal model to estimate a sensitive distribution in complex sample surveys. Journal of Statistical Theory and Practice, 8 (2):319–342, June 2014. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.793630.

Cordeiro:2014:MOF

[383] Gauss M. Cordeiro, Artur J. Lemonte, and Edwin M. M. Ortega. The Marshall–Olkin family of distributions: Mathematical properties and new models. Journal of Statistical Theory and Practice, 8(2):343–366, June 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.802659.

Bush:2014:ODS

[384] Stephen Bush. Optimal designs for stated choice experiments generated from fractional factorial designs. Journal of Statistical Theory and Practice, 8(2):367–381, June 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.805451.

Koizumi:2014:MJB

[385] Kazuyuki Koizumi, Masashi Hyodo, and Tatjana Pavlenko. Modified Jarque–Bera type tests for multivariate normality in a high-dimensional framework. Journal of Statistical Theory and Practice, 8(2):382–399, June 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.806232.

Hosseini-Nasab:2014:MSM

[386] S. Mohammad E. Hosseini-Nasab. A modification of Silverman’s method for smoothed functional principal components analysis. Journal of Statistical Theory and Practice, 8 (2):400–413, June 2014. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.811131.

Lucas:2014:BRR

[387] Peter J. F. Lucas. Book review: Risk Assessment and Decision Analysis With Bayesian Networks, by Norman Fenton and Martin Neil, Boca Raton, FL, CRC Press, 2012. ISBN: 978-1-4398-0910-5 (hardcover), 524 pp., $83.95. Journal of Statistical Theory and Practice, 8(2):414–416, June 2014. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2014.847770.

Kadilar:2014:P

[388] Cem Kadilar and Hulya Cingi. Preface. Journal of Statistical Theory and Practice, 8(3):417, September 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2014.906148;   http:
                                        //link.springer.com/content/pdf/
                                        10.1080/15598608.2014.906148.pdf.

Morgan:2014:CED

[389] J. P. Morgan and J. W. Stallings. On the A criterion of experimental design.

Journal of Statistical Theory and Practice, 8(3):418–422, September 2014. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.769922.

Bennett:2014:EOP

[390] Christopher J. Bennett and Ricardas Zitikis. Estimation of optimal portfolio weights under parameter uncertainty and user-specified constraints: a perturbation method. Journal of Statistical Theory and Practice, 8 (3):423–438, September 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.795125.

Khan:2014:NDC

[391] Zaheen Khan, Sat Gupta, and Javid Shabbir. A note on diagonal circular systematic sampling. Journal of Statistical Theory and Practice, 8 (3):439–443, September 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.801805.

Katzoff:2014:SFR

[392] Myron Katzoff, Wen Zhou, Diba Khan, Guanhua Lu, and Benjamin Kedem. Out-of-sample fusion in risk prediction. Journal of Statistical Theory and Practice, 8(3):444–459, September 2014. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.806233. See er-

ratum [410].

Balakrishna:2014:EVA

[393] N. Balakrishna and K. Shiji. Extreme value autoregressive model and its applications. Journal of Statistical Theory and Practice, 8(3): 460–481, September 2014. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.813415.

Zhang:2014:ELC

[394] Dong Zhang and Biao Zhang. Empirical likelihood confidence intervals for the difference of areas under two correlated ROC curves. Journal of Statistical Theory and Practice, 8 (3):482–508, September 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.815592.

Turkman:2014:GAF

[395] K. F. Turkman, M. A. Amaral Turkman, P. Pereira, A. Sá, and J. M. C. Pereira. Generating annual fire risk maps using Bayesian hierarchical models. Journal of Statistical Theory and Practice, 8(3): 509–533, September 2014. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2013.820158.

Gelfand:2014:ERT

[396] Alan E. Gelfand and Joao V. D. Monteiro. Explaining return times for wildfires. Journal of Statistical Theory and Practice, 8(3):534– 545, September 2014. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.821047.

Aoki:2014:STP

[397] Satoshi Aoki and Masami Miyakawa. Statistical testing procedure for the interaction effects of several controllable factors in two-valued input– output systems. Journal of Statistical Theory and Practice, 8(3): 546–557, September 2014. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.838721.

Puspaningrum:2014:MFP

[398] Heni Puspaningrum, Yan-Xia Lin, and Chandra Gulati. Mean first passage time for the ESTAR model using the integral equation approach and its applications. Journal of Statistical Theory and Practice, 8 (4):559–577, December 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.814379.

Khan:2014:CSA

[399] Diba Khan, Myron Katzoff, and Benjamin Kedem. Coherence structure and its application in mortality forecasting. Journal of Statistical Theory and Practice, 8(4):578–590, December 2014. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.816645.

Coolen:2014:NPI

[400] Frank P. A. Coolen and Sulafah Bin Himd. Nonparametric predictive inference for reproducibility of basic nonparametric tests. Journal of Statistical Theory and Practice, 8(4): 591–618, December 2014. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.819792.

Gardiner:2014:FHT

[401] Joseph C. Gardiner, Zhehui Luo, Xiaoqin Tang, and R. V. Ramamoorthi. Fitting heavy-tailed distributions to health care data by parametric and Bayesian methods. Journal of Statistical Theory and Practice, 8(4):619–652, December 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.824823.

Chesneau:2014:PWE

[402] Christophe Chesneau and Fabien Navarro. On a plug-in wavelet estimator for convolutions of densities. Journal of Statistical Theory and Practice, 8(4):653–673, December 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.824824.

Gebert:2014:PBT

[403] Deyse Márcia Pacheco Gebert and Daniel Furtado Ferreira. Parametric bootstrap tests for determining the number of principal components. Journal of Statistical Theory and Practice, 8(4):674–691, December 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.828337.

Tiwari:2014:MVO

[404] Neeraj Tiwari and Akhil Chilwal. On minimum variance optimal controlled sampling: a simplified approach. Journal of Statistical Theory and Practice, 8(4):692–706, December 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.828338.

Sud:2014:CAB

[405] U. C. Sud, Hukum Chandra, and V. K. Gupta. Calibration approachbased regression-type estimator for inverse relationship between study and auxiliary variable. Journal of Statistical Theory and Practice, 8 (4):707–721, December 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.832643.

Khardani:2014:CLT

[406] S. Khardani, M. Lemdani, and E. Ould Saïd. On the central limit theorem for a conditional mode estimator of a randomly censored time series. Journal of Statistical Theory and Practice, 8(4):722–742, December 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.835755.

Sreedevi:2014:NTC

[407] E. P. Sreedevi, P. G. Sankaran, and P. Dhanavanthan. A nonparametric test for comparing cumulative incidence functions of current status competing risks data. Journal of Statistical Theory and Practice, 8 (4):743–759, December 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.848335.

Esquivel:2014:SAI

[408] Manuel L. Esquível and Pedro P. Mota. On some auto-induced regime switching double-threshold glued diffusions. Journal of Statistical Theory and Practice, 8(4):760–771, December 2014. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.854184.

Gonzalez:2014:EFP

[409] L. M. González, J. M. Singer, and E. J. Stanek III. Estimation of finite population parameters with auxiliary information and response error. Journal of Statistical Theory and Practice, 8(4):772–791, December 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.856358.

Katzoff:2014:ESF

[410] Myron Katzoff, Wen Zhou, Diba Khan, Guanhua Lu, and Benjamin Kedem. Erratum to: Out-of-Sample fusion in risk prediction. Journal of Statistical Theory and Practice, 8(4):792, December 2014. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.942175;   http:
//link.springer.com/content/pdf/
10.1080/15598608.2014.942175.pdf.

See [392].

Coelho:2015:PSI

[411] Carlos A. Coelho, Filipe J. Marques, and Barry C. Arnold. Preface for the special issue on distribution theory, estimation, and inference. Journal of Statistical Theory and Practice, 9(1):1, March 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.975629;   http:
//link.springer.com/content/pdf/
10.1080/15598608.2015.975629.pdf.

Balakrishnan:2015:LTC

[412] Narayanaswamy Balakrishnan and Milan Stehlík. Likelihood testing with censored and missing duration data. Journal of Statistical Theory and Practice, 9(1):2–22, March 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.927811.

Hayter:2015:SST

[413] A. J. Hayter and Jongphil Kim. Smallsample tests for the equality of two normal cumulative probabilities, coefficients of variation, and Sharpe ratios. Journal of Statistical Theory and Practice, 9(1):23–36, March 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.881762.

Marques:2015:NED

[414] Filipe J. Marques and Carlos A. Coelho. Near-exact distributions for the likelihood ratio test statistic for testing multisample independence — the real and complex cases. Journal of Statistical Theory and Practice, 9(1):37–58, March 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.915157.

Balakrishnan:2015:CEP

[415] N. Balakrishnan, C. Koukouvinos, and C. Parpoula. On the computation of entropy prior complexity and marginal prior distribution for the Bernoulli model. Journal of Statistical Theory and Practice, 9(1):59–72, March 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.897139.

Silva:2015:FIG

[416] Rodrigo B. Silva, Marcelo Bourguignon, and Gauss M. Cordeiro. Fréchet and inverse gamma distributions: Correct selection and minimum sample size to discriminate them. Journal of Statistical Theory and Practice, 9(1):73–87, March 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.893730.

Leiva:2015:EBS

[417] Ricardo Leiva, Anuradha Roy, Rubén Bageta, and Juan Carlos Pina. An extension of the Birnbaum–Saunders distribution as a model for fatigue failure due to multiple cracks. Journal of Statistical Theory and Practice, 9(1):88–121, March 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.910149.

Pappas:2015:GEL

[418] Vasileios Pappas, Konstantinos Adamidis, and Sotirios Loukas. A generalization of the exponential–logarithmic distribution. Journal of Statistical Theory and Practice, 9(1): 122–133, March 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.898604.

Cepeda-Cuervo:2015:BRM

[419] Edilberto Cepeda-Cuervo. Beta regression models: Joint mean and variance modeling. Journal of Statistical Theory and Practice, 9(1): 134–145, March 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.890983.

Denker:2015:LQE

[420] Manfred Denker and Lucia Tabacu. Logarithmic quantile estimation for rank statistics. Journal of Statistical Theory and Practice, 9(1): 146–170, March 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.886312. See er-

ratum [526].

vanZyl:2015:ESP

[421] J. Martin van Zyl. Estimation of the shape parameter of a generalized Pareto distribution based on a transformation to Pareto distributed variables.

Journal of Statistical Theory and Practice, 9(1):171–183, March 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.865106.

Neves:2015:MEE

[422] M. Manuela Neves, M. Ivette Gomes, Fernanda Figueiredo, and Dora Prata Gomes. Modeling extreme events: Sample fraction adaptive choice in parameter estimation. Journal of Statistical Theory and Practice, 9 (1):184–199, March 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.890984.

Caeiro:2015:RML

[423] Frederico Caeiro and M. Ivette Gomes. Revisiting the maximum likelihood estimation of a positive extreme value index. Journal of Statistical Theory and Practice, 9(1):200–218, March 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.909754.

Yadav:2015:IFE

[424] Subhash Kumar Yadav, Cem Kadilar, Javid Shabbir, and Sat Gupta. Improved family of estimators of population variance in simple random sampling. Journal of Statistical Theory and Practice, 9(2):219–226, June 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.856359.

Wooff:2015:TWM

[425] David A. Wooff and Jillian M. Anderson. Time-weighted multi-touch attribution and channel relevance in the customer journey to online purchase. Journal of Statistical Theory and Practice, 9(2):227–249, June 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.862753.

Dey:2015:WWD

[426] Sanku Dey, Tanujit Dey, and M. Z. Anis. Weighted Weibull distribution: Properties and estimation. Journal of Statistical Theory and Practice, 9(2):250–265, June 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.875966.

Tang:2015:PMT

[427] Xinyu Tang and Abdus S. Wahed. Pattern-mixture-type estimation and testing of neuroblastoma treatment regimes. Journal of Statistical Theory and Practice, 9(2):266–287, June 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2013.878888.

Coolen-Maturi:2015:NPI

[428] Tahani Coolen-Maturi and Frank P. A. Coolen. Nonparametric predictive inference with combined data under different right-censoring schemes. Journal of Statistical Theory and Practice, 9(2):288–304, June 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2014.886313.

Bevrani:2015:ABD

[429] H. Bevrani and S. Sharifi Far. An approximation to binomial distribution. Journal of Statistical Theory and Practice, 9(2):305–312, June 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2014.892445.

Al-Nasser:2015:ITW

[430] Amjad D. Al-Nasser and Amer I. Al-Omari. Information theoretic weighted mean based on truncated ranked set sampling. Journal of Statistical Theory and Practice, 9(2):313–329, June 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2014.897278.

Burgess:2015:CDC

[431] Leonie Burgess, Stephanie A. Knox, Deborah J. Street, and Richard Norman. Comparing designs constructed with and without priors for choice experiments: a case study. Journal of Statistical Theory and Practice, 9(2):330–360, June 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2014.905223.

Ahmad:2015:ERS

[432] Tanvir Ahmad and Munir Akhtar. Efficient response surface designs for the second-order multivariate polynomial model robust to missing observation.

Journal of Statistical Theory and Practice, 9(2):361–375, June 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.910479.

Dette:2015:ODR

[433] Holger Dette and Christine Kiss. Optimal designs for rational regression models. Journal of Statistical Theory and Practice, 9(2):376–394, June 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.910480.

Arshad:2015:EAS

[434] Mohd. Arshad, Neeraj Misra, and P. Vellaisamy. Estimation after selection from gamma populations with unequal known shape parameters. Journal of Statistical Theory and Practice, 9(2):395–418, June 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.912601.

Gutman:2015:SDM

[435] Alex J. Gutman, Dursun A. Bulutoglu, and Edward D. White. Supersaturated designs with the maximum number of factors for a given resolutionrank. Journal of Statistical Theory and Practice, 9(2):419–435, June 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.916242.

Jacroux:2015:JUF

[436] Mike Jacroux and Bonni Kealy-Dichone. On the joint use of the foldover and partial confounding for the construction of follow-up two-level blocked fractional factorial designs. Journal of Statistical Theory and Practice, 9(2):436–462, June 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.929458.

Hilow:2015:TCM

[437] Hisham Hilow. Three categories of minimum cost systematic full 2^{n} factorial designs. Journal of Statistical Theory and Practice, 9(3): 463–478, September 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.932728.

Meintanis:2015:ADF

[438] Simos G. Meintanis and Efthymios Tsionas. Approximately distributionfree diagnostic tests for regressions with survival data. Journal of Statistical Theory and Practice, 9(3): 479–488, September 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.935616.

Kumar:2015:MGS

[439] C. Satheesh Kumar and M. R. Anusree. On modified generalized skew normal distribution and some of its properties. Journal of Statistical Theory and Practice, 9(3):489–505, September 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.935617.

Singh:2015:AEM

[440] G. N. Singh, M. Khetan, and S. Maurya. Assessment of exponential methods of estimation under nonresponse in two-occasion successive sampling. Journal of Statistical Theory and Practice, 9(3):506–523, September 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.938377.

Best:2015:APT

[441] D. J. Best and J. C. W. Rayner. An alternative to Page’s test permitting both tied and missing data. Journal of Statistical Theory and Practice, 9(3):524–536, September 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.940098.

Nguyen:2015:COB

[442] Nam-Ky Nguyen and Aloke Dey. A catalog of orthogonally blocked threelevel second-order designs with run sizes <= 100. Journal of Statistical Theory and Practice, 9(3): 537–543, September 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.948696.

Ahmad:2015:TGL

[443] M. Rauf Ahmad. U-tests of general linear hypotheses for high-dimensional data under nonnormality and heteroscedasticity. Journal of Statistical Theory and Practice, 9(3):544– 570, September 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2014.968940.

Henriques-Rodrigues:2015:ESS

[444] Lígia Henriques-Rodrigues, M. Ivette Gomes, and B. G. Manjunath. Estimation of a scale second-order parameter related to the PORT methodology. Journal of Statistical Theory and Practice, 9(3):571–599, September 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2014.969816.

Lin:2015:EPE

[445] Y. Lin, A. J. Hayter, and W. Liu. Establishing practical equivalence between three treatments. Journal of Statistical Theory and Practice, 9 (3):600–607, September 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2014.971985.

Hashimoto:2015:MEW

[446] Elizabeth M. Hashimoto, Edwin M. M. Ortega, Gauss M. Cordeiro, and Marcelino A. R. Pascoa. The Mc-Donald extended Weibull distribution. Journal of Statistical Theory and Practice, 9(3):608–632, September 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2014.977980.

Kalucha:2015:REF

[447] Geeta Kalucha, Sat Gupta, and B. K. Dass. Ratio estimation of finite population mean using optional randomized response models. Journal of Statistical Theory and Practice, 9 (3):633–645, September 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.977982.

Davarzani:2015:SIM

[448] Nasser Davarzani, Ahmad Parsian, and Ralf Peeters. Statistical inference on middle-censored data in a dependent setup. Journal of Statistical Theory and Practice, 9(3): 646–657, September 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.978959.

Benavoli:2015:IDP

[449] Alessio Benavoli, Francesca Mangili, Fabrizio Ruggeri, and Marco Zaffalon. Imprecise Dirichlet process with application to the hypothesis test on the probability that X <= Y. Journal of Statistical Theory and Practice, 9 (3):658–684, September 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.985997.

Chakraborty:2015:CRT

[450] Hrishikesh Chakraborty and Genevieve Lyons. Cluster randomized trials: Considerations for design and analysis. Journal of Statistical Theory and Practice, 9(3):685–698, September 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2014.992081.

Kumar:2015:DHM

[451] T. L. Mohan Kumar and Prajneshu. Development of hybrid models for forecasting time-series data using nonlinear SVR enhanced by PSO. Journal of Statistical Theory and Practice, 9(4):699–711, December 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2014.977981.

Li:2015:ESM

[452] Yifang Li and Sujit K. Ghosh. Efficient sampling methods for truncated multivariate normal and studentt distributions subject to linear inequality constraints. Journal of Statistical Theory and Practice, 9(4): 712–732, December 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2014.996690.

Balazs:2015:QIF

[453] Barbara Balázs. m-qualitatively independent families. Journal of Statistical Theory and Practice, 9(4): 733–740, December 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2014.1001046.

Haq:2015:VRS

[454] Abdul Haq, Jennifer Brown, Elena Moltchanova, and Amer Ibrahim Al-Omari. Varied L ranked set sampling scheme. Journal of Statistical Theory and Practice, 9(4): 741–767, December 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.1008606.

Sen:2015:SPR

[455] Sumen Sen, Norou Diawara, and Khan M. Iftekharuddin. Statistical pattern recognition using Gaussian copula. Journal of Statistical Theory and Practice, 9(4):768–777, December 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.1008607.

Singh:2015:CPB

[456] Poonam Singh, Mithilesh Kumar Jha, and Garima Priyadarshini. Constructions of partially balanced crossover designs based on two and higher order association schemes. Journal of Statistical Theory and Practice, 9 (4):778–796, December 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.1019165.

Khattree:2015:MEI

[457] Ravindra Khattree. Mixture experiments in the interior: Yantram designs. Journal of Statistical Theory and Practice, 9(4):797–822, December 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.1029148.

Derrar:2015:NEF

[458] Saliha Derrar, Ali Laksaci, and Elias Ould Saïd. On the nonparametric estimation of the functional ψ-regression for a random lefttruncation model. Journal of Statistical Theory and Practice, 9(4): 823–849, December 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.1032455.

Jensen:2015:ODU

[459] D. R. Jensen and D. E. Ramirez. Outlier detection under star-contoured errors. Journal of Statistical Theory and Practice, 9(4):850–867, December 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic).

URL http://link.springer.com/
article/10.1080/15598608.2015.1034386.

Park:2015:STM

[460] Hyo-Il Park. Simultaneous test for the mean and variance with an application to the statistical process control. Journal of Statistical Theory and Practice, 9(4):868–881, December 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.1034387.

Johanis:2015:TPS

[461] Michal Johanis and Jan Rychtář. A three-player singled out game. Journal of Statistical Theory and Practice, 9(4):882–895, December 2015. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.1042174.

Heyer:2015:IFA

[462] Herbert Heyer. Information functionals with applications to random walk and statistics. Journal of Statistical Theory and Practice, 9(4):

896–933, December 2015. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.1053638.

Saigusa:2016:DSU

[463] Yusuke Saigusa, Kouji Tahata, and Sadao Tomizawa. Decompositions of symmetry using extended palindromic symmetry models for square contingency tables. Journal of Statistical Theory and Practice, 10 (1):1–6, March 2016. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.1048396.

Khan:2016:RBE

[464] Md Hasinur Rahaman Khan and J. Ewart H. Shaw. Robust bias estimation for Kaplan–Meier survival estimator with jackknifing. Journal of Statistical Theory and Practice, 10(1):7–19, March 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.1062833.

Georgiadis:2016:NEI

[465] Stylianos Georgiadis and Nikolaos Limnios. Nonparametric estimation of interval reliability for discretetime semi-Markov systems. Journal of Statistical Theory and Practice, 10(1):20–39, March 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.1064049.

Pathak:2016:NGF

[466] A. K. Pathak and P. Vellaisamy. A note on generalized Farlie–Gumbel-Morgenstern copulas. Journal of Statistical Theory and Practice, 10 (1):40–58, March 2016. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2015.1064838.

Aoki:2016:MCM

[467] Satoshi Aoki, Takayuki Hibi, and Hidefumi Ohsugi. Markov-chain Monte Carlo methods for the Box–Behnken designs and centrally symmetric configurations. Journal of Statistical Theory and Practice, 10(1):59–72, March 2016. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2015.1067172.

Sangpoom:2016:DAL

[468] Suttida Sangpoom and Winai Bodhisuwan. The discrete asymmetric Laplace distribution. Journal of Statistical Theory and Practice, 10 (1):73–86, March 2016. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2015.1067659.

Tcheuko:2016:UAR

[469] Lucas Tcheuko, Brandon Gallas, and Frank Samuelson. Using ANOVA/ random-effects variance estimates to compute a two-sample U-statistic of order (1, 1) estimate of variance. Journal of Statistical Theory and Practice, 10(1):87–99, March 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.1077759.

Chen:2016:NRR

[470] Kun Chen and Kung-Sik Chan. A note on rank reduction in sparse multivariate regression. Journal of Statistical Theory and Practice, 10 (1):100–120, March 2016. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.1081573.

Dong:2016:CBL

[471] Yuexiao Dong and Chaozheng Yang. Cluster-based least absolute deviation regression for dimension reduction. Journal of Statistical Theory and Practice, 10(1):121–132, March 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.1095136.

Zougab:2016:CSB

[472] Nabil Zougab, Smail Adjabi, and Célestin C. Kokonendji. Comparison study to bandwidth selection in binomial kernel estimation using Bayesian approaches. Journal of Statistical Theory and Practice, 10(1):133–153, March 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.1098579.

Mazumder:2016:BND

[473] Satyaki Mazumder and Sourabh Bhattacharya. Bayesian nonparametric dynamic state space modeling with circular latent states. Journal of Statistical Theory and Practice, 10 (1):154–178, March 2016. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.1100562.

Ristic:2016:GEG

[474] Miroslav M. Ristić and Debasis Kundu. Generalized exponential geometric extreme distribution. Journal of Statistical Theory and Practice, 10(1): 179–201, March 2016. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.1101628.

Sengupta:2016:BDI

[475] Pooja Sengupta, N. Rao Chaganty, and Roy T. Sabo. Bivariate doubly inflated Poisson models with applications. Journal of Statistical Theory and Practice, 10(1):202–215, March 2016. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.1105166.

Sharp:2016:IPV

[476] Julia L. Sharp, James Wrenn, and Patrick D. Gerard. Identifying the perceived value of statistical consulting in a university setting. Journal of Statistical Theory and Practice, 10(1):216–225, March 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.1108254.

Chen:2016:NTB

[477] Yuhui Chen. A new type of Bayesian nonparametric control charts for individual measurements. Journal of Statistical Theory and Practice, 10 (1):226–238, March 2016. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.1117405.

Lee:2016:STS

[478] Gi-Sung Lee, Ki-Hak Hong, and Chang-Kyoon Son. A stratified twostage unrelated randomized response model for estimating a rare sensitive attribute based on the Poisson distribution. Journal of Statistical Theory and Practice, 10(2):239–262, June 2016. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.1126696.

Li:2016:RPM

[479] Dandan Li and Siva Sivaganesan. On the role of the prior in multiplicity adjustment. Journal of Statistical Theory and Practice, 10 (2):263–290, June 2016. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.1128858.

Dey:2016:TPM

[480] Sanku Dey, Tanujit Dey, Sajid Ali, and Madhuri S. Mulekar. Two-parameter Maxwell distribution: Properties and different methods of estimation. Journal of Statistical Theory and Practice, 10(2):291–310, June 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2015.1135090.

Braga:2016:OLL

[481] Altemir da Silva Braga, Gauss M. Cordeiro, Edwin M. M. Ortega, and José Nilton da Cruz. The odd log-logistic normal distribution: Theory and applications in analysis of experiments. Journal of Statistical Theory and Practice, 10 (2):311–335, June 2016. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1141127.

Sengupta:2016:ODU

[482] S. Sengupta. Optimal design unbiased strategy for estimating dichotomous finite population variance under Warner’s randomized response plan. Journal of Statistical Theory and Practice, 10(2):336–341, June 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1141128.

Fischer:2016:MSR

[483] Matthias Fischer, Christoph Köstler, and Kevin Jakob. Modeling stochastic recovery rates and dependence between default rates and recovery rates within a generalized credit portfolio framework. Journal of Statistical Theory and Practice, 10(2):342–356, June 2016. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1141733.

Santos:2016:PCM

[484] Rui Santos, Miguel Felgueiras, and João Paulo Martins. Pseudo-convex mixtures generated by shape-extended stable distributions for extremes. Journal of Statistical Theory and Practice, 10(2):357–374, June 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1146929.

Raza:2016:SAM

[485] Mahdi Saber Raza and Mark Broom. Survival analysis modeling with hidden censoring. Journal of Statistical Theory and Practice, 10 (2):375–388, June 2016. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1152205.

Stiburek:2016:SID

[486] David Stiburek. Statistical inference on the drift parameter in symmetric stable Lévy process with a deterministic drift. Journal of Statistical Theory and Practice, 10(2):389–404, June 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1156038.

Kumar:2016:BAG

[487] C. Satheesh Kumar and B. Unnikrishnan Nair. Bivariate alphageneralized hyper-Poisson distribution: Properties and applications. Journal of Statistical Theory and Practice, 10(2):405–419, June 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2016.1156039.

Fuchs:2016:MEV

[488] Mathias Fuchs and Norbert Krautenbacher. Minimization and estimation of the variance of prediction errors for cross-validation designs. Journal of Statistical Theory and Practice, 10(2):420–443, June 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2016.1158675.

Sharma:2016:ECP

[489] M. K. Sharma and B. V. S. Sisodia. Estimation of crop production for smaller geographical area: an application of discriminant function analysis. Journal of Statistical Theory and Practice, 10(2):444–455, June 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2016.1158676.

Some:2016:EAK

[490] Sobom M. Somé and Célestin C. Kokonendji. Effects of associated kernels in nonparametric multiple regressions. Journal of Statistical Theory and Practice, 10(2):456–471, June 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2016.1160010.

Traulsen:2016:BMG

[491] Arne Traulsen. Biological models in game theory. Journal of Statistical Theory and Practice, 10 (2):472–474, June 2016. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1172462.

Ardjoun:2016:RKE

[492] Fatima Zohra Ardjoun, Larbi Ait Hennani, and Ali Laksaci. A recursive kernel estimate of the functional modal regression under ergodic dependence condition. Journal of Statistical Theory and Practice, 10 (3):475–496, September 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1183245.

Sinha:2016:SDR

[493] Sanjoy K. Sinha and Xiaojian Xu. Sequential designs for repeated-measures experiments. Journal of Statistical Theory and Practice, 10(3):497– 514, September 2016. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1184111.

Coolen-Maturi:2016:PIB

[494] Tahani Coolen-Maturi, Frank P. A. Coolen, and Noryanti Muhammad. Predictive inference for bivariate data: Combining nonparametric predictive inference for marginals with an estimated copula. Journal of Statistical Theory and Practice, 10(3): 515–538, September 2016. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1184112.

Chaturvedi:2016:SPC

[495] Anoop Chaturvedi, Ashutosh Kumar Dubey, and Chandra Gulati. Statistical process control for autocorrelated data on grid. Journal of Statistical Theory and Practice, 10 (3):539–549, September 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2016.1185387.

Mourya:2016:CAE

[496] K. K. Mourya, B. V. S. Sisodia, and Hukum Chandra. Calibration approach for estimating finite population parameters in two-stage sampling. Journal of Statistical Theory and Practice, 10 (3):550–562, September 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2016.1189861.

Kumar:2016:ABD

[497] Narinder Kumar and Manesh Chawla. Algorithm-based distribution of twosample statistics. Journal of Statistical Theory and Practice, 10(3): 563–573, September 2016. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2016.1191046.

Wang:2016:AND

[498] Wei Wang and Bo Zhang. Assessing natural direct and indirect effects for a continuous exposure and a dichotomous outcome. Journal of Statistical Theory and Practice, 10 (3):574–587, September 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1203843.

Marzban:2016:CIV

[499] Samira Marzban and Tom Lahmer. Conceptual implementation of the variance-based sensitivity analysis for the calculation of the first-order effects. Journal of Statistical Theory and Practice, 10(4):589–611, December 2016. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1207578.

Salloum:2016:ELC

[500] Zahraa Salloum. Empirical likelihood confidence regions for the parameters of a two phases nonlinear model with and without missing response data. Journal of Statistical Theory and Practice, 10(4):612–638, December 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1212749.

Hughes:2016:RDA

[501] Gary B. Hughes. A review of dimensional assessment for statistics and probability. Journal of Statistical Theory and Practice, 10(4): 639–655, December 2016. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1214852.

Slaoui:2016:CSP

[502] Yousri Slaoui. On the choice of smoothing parameters for semirecursive nonparametric hazard estimators. Journal of Statistical Theory and Practice, 10(4):656–672, December 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2016.1214853.

Hosoya:2016:LRT

[503] Miki Hosoya and Takashi Seo. On the likelihood ratio test for the equality of multivariate normal populations with two-step monotone missing data. Journal of Statistical Theory and Practice, 10(4):673–692, December 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2016.1215942.

Ghosh:2016:SAB

[504] Indranil Ghosh and Samik Ray. Some alternative bivariate Kumaraswamytype distributions via copula with application in risk management. Journal of Statistical Theory and Practice, 10(4):693–706, December 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2016.1215943.

Kumar:2016:IEP

[505] Sunil Kumar. Improved estimation of population mean in presence of nonresponse and measurement error. Journal of Statistical Theory and Practice, 10(4):707–720, December 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2016.1216488. See

note [634].

Al-Omari:2016:NME

[506] Amer Ibrahim Al-Omari. A new measure of entropy of continuous random variable. Journal of Statistical Theory and Practice, 10(4): 721–735, December 2016. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1217444.

Tuyl:2016:CPM

[507] Frank Tuyl, Richard Gerlach, and Kerrie Mengersen. Consensus priors for multinomial and binomial ratios. Journal of Statistical Theory and Practice, 10(4):736–754, December 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1219684.

Khan:2016:IRT

[508] Lakhkar Khan, Javid Shabbir, and Sat Gupta. Improved ratio-type estimators using stratified doubleranked set sampling. Journal of Statistical Theory and Practice, 10 (4):755–767, December 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1219685.

Park:2016:RMA

[509] Jason Hong Jae Park. A random measure algebra under convolution. Journal of Statistical Theory and Practice, 10(4):768–779, December 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1224745.

Ortega:2016:OBS

[510] Edwin M. M. Ortega, Artur J. Lemonte, Gauss M. Cordeiro, and José Nilton da Cruz. The odd Birnbaum–Saunders regression model with applications to lifetime data. Journal of Statistical Theory and Practice, 10(4):780–804, December 2016. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2016.1224746.

Esquivel:2016:SSM

[511] Manuel L. Esquível, Pedro P. Mota, and João Tiago Mexia. On some statistical models with a random number of observations. Journal of Statistical Theory and Practice, 10 (4):805–823, December 2016. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2016.1227735.

Mostafa:2016:RLS

[512] Sayed A. Mostafa and Ibrahim A. Ahmad. Remainder linear systematic sampling with multiple random starts. Journal of Statistical Theory and Practice, 10(4):824–851, December 2016. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2016.1231094.

Biswas:2016:CAR

[513] Atanu Biswas and Rahul Bhattacharya. A covariate-adjusted response-adaptive allocation for a general class of continuous responses. Journal of Statistical Theory and Practice, 10(4):852–863, December 2016.

CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1232207.

Krenek:2017:DSC

[514] Russell Krenek, Jinho Cha, Byung Rae Cho, and Julia L. Sharp. Development of statistical convolutions of truncated normal and truncated skew normal distributions with applications. Journal of Statistical Theory and Practice, 11(1):1–25, March 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1244022.

Saghir:2017:MLB

[515] Aamir Saghir, Aneeqa Khadim, and Zhengyan Lin. The Maxwell lengthbiased distribution: Properties and estimation. Journal of Statistical Theory and Practice, 11(1):26–40, March 2017. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1246266.

Singh:2017:WTB

[516] Charan Jeet Singh, Sandeep Kaur, and Madhu Jain. Waiting time of bulk arrival unreliable queue with balking and Bernoulli feedback using maximum entropy principle. Journal of Statistical Theory and Practice, 11(1):41–62, March 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1251365.

Mishra:2017:NVB

[517] Nutan Mishra. Nonbinary variancebalanced designs-part i, optimality. Journal of Statistical Theory and Practice, 11(1):63–75, March 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2016.1258348.

Khan:2017:IMU

[518] N. Mamode Khan, Y. Sunecher, and V. Jowaheer. Inferential methods for an unconstrained nonstationary BINMA time series process with Poisson innovations. Journal of Statistical Theory and Practice, 11(1):76–106, March 2017. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2016.1258600.

Ciuperca:2017:AFL

[519] Gabriela Ciuperca. Adaptive fused LASSO in grouped quantile regression. Journal of Statistical Theory and Practice, 11(1):107–125, March 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2016.1258601.

Chen:2017:SRC

[520] Yuhui Chen and Timothy Hanson. Semiparametric regression control charts. Journal of Statistical Theory and Practice, 11(1): 126–144, March 2017. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2016.1260502.

Shin:2017:CSD

[521] Seung Jun Shin and Sujit K. Ghosh. A comparative study of the doseresponse analysis with application to the target dose estimation. Journal of Statistical Theory and Practice, 11(1):145–162, March 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1261260.

Xue:2017:ODE

[522] Xiaoqiang Xue and Valerii V. Fedorov. Optimal design of experiments with the observation censoring driven by random enrollment of subjects. Journal of Statistical Theory and Practice, 11(1):163–178, March 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1263808.

Alizadeh:2017:GGF

[523] Morad Alizadeh, Gauss M. Cordeiro, Luis Gustavo Bastos Pinho, and Indranil Ghosh. The Gompertz-g family of distributions. Journal of Statistical Theory and Practice, 11(1): 179–207, March 2017. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1267668.

Rayner:2017:EA

[524] J. C. W. Rayner. Extended ANOVA. Journal of Statistical Theory and Practice, 11(1):208–219, March 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1268550.

Gamage:2017:AEL

[525] Ramadha D. Piyadi Gamage, Wei Ning, and Arjun K. Gupta. Adjusted empirical likelihood for longmemory time-series models. Journal of Statistical Theory and Practice, 11(1):220–233, March 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1271373.

Denker:2017:ELQ

[526] Manfred Denker and Lucia Tabacu. Erratum to: Logarithmic quantile estimation for rank statistics. Journal of Statistical Theory and Practice, 11(1):234, March 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1254693;   http:
//link.springer.com/content/pdf/
10.1080/15598608.2016.1254693.pdf.

See [420].

Stufken:2017:P

[527] John Stufken, Abhyuday Mandal, and Manohar Aggarwal. Preface. Journal of Statistical Theory and Practice, 11(2):235, June 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1321281;   http:
//link.springer.com/content/pdf/
10.1080/15598608.2017.1321281.pdf.

Grossmann:2017:PPC

[528] Heiko Großmann. Partial-profile choice designs for estimating main effects and interactions of two-level attributes from paired comparison data. Journal of Statistical Theory and Practice, 11(2):236–253, June 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1197868.

Tian:2017:ECE

[529] T. Tian and M. Yang. Efficiency of the coordinate-exchange algorithm in constructing exact optimal discrete choice experiments. Journal of Statistical Theory and Practice, 11(2):254–268, June 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1203842.

Zhang:2017:AIM

[530] Wei Zhang, Abhyuday Mandal, and John Stufken. Approximations of the information matrix for a panel mixed logit model. Journal of Statistical Theory and Practice, 11 (2):269–295, June 2017. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1219288.

Ribeiro:2017:PDI

[531] Tiago Ribeiro, Richard Carson, Jordan J. Louviere, and John M. Rose. Possible design-induced artifacts associated with designs for discrete choice experiments. Journal of Statistical Theory and Practice, 11(2):296–321, June 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2016.1258680.

Sun:2017:EDC

[532] Fangfang Sun and Angela Dean. Aefficient discrete choice designs for attributes with unequal numbers of levels. Journal of Statistical Theory and Practice, 11(2):322–338, June 2017. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2017.1292482.

Jaynes:2017:MAD

[533] Jessica Jaynes, Hongquan Xu, and Weng Kee Wong. Minimum aberration designs for discrete choice experiments. Journal of Statistical Theory and Practice, 11(2):339–360, June 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2017.1299055.

Benda:2017:ADQ

[534] Norbert Benda, Paul-Christian Bürkner, Fritjof Freise, Heinz Holling, and Rainer Schwabe. Adaptive designs for quantal dose-response experiments with false answers. Journal of Statistical Theory and Practice, 11(3): 361–374, September 2017. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2016.1213209.

Steland:2017:AFS

[535] Ansgar Steland. On the accuracy of fixed sample and fixed width confidence intervals based on the vertically weighted average. Journal of Statistical Theory and Practice, 11 (3):375–392, September 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1263809.

Rasch:2017:HRM

[536] Dieter Rasch and Takuya Yanagida. How robust is the modified sequential triangular test of a correlation coefficient against nonnormality of the basic variables? Journal of Statistical Theory and Practice, 11 (3):393–401, September 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1263810.

Yanagida:2017:SRP

[537] Takuya Yanagida. seqtest: an R package for sequential triangular tests. Journal of Statistical Theory and Practice, 11(3):402–406, September 2017. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1268549.

Andronov:2017:RRE

[538] Alexander Andronov. On a reward rate estimation for the finite irreducible continuous-time Markov chain. Journal of Statistical Theory and Practice, 11(3):407–417, September 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1282895.

Gares:2017:FHT

[539] Valérie Garès, Sandrine Andrieu, Jean-François Dupuy, and Nicolas Savy.

On the Fleming–Harrington test for late effects in prevention randomized controlled trials. Journal of Statistical Theory and Practice, 11 (3):418–435, September 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1295889.

Arboretti:2017:MEU

[540] Rosa Arboretti, Eleonora Carrozzo, Fortunato Pesarin, and Luigi Salmaso. A multivariate extension of unionintersection permutation solution for two-sample testing. Journal of Statistical Theory and Practice, 11(3): 436–448, September 2017. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1295891.

Tarima:2017:SEP

[541] Sergey Tarima. Statistical estimation in the presence of possibly incorrect model assumptions. Journal of Statistical Theory and Practice, 11 (3):449–467, September 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1299056.

Happ:2017:HDR

[542] Martin Happ, Solomon W. Harrar, and Arne C. Bathke. High-dimensional repeated measures. Journal of Statistical Theory and Practice, 11(3): 468–477, September 2017. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1307792;   http:

//link.springer.com/content/pdf/
10.1080/15598608.2017.1307792.pdf.

Einbeck:2017:SCB

[543] Jochen Einbeck and Simos Meintanis. Self-consistency-based tests for bivariate distributions. Journal of Statistical Theory and Practice, 11 (3):478–492, September 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1318098.

Yanagida:2017:RTP

[544] Takuya Yanagida, Dieter Rasch, Klaus D. Kubinger, and Berthold Schneider. Robustness of the test of a product moment correlation coefficient under nonnormality. Journal of Statistical Theory and Practice, 11 (3):493–502, September 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1329102.

Barbiero:2017:EMM

[545] Alessandro Barbiero. Estimating a multivariate model with discrete Weibull margins. Journal of Statistical Theory and Practice, 11(4): 503–514, December 2017. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1292483.

Chen:2017:CRM

[546] Yuhui Chen and Timothy Hanson. Copula regression models for discrete and mixed bivariate responses. Journal of Statistical Theory and Practice, 11(4):515–530, December 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2016.1278059.

Hoxhaj:2017:BPS

[547] Valmira Hoxhaj and Ravindra Khattree. Beyond Q–Q plots: Some new graphical tools for the assessment of distributional assumptions and the tail behavior of the data. Journal of Statistical Theory and Practice, 11 (4):531–552, December 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1299057.

Djerroud:2017:BAB

[548] Lamia Djerroud, Nabil Zougab, and Smail Adjabi. Bayesian approach to bandwidth selection for multivariate count regression function estimation by associated discrete kernel. Journal of Statistical Theory and Practice, 11(4):553–572, December 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1281180.

Balakrishnan:2017:NCS

[549] N. Balakrishnan, Helton Saulo, and Jeremias Leão. On a new class of skewed Birnbaum–Saunders models. Journal of Statistical Theory and Practice, 11(4):573–593, December 2017. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1286275.

Tilbi:2017:CSG

[550] D. Tilbi and N. Seddik-Ameur. Chi-squared goodness-of-fit tests for the generalized Rayleigh distribution. Journal of Statistical Theory and Practice, 11(4):594–603, December 2017. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1286276.

Chawsheen:2017:STS

[551] Tara Ahmed Chawsheen and Mark Broom. Seasonal time-series modeling and forecasting of monthly mean temperature for decision making in the Kurdistan Region of Iraq. Journal of Statistical Theory and Practice, 11(4):604–633, December 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1292484.

Miranda:2017:FC

[552] Enrique Miranda and Marco Zaffalon. Full conglomerability. Journal of Statistical Theory and Practice, 11(4):634–669, December 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1295890.

Muralidharan:2017:ALM

[553] K. Muralidharan and Bavagosai Pratima. Analysis of lifetime model with discrete mass at zero and one. Journal of Statistical Theory and Practice, 11(4):670–692, December 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2017.1303407.

Mazumder:2017:NDS

[554] Satyaki Mazumder and Sourabh Bhattacharya. Nonparametric dynamic state space modeling of observed circular time series with circular latent states: a Bayesian perspective. Journal of Statistical Theory and Practice, 11(4):693–718, December 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2017.1305922.

Biswas:2017:AAD

[555] Atanu Biswas, Rahul Bhattacharya, and Taranga Mukherjee. An adaptive allocation design for circular treatment outcome. Journal of Statistical Theory and Practice, 11(4): 719–730, December 2017. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2017.1307147.

Ahmad:2017:LIT

[556] M. Rauf Ahmad. Location-invariant tests of homogeneity of large-dimensional covariance matrices. Journal of Statistical Theory and Practice, 11(4): 731–745, December 2017. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2017.1308895.

Ghosh:2017:GGM

[557] Himadri Ghosh and Prajneshu. Gompertz growth model in random environment with time-dependent diffusion.

Journal of Statistical Theory and Practice, 11(4):746–758, December 2017. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1309307.

Cabrera:2017:ODM

[558] Jesús Alonso Cabrera and Rainer Schwabe. D-optimal design for a model with interaction between a qualitative and a quantitative factor in the presence of random block effects. Journal of Statistical Theory and Practice, 11(4):759–770, December 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1321070.

Chahad:2017:FLL

[559] Abdelkader Chahad, Larbi Ait-Hennani, and Ali Laksaci. Functional local linear estimate for functional relative-error regression. Journal of Statistical Theory and Practice, 11(4):771–789, December 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1321071.

Singh:2017:ESR

[560] Housila P. Singh and Swarangi M. Gorey. An efficient stratified randomized response model. Journal of Statistical Theory and Practice, 11 (4):790–809, December 2017. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1350607.

Coelho:2018:P

[561] Carlos A. Coelho, Thomas Mathew, and Nitis Mukhopadhyay. Preface. Journal of Statistical Theory and Practice, 12(1):1–2, March 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2018.1429200.

Knapp:2018:ECI

[562] Guido Knapp. An exact confidence interval for a common effect size. Journal of Statistical Theory and Practice, 12(1):3–11, March 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2016.1278060.

Nummi:2018:SMR

[563] Tapio Nummi, Janne Salonen, Lasse Koskinen, and Jianxin Pan. A semiparametric mixture regression model for longitudinal data. Journal of Statistical Theory and Practice, 12 (1):12–22, March 2018. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2017.1298062.

Wu:2018:ACE

[564] Jing Wu, Mário de Castro, Elizabeth D. Schifano, and Ming-Hui Chen. Assessing covariate effects using Jeffreys-type prior in the Cox model in the presence of a monotone partial likelihood. Journal of Statistical Theory and Practice, 12(1):23–41, March 2018. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1299058.

Gibbons:2018:OPL

[565] Robert D. Gibbons. OxyELITE Pro and liver disease: Statistical assessment of an apparent association. Journal of Statistical Theory and Practice, 12(1):42–47, March 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1305923.

Tanaka:2018:IEU

[566] Hidekazu Tanaka, Nabendu Pal, and Wooi K. Lim. On improved estimation under Weibull model. Journal of Statistical Theory and Practice, 12(1):48–65, March 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1305921.

Yang:2018:SAN

[567] Ye Yang and Thomas Mathew. The simultaneous assessment of normality and homoscedasticity in linear fixed effects models. Journal of Statistical Theory and Practice, 12 (1):66–81, March 2018. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1320243.

Chai:2018:OTL

[568] Feng-Shun Chai, Ashish Das, and Rakhi Singh. Optimal two-level choice designs for estimating main and specified two-factor interaction effects. Journal of Statistical Theory and Practice, 12(1):82–92, March 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1329101.

Chattopadhyay:2018:CPE

[569] Gaurangadeb Chattopadhyay and Bikas K. Sinha. Coverage probability and exact inference. Journal of Statistical Theory and Practice, 12(1):93–99, March 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1329674.

Klein:2018:SDC

[570] Martin Daniel Klein and Gauri Sankar Datta. Statistical disclosure control via sufficiency under the multiple linear regression model. Journal of Statistical Theory and Practice, 12(1):100–110, March 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1350606.

Mukhopadhyay:2018:MLT

[571] Nitis Mukhopadhyay and Shelemyahu Zacks. Modified Linex two-stage and purely sequential estimation of the variance in a normal distribution with illustrations using horticultural data. Journal of Statistical Theory and Practice, 12(1):111–135, March 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1350608.

SahaRay:2018:OBM

[572] Rita SahaRay and Ganesh Dutta. On the optimality of blocked main effects plans with even number of runs. Journal of Statistical Theory and Practice, 12(1):136–150, March 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1378138.

Alvo:2018:PEN

[573] Mayer Alvo, Tze Leung Lai, and Philip L. H. Yu. Parametric embedding of nonparametric inference problems. Journal of Statistical Theory and Practice, 12(1):151–164, March 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1399840.

Raqab:2018:PRT

[574] M. Z. Raqab, O. M. Bdair, M. T. Madi, and F. A. Alqallaf. Prediction of the remaining testing time for the generalized Pareto progressive censoring samples with applications to extreme hydrology events. Journal of Statistical Theory and Practice, 12(2):165–187, June 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1338168.

Guha:2018:CAB

[575] Saurav Guha, U. C. Sud, and Hukum Chandra. Calibration approach-based chain ratio type and chain product type estimators in two-phase sampling involving two auxiliary variables. Journal of Statistical Theory and Practice, 12(2):188–205, June 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1341354.

Henriques-Rodrigues:2018:LIR

[576] Lígia Henriques-Rodrigues and M. Ivette Gomes. Location-invariant reducedbias tail index estimation under a third-order framework. Journal of Statistical Theory and Practice, 12 (2):206–230, June 2018. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1342577.

Sel:2018:BML

[577] Sanghun Sel, Myoungjin Jung, and Younshik Chung. Bayesian and maximum likelihood estimations from parameters of McDonald extended Weibull model based on progressive type-II censoring. Journal of Statistical Theory and Practice, 12 (2):231–254, June 2018. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1343693.

Baragona:2018:ELO

[578] Roberto Baragona, Francesco Battaglia, and Domenico Cucina. Empirical likelihood for outlier detection in regression models. Journal of Statistical Theory and Practice, 12 (2):255–281, June 2018. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1344940.

Haq:2018:ICE

[579] Abdul Haq and Javid Shabbir. An improved class of estimators of finite population mean in simple random sampling using an auxiliary attribute. Journal of Statistical Theory and Practice, 12(2):282–289, June 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1349012.

A:2018:RDS

[580] Sayed A and Ibrahim A. Recent developments in systematic sampling: a review. Journal of Statistical Theory and Practice, 12(2):290–310, June 2018. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1353456.

Khalil:2018:GEF

[581] Sadia Khalil, Sat Gupta, and Muhammad Hanif. A generalized estimator for finite population mean in the presence of measurement errors in stratified random sampling. Journal of Statistical Theory and Practice, 12(2):311–324, June 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1370621.

Shukla:2018:DSR

[582] Dhyanesh Shukla and B. V. S. Sisodia. Double sampling regression estimator of population mean subsequent to preliminary test of significance in finite population survey sampling. Journal of Statistical Theory and Practice, 12(2):325–335, June 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2017.1377652.

Li:2018:CSR

[583] Zhiming Li, Zhidong Teng, Lijun Wu, and Runchu Zhang. Construction of some 3^{n−m} regular designs with general minimum lower order confounding. Journal of Statistical Theory and Practice, 12(2):336–355, June 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2017.1381057.

Wang:2018:NAM

[584] Dabuxilatu Wang. A note on autoregressive models with fuzzy random variables. Journal of Statistical Theory and Practice, 12 (2):356–369, June 2018. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2017.1389664.

Xu:2018:PMR

[585] Xiaojian Xu and Xiaoyu Chen. A practical method of robust estimation in case of asymmetry. Journal of Statistical Theory and Practice, 12(2):370–396, June 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2017.1393779.

Seddik-Ameur:2018:TFA

[586] Nacira Seddik-Ameur and Wafa Treidi. On testing the fit of accelerated failure time and proportional hazard Weibull extension models. Journal of Statistical Theory and Practice, 12(2):397–411, June 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1397568.

Khan:2018:CAU

[587] Zaheen Khan. A class of almost unbiased estimators of population mean using auxiliary variable and auxiliary attribute in systematic sampling. Journal of Statistical Theory and Practice, 12(2):412–422, June 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1399841.

Zaslavsky:2018:MTC

[588] Boris G. Zaslavsky. Multiple testing with close to equally correlated structure. Journal of Statistical Theory and Practice, 12(2):423–435, June 2018. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1408508.

Ghosh:2018:RBD

[589] Indranil Ghosh and Gholamhossein Hamedani. On the ristic–Balakrishnan distribution: Bivariate extension and characterizations. Journal of Statistical Theory and Practice, 12 (2):436–449, June 2018. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1410264.

Lee:2018:DRG

[590] Jaylen Lee, Shannon Ciccarello, Mithun Acharjee, and Kumer Das. Dimension reduction of gene expression data. Journal of Statistical Theory and Practice, 12(2):450–461, June 2018. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2017.1413456.

Felix-Medina:2018:CCS

[591] Martín H. Félix-Medina. Combining cluster sampling and link-tracing sampling to estimate the size of a hidden population: Asymptotic properties of the estimators. Journal of Statistical Theory and Practice, 12 (3):463–496, September 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2017.1405374.

Chandra:2018:SAE

[592] Hukum Chandra, Kaustav Aditya, and Sushil Kumar. Small-area estimation under a log-transformed area-level model. Journal of Statistical Theory and Practice, 12(3): 497–505, September 2018. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2017.1415174.

Gupta:2018:UMR

[593] Sat Gupta, Samridhi Mehta, Javid Shabbir, and Sadia Khalil. A unified measure of respondent privacy and model efficiency in quantitative RRT models. Journal of Statistical Theory and Practice, 12(3):

506–511, September 2018. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2017.1415175.

Al-Omari:2018:REU

[594] Amer Ibrahim Al-Omari and Amjad D. Al-Nasser. Ratio estimation using multistage median ranked set sampling approach. Journal of Statistical Theory and Practice, 12 (3):512–529, September 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2018.1425168.

Samawi:2018:RSS

[595] Hani M. Samawi, Amal Helu, Haresh Rochani, JingJing Yin, Lili Yu, and Robert Vogel. Reducing sample size needed for accelerated failure time model using more efficient sampling methods. Journal of Statistical Theory and Practice, 12(3): 530–541, September 2018. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2018.1431574.

El-Sagheer:2018:EMW

[596] Rashad M. El-Sagheer and Nagwa M. Mohamed. On estimation of modified Weibull parameters in presence of accelerated life test. Journal of Statistical Theory and Practice, 12 (3):542–560, September 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2018.1431575.

Rayner:2018:ECM

[597] J. C. W. Rayner and D. J. Best. Extensions to the Cochran–Mantel–Haenszel mean scores and correlation tests. Journal of Statistical Theory and Practice, 12(3):561–574, September 2018. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2018.1439424.

Street:2018:OOA

[598] Deborah J. Street and Emily M. Bird. D-optimal orthogonal array minus t run designs. Journal of Statistical Theory and Practice, 12 (3):575–594, September 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2018.1441081.

Kundu:2018:UBG

[599] Debasis Kundu and Vahid Nekoukhou. Univariate and bivariate geometric discrete generalized exponential distributions. Journal of Statistical Theory and Practice, 12 (3):595–614, September 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1080/15598608.2018.1441082.

Sankaran:2018:NCQ

[600] Paduthol Godan Sankaran and M. Dileep Kumar. A new class of quantile functions useful in reliability analysis. Journal of Statistical Theory and Practice, 12(3):615–634, September 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2018.1448732.

Kim:2018:CCR

[601] Yoonsang Kim and Dulal K. Bhaumik. Calibration confidence regions for asbestos fibers with heteroscedasticity and interlaboratory variability. Journal of Statistical Theory and Practice, 12(3):635–656, September 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2018.1449685.

Kawasaki:2018:ADT

[602] Tamae Kawasaki, Nobumichi Shutoh, and Takashi Seo. On the asymptotic distribution of T^{2}-type statistic with two-step monotone missing data. Journal of Statistical Theory and Practice, 12(3):657–668, September 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2018.1450795.

Awwad:2018:OTS

[603] Raed R. Abu Awwad, Omar M. Bdair, and Ghassan K. Abufoudeh. One- and two-sample prediction for the progressively censored Rayleigh residual data. Journal of Statistical Theory and Practice, 12(4):669–687, December 2018. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2018.1450796.

Martinez:2018:NSD

[604] Edson Z. Martinez and Jorge A. Achcar. A new straightforward defective distribution for survival analysis in the presence of a cure fraction. Journal of Statistical Theory and Practice, 12(4):688–703, December 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2018.1460885.

Sohail:2018:CRT

[605] Muhammad Umair Sohail, Javid Shabbir, and Shakeel Ahmed. A class of ratio type estimators for imputing the missing values under rank set sampling. Journal of Statistical Theory and Practice, 12(4):704–717, December 2018. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2018.1460886.

daCunha:2018:SIV

[606] Enai Taveira da Cunha, Klaus L. P. Vasconcellos, and Marcelo Bourguignon. A skew integer-valued timeseries process with generalized Poisson difference marginal distribution. Journal of Statistical Theory and Practice, 12(4):718–743, December 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2018.1470046.

Wu:2018:SST

[607] Yanhong Wu. Supplementary score test for sparse signals in large-scale truncated one-sided sequential tests. Journal of Statistical Theory and Practice, 12(4):744–756, December 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2018.1472681.

Jensen:2018:SOS

[608] Donald R. Jensen and Donald E. Ramirez. Seeking outlying subsets under star-contoured errors. Journal of Statistical Theory and Practice, 12(4):757–775, December 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2018.1473183.

Malekzadeh:2018:TEQ

[609] Ahad Malekzadeh and Ali Akbar Jafari. Testing equality of quantiles of two-parameter exponential distributions under progressive type II censoring. Journal of Statistical Theory and Practice, 12(4): 776–793, December 2018. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2018.1474147.

Nadeb:2018:SCS

[610] Hossein Nadeb and Hamzeh Torabi. Stochastic comparisons of series systems with independent heterogeneous Lomax–exponential components. Journal of Statistical Theory and Practice, 12(4):794–812, December 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2018.1474819.

Bhattacharya:2018:TSA

[611] Rahul Bhattacharya and Madhumita Shome. A two-stage adaptive allocation design for survival outcome with informative censoring. Journal of Statistical Theory and Practice, 12 (4):813–836, December 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2018.1479992.

Basalamah:2018:BSD

[612] Doaa Basalamah, Wei Ning, and Arjun Gupta. The beta skew t distribution and its properties. Journal of Statistical Theory and Practice, 12(4):837–860, December 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2018.1481468.

Marques:2018:STE

[613] Filipe J. Marques and Carlos A. Coelho. The simultaneous test of equality and circularity of several covariance matrices. Journal of Statistical Theory and Practice, 12(4): 861–885, December 2018. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2018.1482806.

Aragon:2018:MLB

[614] Davi C. Aragon, Jorge A. Achcar, and Edson Z. Martinez. Maximum likelihood and Bayesian estimators for the double Poisson distribution. Journal of Statistical Theory and Practice, 12(4):886–911, December 2018. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1080/15598608.2018.1489919.

Ghosh:2019:OFR

[615] Himadri Ghosh and Prajneshu. Optimum fitting of Richards growth model in random environment. Journal of Statistical Theory and Practice, 13(1): ??, March 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-018-
0004-9.

Haouas:2019:ESO

[616] Nawel Haouas, Abdelhakim Necir, and Brahim Brahimi. Estimating the second-order parameter of regular variation and bias reduction in tail index estimation under random truncation. Journal of Statistical Theory and Practice, 13(1):??, March 2019. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-018-0017-4.

Rapallo:2019:ACR

[617] Fabio Rapallo and Maria-Piera Rogantin. Algebraic characterization of C-regular fractions under level permutations. Journal of Statistical Theory and Practice, 13(1):??, March 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-018-0012-9.

Priyanka:2019:SCE

[618] Kumari Priyanka and Pidugu Trisandhya. Some classes of estimators for sensitive population mean on successive moves. Journal of Statistical Theory and Practice, 13(1):??, March 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-018-0008-5.

Dey:2019:SVE

[619] Arabin Kumar Dey and Biplab Paul. Some variations of EM algorithms for Marshall–Olkin bivariate Pareto distribution with location and scale. Journal of Statistical Theory and Practice, 13(1):??, March 2019. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s42519-018-0009-4.

Shirazi:2019:NND

[620] Esmaeil Shirazi and Yogendra P. Chaubey. Non-negative density estimation via wavelet block thresholding for biased data. Journal of Statistical Theory and Practice, 13(1):??, March 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s42519-018-0019-2.

Alqifari:2019:RNP

[621] Hana N. Alqifari and Frank P. A. Coolen. Robustness of nonparametric predictive inference for future order statistics. Journal of Statistical Theory and Practice, 13(1):??, March 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s42519-018-0011-x;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s42519-018-0011-x.pdf.

Bordes:2019:EPD

[622] Laurent Bordes, Christian Paroissin, and Ali Salami. Extension of the parametric delta method with an application to an inference method for a degradation model. Journal of Statistical Theory and Practice, 13(1):??, March 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-018-
0001-z.

El-Sagheer:2019:IWE

[623] Rashad M. El-Sagheer, Mohamed A. W. Mahmoud, and Heba Nagaty. Inferences for Weibull–exponential distribution based on progressive Type– II censoring under step–stress partially accelerated life test model. Journal of Statistical Theory and Practice, 13(1): ??, March 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-018-
0018-3.

Marques:2019:INP

[624] Filipe J. Marques, Frank P. A. Coolen, and Tahani Coolen-Maturi. Introducing nonparametric predictive inference methods for reproducibility of likelihood ratio tests. Journal of Statistical Theory and Practice, 13(1):??, March 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-018-
0020-9.

Santana:2019:GBW

[625] Tiago V. F. Santana, Edwin M. M. Ortega, and Gauss M. Cordeiro. Generalized beta Weibull linear model: Estimation, diagnostic tools and residual analysis. Journal of Statistical Theory and Practice, 13(1):??, March 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-018-0022-7.

Marques:2019:ALR

[626] Filipe Marques, Frank Coolen, and Tahani Coolen-Maturi. Approximations for the likelihood ratio statistic for hypothesis testing between two beta distributions. Journal of Statistical Theory and Practice, 13(1):??, March 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s42519-018-
                                        0021-8.

Barczy:2019:LSE

[627] Mátyás Barczy, Balázs Nyul, and Gyula Pap. Least-squares estimation for the subcritical Heston model based on continuous-time observations. Journal of Statistical Theory and Practice, 13(1):??, March 2019. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s42519-018-0007-6.

Singh:2019:EPM

[628] G. N. Singh and S. Suman. Estimation of population mean using imputation methods for missing data under two-phase sampling design. Journal of Statistical Theory and Practice, 13(1): ??, March 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s42519-018-
                                        0016-5.

Kpanzou:2019:BIE

[629] Tchilabalo A. Kpanzou, Kokouvi M. Gamado, and Hippolyte Hounnon. A Beran-inspired estimator for the Weibull-type tail coefficient. Journal of Statistical Theory and Practice, 13(1):

??, March 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-018-
0013-8.

Davarzani:2019:IMO

[630] Nasser Davarzani, Leila Golparvar, and Ahmad Parsian. Inference on Marshall–Olkin bivariate exponential in the presence of dependent left censoring. Journal of Statistical Theory and Practice, 13(1):??, March 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-018-0003-x.

Kumari:2019:ETP

[631] Taruna Kumari, Ajit Chaturvedi, and Anupam Pathak. Estimation and testing procedures for the reliability functions of Kumaraswamy-g distributions and a characterization based on records. Journal of Statistical Theory and Practice, 13(1):??, March 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-018-0014-7.

Hussain:2019:NFE

[632] Ishfaq Hussain and Abdul Haq. A new family of estimators for population mean with dual use of the auxiliary information. Journal of Statistical Theory and Practice, 13(1):??, March 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-018-0023-6.

Sevilimedu:2019:AMS

[633] Varadan Sevilimedu, Lili Yu, Hani Samawi, and Haresh Rochani. Application of the misclassification simulation extrapolation procedure to log– logistic accelerated failure time models in survival analysis. Journal of Statistical Theory and Practice, 13(1):??, March 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s42519-018-
                                        0024-5.

Sabir:2019:NKI

[634] Shaista Sabir and Aamir Sanaullah. A note on Kumar (2016): Improved Estimation of Population Mean in the Presence of Measurement Error and Nonresponse. Journal of Statistical Theory and Practice, 13(1):??, March 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s42519-018-0025-4. See [505].

Weiss:2019:MDP

[635] Christian H. Weiß and Boris Aleksandrov. Model diagnostics for Poisson INARMA processes using bivariate dispersion indexes. Journal of Statistical Theory and Practice, 13(2):??, June 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s42519-018-
                                        0028-1.

Hashimoto:2019:LBX

[636] Elizabeth M. Hashimoto, Giovana O. Silva, Edwin M. M. Ortega, and Gauss M. Cordeiro. Log-Burr XII gamma–Weibull regression model with random effects and censored data. Journal of Statistical Theory and Practice, 13(2):??, June 2019. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-018-0026-3.

Cruz:2019:WSE

[637] Juan Alberto Rojas Cruz and Iesus C. Diniz. On weak and strong ergodicity. Journal of Statistical Theory and Practice, 13(2):??, June 2019. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-018-0029-0.

Cahoy:2019:ISP

[638] Dexter Cahoy and Joseph Sedransk. Inverse stable prior for exponential models. Journal of Statistical Theory and Practice, 13(2):??, June 2019. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-018-0027-2.

Samawi:2019:KBM

[639] Hani Samawi, Haresh Rochani, JingJing Yin, and Robert Vogel. On kernelbased mode estimation using different stratified sampling designs. Journal of Statistical Theory and Practice, 13(2): ??, June 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-018-
0034-3.

Rasheed:2019:MRM

[640] Muhammad Rasheed, Muhammad Ismail, Farrukh Shehzad, Rashid Ahmed, and Akbar Ali. Minimal repeated measurements designs balanced for residual effects in circular periods of three different sizes. Journal of Statistical Theory and Practice, 13(2):??, June 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-018-0030-7.

Mohanty:2019:ISG

[641] Bigyanananda Mohanty and A. K. P. C. Swain. Identifying social groupwise household-level determinants of poverty in rural Odisha using logistic regression model. Journal of Statistical Theory and Practice, 13(2):??, June 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-018-
0032-5.

Sango:2019:EVM

[642] Joël Sango and Pierre Duchesne. Evaluating vector multiplicative error models with the Hosking–Ljung–Box portmanteau test and kernel-based test statistics. Journal of Statistical Theory and Practice, 13(2):??, June 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-018-0036-1.

Velina:2019:ELB

[643] Mara Velina, Janis Valeinis, and George Luta. Empirical likelihoodbased inference for the difference of two location parameters using smoothed M-estimators. Journal of Statistical Theory and Practice, 13(2):??, June 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-019-
0037-8.

Gribkova:2019:ATF

[644] Nadezhda Gribkova and Ricardas Zitikis. Assessing transfer functions in control systems. Journal of Statistical Theory and Practice, 13(2):??, June 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-018-
0035-2.

Khardani:2019:RKD

[645] Salah Khardani and Yousri Slaoui. Recursive kernel density estimation and optimal bandwidth selection under α: Mixing data. Journal of Statistical Theory and Practice, 13(2):??, June 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-018-
0031-6.

Thiuthad:2019:PEL

[646] Phontita Thiuthad and Nabendu Pal. Point estimation of the location parameter of a skew-normal distribution: Some fixed sample and asymptotic results. Journal of Statistical Theory and Practice, 13(2):??, June 2019. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-018-0033-4.

Alabdulhadi:2019:NPC

[647] Manal H. Alabdulhadi, Frank P. A. Coolen, and Tahani Coolen-Maturi. Nonparametric predictive comparison of two diagnostic tests based on total numbers of correctly diagnosed individuals. Journal of Statistical Theory and Practice, 13(3):??, September 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-019-0039-6;   http:
//link.springer.com/content/pdf/
10.1007/s42519-019-0039-6.pdf.

Maurya:2019:EPP

[648] Raj Kamal Maurya, Yogesh Mani Tripathi, and Manoj Kumar Rastogi. Estimation and prediction for a progressively first-failure censored inverted exponentiated Rayleigh distribution. Journal of Statistical Theory and Practice, 13(3):??, September 2019. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-019-0038-7.

Chaudhury:2019:VRU

[649] Prabir Chaudhury and Diganta Mukherjee. Are values reported using quantities and prices in consumption expenditure data? Journal of Statistical Theory and Practice, 13(3):??, September 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-019-
0040-0.

Barale:2019:NCC

[650] M. S. Barale and D. T. Shirke. Nonparametric control charts based on data depth for location parameter. Journal of Statistical Theory and Practice, 13(3):??, September 2019. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-019-0041-z.

Wang:2019:DFE

[651] Qing Wang and Lucia Tabacu. On distribution function estimation using log–odds interpolation. Journal of Statistical Theory and Practice, 13(3):??, September 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-019-
0044-9.

Ahmad:2019:TZC

[652] M. Rauf Ahmad. Tests of zero correlation using modified RV coefficient for high-dimensional vectors. Journal of Statistical Theory and Practice, 13(3):??, September 2019. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-019-0043-x;   http:
//link.springer.com/content/pdf/
10.1007/s42519-019-0043-x.pdf.

Ghatak:2019:CMN

[653] Debolina Ghatak and Bimal K. Roy. Conditional masking to numerical data. Journal of Statistical Theory and Practice, 13(3):??, September 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-019-0042-y.

Christianson:2019:APS

[654] Annette Christianson and Siva Sivaganesan. An approximate posterior simulation for GLMM with large samples. Journal of Statistical Theory and Practice, 13(3):??, September 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s42519-019-0045-8.

Wang:2019:DCI

[655] HaiYing Wang. Divide-and-conquer information-based optimal subdata selection algorithm. Journal of Statistical Theory and Practice, 13(3):??, September 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s42519-019-
                                        0048-5.

Najarzadegan:2019:DBD

[656] H. Najarzadegan, M. H. Alamatsaz, and I. Kazemi. Discrete bivariate distributions generated by copulas: DBEEW distribution. Journal of Statistical Theory and Practice, 13(3):??, September 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s42519-019-
                                        0046-7.

Shahbaz:2019:LWP

[657] Muhammad Qaiser Shahbaz. Life and work of Prof. Muhammad Hanif. Journal of Statistical Theory and Practice, 13(3):??, September 2019. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s42519-019-0047-6.

Shen:2019:RPI

[658] Changyu Shen and Huiping Xu. Randomized phase III oncology trials: a survey and empirical Bayes inference.

Journal of Statistical Theory and Practice, 13(3):??, September 2019. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-019-0049-4.

Trong:2019:HDC

[659] Dang Duc Trong, Ton That Quang Nguyen, and Nguyen Nhu Lan. Heteroscedastic deconvolution of P(X < Y) with compactly supported error densities. Journal of Statistical Theory and Practice, 13(3):??, September 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-019-
0050-y.

Nyarko:2019:ODS

[660] Eric Nyarko and Rainer Schwabe. Optimal designs for second-order interactions in paired comparison experiments with binary attributes. Journal of Statistical Theory and Practice, 13(4):??, December 2019. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-019-0051-x.

Kayal:2019:ICD

[661] Tanmay Kayal, Yogesh Mani Tripathi, and Liang Wang. Inference for the Chen distribution under progressive first-failure censoring. Journal of Statistical Theory and Practice, 13(4):??, December 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-019-
0052-9.

Bellahnid:2019:SKD

[662] Abdellatif Bellahnid and Amadou Sarr. Skewed Kotz distribution with application to financial stock returns. Journal of Statistical Theory and Practice, 13(4):??, December 2019. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s42519-019-0054-7.

Bobotas:2019:ODS

[663] Panayiotis Bobotas and Maria Kateri. Optimal designs for step–stress models under interval censoring. Journal of Statistical Theory and Practice, 13(4):??, December 2019. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s42519-019-0055-6.

Labopin-Richard:2019:CQS

[664] Tatiana Labopin-Richard, Fabrice Gamboa, Aurélien Garivier, and Jérôme Stenger. Conditional quantile sequential estimation for stochastic codes. Journal of Statistical Theory and Practice, 13(4):??, December 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s42519-019-0053-8.

Nystrom:2019:ICI

[665] Emily Nystrom, Julia L. Sharp, and William C. Bridges. The impact of correlated and/or interacting predictor omission on estimated regression coefficients in linear regression. Journal of Statistical Theory and Practice, 13(4):??, December 2019. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-019-0056-5.

Aguilar:2019:ZTP

[666] Guilherme A. S. Aguilar, Fernando A. Moala, and Gauss M. Cordeiro. Zero-truncated Poisson exponentiated gamma distribution: Application and estimation methods. Journal of Statistical Theory and Practice, 13(4):??, December 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-019-
0059-2.

Mostafa:2019:FPM

[667] Sayed A. Mostafa and Qingsong Shan. Finite population model-assisted estimation using combined parametric and nonparametric regression smoothers. Journal of Statistical Theory and Practice, 13(4):??, December 2019. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-019-0060-9.

Bagchi:2019:CCT

[668] Sunanda Bagchi and John P. Morgan. A complete class of type 1 optimal block designs with unequal replications. Journal of Statistical Theory and Practice, 13(4):??, December 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-019-0057-4.

Spencer:2019:ICF

[669] Neil A. Spencer, Pritam Ranjan, and Franklin Mendivil. Isomorphism check for 2^{n} factorial designs with randomization restrictions. Journal of Statistical Theory and Practice, 13(4):??, December 2019. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-019-
0064-5.

Pokhilko:2019:ODN

[670] Victoria Pokhilko, Qiong Zhang, Lulu Kang, and D’arcy P. Mays. D-optimal design for network A/B testing. Journal of Statistical Theory and Practice, 13(4):??, December 2019. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-019-0058-3.

Arnab:2019:ARE

[671] Raghunath Arnab, J. O. Olaomi, and B. C. Rodŕiguez. An addendum to “Regression Estimator in Ranked Set Sampling by Yu and Lam”. Journal of Statistical Theory and Practice, 13(4):??, December 2019. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-019-0065-4.

Sabbaghi:2019:EEC

[672] Arman Sabbaghi. An evaluation of estimation capacity under the conditional main effect parameterization. Journal of Statistical Theory and Practice, 13(4):??, December 2019. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-019-0061-8.

Kang:2019:BOD

[673] Lulu Kang and Xiao Huang. Bayesian A-optimal design of experiment with quantitative and qualitative responses. Journal of Statistical Theory and Practice, 13(4):??, December 2019. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-019-0063-6.

El-Sagheer:2019:BEU

[674] Rashad M. El-Sagheer, Mohamed A. W. Mahmoud, and Hasaballah M. Hasaballah. Bayesian estimations using MCMC approach under threeparameter Burr–XII distribution based on unified hybrid censored scheme. Journal of Statistical Theory and Practice, 13(4):??, December 2019. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-019-0066-3.

Zhu:2020:IEO

[675] Yuanyuan Zhu, Philip L. H. Yu, and Thomas Mathew. Improved estimation of optimal portfolio with an application to the US stock market. Journal of Statistical Theory and Practice, 14(1): ??, March 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-019-
0067-2.

Tayyab:2020:CCU

[676] Muhammad Tayyab, Muhammad Noor ul Amin, and Muhammad Hanif. A control chart using quartile pair ranked set sampling for monitoring the process mean. Journal of Statistical Theory and Practice, 14(1):??, March 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s42519-019-0069-0.

Shen:2020:DAF

[677] Sumin Shen, Zhiyang Zhang, and Xinwei Deng. On design and analysis of funnel testing experiments in webpage optimization. Journal of Statistical Theory and Practice, 14(1):??, March 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s42519-019-
                                        0068-1.

Dey:2020:DHC

[678] Asim Kumer Dey, Audrene Edwards, and Kumer Pial Das. Determinants of high crude oil price: a nonstationary extreme value approach. Journal of Statistical Theory and Practice, 14(1): ??, March 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s42519-019-
                                        0070-7.

Lin:2020:PCD

[679] Luyao Lin, Derek Bingham, and Ryan Lekivetz. Power considerations in designed experiments. Journal of Statistical Theory and Practice, 14(1):??, March 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s42519-019-
                                        0071-6.

Akhtar:2020:CEM

[680] Yasmeen Akhtar and Frederick Kin Hing Phoa. Cost-efficient mixed-level covering designs for testing experiments. Journal of Statistical Theory and Practice, 14(1):??, March 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-019-0062-7.

Lekivetz:2020:CAU

[681] Ryan Lekivetz and Joseph Morgan. Covering arrays: Using prior information for construction, evaluation and to facilitate fault localization. Journal of Statistical Theory and Practice, 14(1): ??, March 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-019-
0075-2.

Reangsephet:2020:AES

[682] O. Reangsephet, S. Lisawadi, and S. E. Ahmed. Adaptive estimation strategies in gamma regression model. Journal of Statistical Theory and Practice, 14(1): ??, March 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-019-
0076-1.

Yang:2020:GFR

[683] F. Yang, C. Devon Lin, and P. Ranjan. Global fitting of the response surface via estimating multiple contours of a simulator. Journal of Statistical Theory and Practice, 14(1):??, March 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-019-0077-0.

Sverdlov:2020:ODC

[684] Oleksandr Sverdlov, Yevgen Ryeznik, and Weng Kee Wong. On optimal designs for clinical trials: an updated review. Journal of Statistical Theory and Practice, 14(1):??, March 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s42519-019-0073-4.

Ciuperca:2020:CDQ

[685] Gabriela Ciuperca and Matús Maciak. Changepoint detection by the quantile LASSO method. Journal of Statistical Theory and Practice, 14(1):??, March 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s42519-019-
                                        0078-z.

Samanta:2020:EIS

[686] Debashis Samanta, Arnab Koley, Arindam Gupta, and Debasis Kundu. Exact inference of a simple step– stress model with hybrid Type–II stress changing time. Journal of Statistical Theory and Practice, 14(1):??, March 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s42519-019-
                                        0072-5.

Bagchi:2020:NRV

[687] Sunanda Bagchi. Nearly resolution V plans on blocks of small size. Journal of Statistical Theory and Practice, 14(1): ??, March 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s42519-019-
0074-3.

Singh:2020:IRR

[688] G. N. Singh and M. Usman. Improved regression–cum-ratio estimators using information on two auxiliary variables dealing with subsampling technique of non-response. Journal of Statistical Theory and Practice, 14(1):??, March 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-019-
0082-3.

Nicolussi:2020:GCL

[689] Federica Nicolussi and Maria Grazia Zoia. Gram–Charlier-like expansions of the convoluted hyperbolic–secant density. Journal of Statistical Theory and Practice, 14(1):??, March 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-019-0081-4.

Chowdhury:2020:CFE

[690] Shrabanti Chowdhury, Joshua Lukemire, and Abhyuday Mandal. A-ComVar: a flexible extension of common variance designs. Journal of Statistical Theory and Practice, 14(1):??, March 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-019-0079-y.

Samawi:2020:FIP

[691] Hani M. Samawi, Xinyan Zhang, and Haresh Rochani. Further improving the performance of logistic regression analysis using double extreme ranking. Journal of Statistical Theory and Practice, 14(1):??, March 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-0084-1.

Tarima:2020:ECU

[692] Sergey Tarima, Bonifride Tuyishimire, Rodney Sparapani, Lisa Rein, and John Meurer. Estimation combining unbiased and possibly biased estimators. Journal of Statistical Theory and Practice, 14(2):??, June 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-0083-2.

Wang:2020:LOD

[693] Zhongshen Wang and John Stufken. Locally D-optimal designs for binary responses in the presence of factorial effects. Journal of Statistical Theory and Practice, 14(2):??, June 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-0085-0.

Fushiki:2020:NBA

[694] Tadayoshi Fushiki and Tadahiko Maeda. Nonresponse bias adjustment in regression analysis. Journal of Statistical Theory and Practice, 14(2):??, June 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s42519-020-
                                        0086-z;   http://link.springer.
                                        com/content/pdf/10.1007/s42519-
                                        020-0086-z.pdf.

Portnoy:2020:TEP

[695] Stephen Portnoy. The two-envelope problem for general distributions. Journal of Statistical Theory and Practice, 14(2):??, June 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-020-0087-y.

Halvarsson:2020:MLE

[696] Daniel Halvarsson. Maximum likelihood estimation of asymmetric double Type II Pareto distributions. Journal of Statistical Theory and Practice, 14 (2):??, June 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-019-
0080-5.

Jankar:2020:OCD

[697] Jeevan Jankar, Abhyuday Mandal, and Jie Yang. Optimal crossover designs for generalized linear models. Journal of Statistical Theory and Practice, 14(2): ??, June 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-020-
00089-5.

Prus:2020:ODH

[698] Maryna Prus, Norbert Benda, and Rainer Schwabe. Optimal design in hierarchical random effect models for individual prediction with application in precision medicine. Journal of Statistical Theory and Practice, 14(2):??, June 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s42519-020-

00090-y. See correction [813].

Waggoner:2020:SMP

[699] Philip D. Waggoner. A simple method for purging mediation effects. Journal of Statistical Theory and Practice, 14 (2):??, June 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s42519-020-
                                        00095-7.

Coolen:2020:NPIa

[700] Frank P. A. Coolen and Sulafah Bin Himd. Nonparametric predictive inference bootstrap with application to reproducibility of the two-sample Kolmogorov–Smirnov test. Journal of Statistical Theory and Practice, 14(2): ??, June 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s42519-020-
                                        00097-5;   http://link.springer.
                                        com/content/pdf/10.1007/s42519-
                                        020-00097-5.pdf.

Lennartz:2020:MVC

[701] Jens Marcel Lennartz, Stefan Bedbur, and Udo Kamps. Minimum volume confidence regions for parameters of exponential distributions from different samples. Journal of Statistical Theory and Practice, 14(2):??, June 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00096-6.

Thomas:2020:NBD

[702] P. Yageen Thomas and Jitto Jose. A new bivariate distribution with Rayleigh and Lindley distributions as marginals. Journal of Statistical Theory and Practice, 14(2):??, June 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-020-00093-9.

Ghahramani:2020:TSR

[703] M. Ghahramani and S. S. White. Time series regression for zero-inflated and overdispersed count data: a functional response model approach. Journal of Statistical Theory and Practice, 14(2): ??, June 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-020-
00094-8.

Sousa:2020:PPR

[704] J. Beleza Sousa, Manuel L. Esquível, and Raquel M. Gaspar. Pulled-to-par returns for zero-coupon bonds historical simulation value at risk. Journal of Statistical Theory and Practice, 14(2): ??, June 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-020-
00092-w.

Aryuyuen:2020:TIT

[705] Sirinapa Aryuyuen and Winai Bodhisuwan. The Type II Topp Leonepower Lomax distribution with analysis in lifetime data. Journal of Statistical Theory and Practice, 14(2):??, June 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-020-
00091-x.

Stufken:2020:SSG

[706] John Stufken. Standing on the shoulders of a giant: The life and work of Samad Hedayat. Journal of Statistical Theory and Practice, 14(2):??, June 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s42519-020-
                                        00101-y.

Tomer:2020:RIM

[707] Sanjeev K. Tomer and M. S. Panwar. A review on inverse Maxwell distribution with its statistical properties and applications. Journal of Statistical Theory and Practice, 14(3):??, September 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s42519-020-
                                        00100-z.

Lee:2020:SMI

[708] Bu Hyoung Lee and Junyong Park. A spectral measure for the information loss of temporal aggregation. Journal of Statistical Theory and Practice, 14(3):??, September 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00099-3.

Grover:2020:IRT

[709] Lovleen Kumar Grover and Amanpreet Kaur. An improved regression type estimator of population mean with two auxiliary variables in stratified double sampling. Journal of Statistical Theory and Practice, 14(3):??, September 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s42519-020-
00107-6.

Michailidis:2020:MEP

[710] Georgios E. Michailidis, Ioannis I. Spyroglou, Dimitrios Zaridis, and Alexandros G. Rigas. A method of estimating the partial power spectrum of a bivariate point process and an application to a neurophysiological data set. Journal of Statistical Theory and Practice, 14(3):??, September 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-020-00105-8.

Bartok:2020:HAR

[711] Larissa Bartok and Matthias A. Burzler. How to assess rater rankings? a theoretical and a simulation approach using the sum of the pairwise absolute row differences (PARDs). Journal of Statistical Theory and Practice, 14(3):??, September 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-020-00103-w;   http:
//link.springer.com/content/pdf/
10.1007/s42519-020-00103-w.pdf.

Chen:2020:EIH

[712] Kedong Chen, William Li, and Sijian Wang. An easy-to-implement hierarchical standardization for variable selection under strong heredity constraint. Journal of Statistical Theory and Practice, 14(3):??, September 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00102-x.

Nik:2020:PMF

[713] A. Saadati Nik, A. Asgharzadeh, and M. Z. Raqab. Prediction methods for future failure times based on Type-II right-censored samples from new Pareto-type distribution. Journal of Statistical Theory and Practice, 14(3):??, September 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00106-7.

Saulo:2020:ARM

[714] Helton Saulo, Roberto Vila, Filidor Vilca, and José L. Martínez. On asymmetric regression models with allowance for temporal dependence. Journal of Statistical Theory and Practice, 14(3):??, September 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00104-9.

Schott:2020:ACC

[715] Dieter Schott. About the calculation of the coefficients of the Fleishman distributions. Journal of Statistical Theory and Practice, 14(3):??, September 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s42519-020-
                                        00111-w.

Ratnasingam:2020:CDS

[716] Suthakaran Ratnasingam and Wei Ning. Confidence distributions for skew normal change-point model based on modified information criterion. Journal of Statistical Theory and Practice, 14(3):??, September 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-020-00108-5.

Yuan:2020:SMR

[717] Linden Yuan, Lili Zhou, and Ao Yuan. Semiparametric mixture of regression models under unimodal error distribution. Journal of Statistical Theory and Practice, 14(3):??, September 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-020-00113-8.

Papamichail:2020:RAS

[718] Chrysanthi Papamichail, Salim Bouzebda, and Nikolaos Limnios. Regression analysis of stochastic fatigue crack growth model in a martingale difference framework. Journal of Statistical Theory and Practice, 14(3):??, September 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-020-00110-x.

DeYoreo:2020:BMM

[719] Maria DeYoreo and Jerome P. Reiter. Bayesian mixture modeling for multivariate conditional distributions. Journal of Statistical Theory and Practice, 14(3):??, September 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

//link.springer.com/article/10.
1007/s42519-020-00109-4.

Chesneau:2020:IML

[720] Christophe Chesneau, Lishamol Tomy, Jiju Gillariose, and Farrukh Jamal. The inverted modified Lindley distribution. Journal of Statistical Theory and Practice, 14(3):??, September 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00116-5.

Jha:2020:EHR

[721] Brijesh Kumar Jha, Ajaya Kumar Mahapatra, and Suchandan Kayal. Estimation of hazard rate of a selected exponential population. Journal of Statistical Theory and Practice, 14(3):??, September 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00112-9.

Nemeth:2020:RET

[722] László Németh and András Zempléni. Regression estimator for the tail index. Journal of Statistical Theory and Practice, 14(3):??, September 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00114-7.

Komaroff:2020:RBV

[723] Eugene Komaroff. Relationships between p-values and Pearson correlation coefficients, Type 1 errors and effect size errors, under a true null hypothesis. Journal of Statistical Theory and Practice, 14(3):??, September 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00115-6.

Ando:2020:DIM

[724] Shuji Ando. Directional index for measuring global symmetry in square tables. Journal of Statistical Theory and Practice, 14(3):??, September 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00117-4.

Uto:2020:BSL

[725] N. P. Uto and R. A. Bailey. Balanced semi-Latin rectangles: Properties, existence and constructions for block size two. Journal of Statistical Theory and Practice, 14(3):??, September 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00118-3.

Chaturvedi:2020:DPT

[726] Ajit Chaturvedi and Anshika Bhatnagar. Development of preliminary test estimators and preliminary test confidence intervals for measures of reliability of Kumaraswamy-G distributions based on progressive Type-II censoring. Journal of Statistical Theory and Practice, 14(3):??, September 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00119-2.

Dai:2020:LBT

[727] Xiaowu Dai and Yuhua Zhu. On large batch training and sharp minima: a Fokker–Planck perspective. Journal of Statistical Theory and Practice, 14(3):??, September 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00120-9.

Rasch:2020:EAT

[728] Dieter Rasch, Jürgen Pilz, and Rob Verdooren. An empirical approach that a two-stage procedure is better than Bechhofer’s approach. Journal of Statistical Theory and Practice, 14(4):??, December 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00098-4.

Reyes:2020:NDH

[729] Jimmy Reyes, Mario A. Rojas, and Héctor W. Gómez. Nakagami distribution with heavy tails and applications to mining engineering data. Journal of Statistical Theory and Practice, 14(4):??, December 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00122-7.

Ruzankin:2020:ACM

[730] Pavel S. Ruzankin. On absolute central moments of Poisson distribution. Journal of Statistical Theory and Practice, 14(4):??, December 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00121-8.

Rasch:2020:DMM

[731] Dieter Rasch and Rob Verdooren. Determination of minimum and maximum experimental size in one-, twoand three-way ANOVA with fixed and mixed models by R. Journal of Statistical Theory and Practice, 14(4):??, December 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00088-6.

Mahto:2020:ERM

[732] Amulya Kumar Mahto, Yogesh Mani Tripathi, and Fatih Kizilaslan. Estimation of reliability in a multicomponent stress-strength model for a general class of inverted exponentiated distributions under progressive censoring. Journal of Statistical Theory and Practice, 14(4):??, December 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00123-6.

Singh:2020:MDC

[733] Poonam Singh, Vandana Sarin, and Neha Midha. Mixture designs constructed using Taguchi’s mixed element orthogonal arrays. Journal of Statistical Theory and Practice, 14(4):??, December 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00126-3.

Evangelaras:2020:CGM

[734] Haralambos Evangelaras and Christos Peveretos. Construction of generalized minimum aberration threelevel orthogonal arrays with three, four and five columns. Journal of Statistical Theory and Practice, 14 (4):??, December 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00125-4.

Chai:2020:OFD

[735] Feng-Shun Chai and Ashish Das. A-optimal factorial designs for test versus control comparisons. Journal of Statistical Theory and Practice, 14(4):??, December 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00124-5.

Coolen:2020:NPIb

[736] Frank P. A. Coolen and Filipe J. Marques. Nonparametric predictive inference for test reproducibility by sampling future data orderings. Journal of Statistical Theory and Practice, 14(4):??, December 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00127-2.

Raju:2020:SRP

[737] David Chris Raju, S. M. Sunoj, and G. Rajesh. Some reliability properties of bivariate cumulative residual Tsallis entropy. Journal of Statistical Theory and Practice, 14(4):??, December 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00129-0.

Phuong:2020:DDN

[738] Cao Xuan Phuong and Le Thi Hong Thuy. Density deconvolution in a nonstandard case of heteroscedastic noises. Journal of Statistical Theory and Practice, 14(4):??, December 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00130-7.

Ghosh:2020:EPR

[739] Subir Ghosh. Exact permutation/ randomization tests algorithms. Journal of Statistical Theory and Practice, 14(4):??, December 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00131-6.

Elkadry:2020:HTW

[740] Alaa Elkadry and Gary C. McDonald. Hypothesis testing when data sources are uncertain. Journal of Statistical Theory and Practice, 14 (4):??, December 2020. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00132-5.

Das:2020:LWP

[741] Ashish Das. Life and work of Professor Aloke Dey: Our beloved Sir. Journal of Statistical Theory and Practice, 14(4):??, December 2020. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00136-1.

Ning:2021:CCP

[742] Wei Ning and Yanhong Wu. Common change-point estimation and changed panel isolation after sequential detection in an exponential family. Journal of Statistical Theory and Practice, 15(1):??, March 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00134-3.

Bera:2021:AVT

[743] Anil K. Bera, Osman Dogan, and Süleyman Taspinar. Asymptotic variance of test statistics in the ML and QML frameworks. Journal of Statistical Theory and Practice, 15(1):??, March 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00137-0.

Zhang:2021:MES

[744] Qi Zhang, Sadia Khalil, and Sat Gupta. Mean estimation of sensitive variables under non-response and measurement errors using optional RRT models. Journal of Statistical Theory and Practice, 15(1):??, March 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00135-2.

Sapam:2021:GLS

[745] Sobita Sapam and Bikas Kumar Sinha. Graeco Latin square designs with neighbor effects. Journal of Statistical Theory and Practice, 15(1):??, March 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00142-3.

Pal:2021:ICE

[746] Surya K. Pal and Housila P. Singh. Investigation of competent estimation procedure for estimating the finite population mean at current occasion addressing non-response in twooccasion successive sampling. Journal of Statistical Theory and Practice, 15(1):??, March 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00140-5.

Gardiner:2021:RMS

[747] Joseph C. Gardiner. Restricted mean survival time estimation: Nonparametric and regression methods. Journal of Statistical Theory and Practice, 15(1):??, March 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00144-1.

Rao:2021:NET

[748] B. L. S. Prakasa Rao. Nonparametric estimation of trend for stochastic differential equations driven by fractional Levy process. Journal of Statistical Theory and Practice, 15(1):??, March 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00138-z.

Zhai:2021:PRM

[749] Xiaoyu Zhai and Tapan K. Nayak. A post-randomization method for rigorous identification risk control in releasing microdata. Journal of Statistical Theory and Practice, 15(1):??, March 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00143-2.

Flores:2021:RVC

[750] Yuri Salazar Flores and Adán Díaz Hernández. Regular variation, conditions of domain of attraction and the existence of the tail dependence function in the general dependence case: a copula approach. Journal of Statistical Theory and Practice, 15(1):??, March 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00128-1.

Ong:2021:BCM

[751] Seng Huat Ong, Ramesh C. Gupta, and Shin Zhu Sim. Bivariate Conway– Maxwell Poisson distributions with given marginals and correlation. Journal of Statistical Theory and Practice, 15(1):??, March 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00141-4.

Al-Moisheer:2021:MTO

[752] A. S. Al-Moisheer, A. F. Daghestani, and K. S. Sultan. Mixture of two one-parameter Lindley distributions: Properties and estimation. Journal of Statistical Theory and Practice, 15(1):??, March 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00133-4.

Rao:2021:UES

[753] T. J. Rao. Unordering of estimators in sampling theory: Revisited. Journal of Statistical Theory and Practice, 15(1):??, March 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00139-y.

Deng:2021:PCM

[754] Yihao Deng and N. R. Chaganty. Paircopula models for analyzing family data. Journal of Statistical Theory and Practice, 15(1):??, March 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00146-z.

Lagergren:2021:DLA

[755] John Lagergren, Kevin Flores, and Semyon Tsynkov. Deep learning approach to the detection of scattering delay in radar images. Journal of Statistical Theory and Practice, 15(1):??, March 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00149-w.

Cengiz:2021:PAH

[756] Cigdem Cengiz, Dietrich von Rosen, and Martin Singull. Profile analysis in high dimensions. Journal of Statistical Theory and Practice, 15(1):??, March 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00154-z.

Bouali:2021:REC

[757] Dalal Lala Bouali, Fatah Benatia, and Christophe Chesneau. Robust estimator of conditional tail expectation of Pareto-type distribution. Journal of Statistical Theory and Practice, 15(1):??, March 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00153-0.

Li:2021:PTP

[758] Yuanhan Li, Jerome P. Keating, and Narayanaswamy Balakrishnan. PMC theorems on PCR-ridge class estimators. Journal of Statistical Theory and Practice, 15(1):??, March 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00150-3.

Shalabh:2021:GFN

[759] Shalabh and Subhra Sankar Dhar. Goodness of fit in nonparametric regression modelling. Journal of Statistical Theory and Practice, 15(1):??, March 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00148-x.

Singh:2021:EPM

[760] Housila P. Singh, Anurag Gupta, and Rajesh Tailor. Estimation of population mean using a difference-type exponential imputation method. Journal of Statistical Theory and Practice, 15(1):??, March 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00151-2.

Krebs:2021:NEI

[761] Johannes Krebs. A note on exponential inequalities in Hilbert spaces for spatial processes with applications to the functional kernel regression model. Journal of Statistical Theory and Practice, 15(1):??, March 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00147-y.

Pal:2021:CRM

[762] Ayan Pal, Shuvashree Mondal, and Debasis Kundu. A cure rate model for exponentially distributed lifetimes with competing risks. Journal of Statistical Theory and Practice, 15(1):??, March 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00145-0.

Bose:2021:KDE

[763] Arup Bose and Madhuchhanda Bhattacharjee. Kernel density estimates in a non-standard situation. Journal of Statistical Theory and Practice, 15(1):??, March 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00161-0.

Murray:2021:DFR

[764] Ryan W. Murray and Ernest Fokoué. Dropout fails to regularize nonparametric learners. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s42519-020-00158-
                                        9.

Mahzarnia:2021:EMP

[765] Ali Mahzarnia, Michael Grabchak, and Jiancheng Jiang. Estimation of the minimum probability of a multinomial distribution. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00163-y.

Kedem:2021:ETP

[766] Benjamin Kedem and Saumyadipta Pyne. Estimation of tail probabilities by repeated augmented reality. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00152-1.

Dutta:2021:OED

[767] Ganesh Dutta and Rita SahaRay. Optimal and efficient designs for comparing a set of test treatments with a set of controls in a heteroscedastic one-way layout with covariates. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00157-w.

Chaubey:2021:EEN

[768] Yogendra P. Chaubey and Nhat Linh Vu. On the estimation of entropy for non-negative data. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00165-4.

Ahlawat:2021:RMR

[769] Nitika Ahlawat and S. C. Malik. Regression modelling on reliability variation of a parallel series system of (m, ni) order with addition and removal of parallel paths. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00155-y.

Singh:2021:RIU

[770] Pushpinder Singh, Abhijit Mandal, and Ayanendranath Basu. Robust inference using the exponential-polynomial divergence. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00162-z.

Opheim:2021:STI

[771] Timothy Opheim and Anuradha Roy. Score tests for intercept and slope parameters of doubly multivariate linear models with skew-normal errors. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00159-8.

Ramanathan:2021:FIC

[772] T. V. Ramanathan and S. C. Pandhare. The focused information criterion for logistic time series regression models under locally biased estimating functions. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00169-0.

Ghosh:2021:DCN

[773] Kaushik Ghosh and Michael M. Nava. Detecting change in the number of modes for circular data. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00166-3.

Shi:2021:ECF

[774] Xiaoping Shi and Yuehua Wu. An empirical-characteristic-function-based change-point test for detection of multiple distributional changes. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00170-7.

Koul:2021:MDL

[775] Hira L. Koul and Indeewara Perera. A minimum distance lack-of-fit test in a Markovian multiplicative error model. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00168-1.

Ghosh:2021:ETB

[776] Malay Ghosh. Exponential tail bounds for chisquared random variables. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00156-x.

From:2021:SNB

[777] Steven G. From and Suthakaran Ratnasingam. Some new bounds for moment generating functions of various life distributions using mean residual life functions. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00176-1.

Kundu:2021:CLM

[778] Debasis Kundu and Rhythm Grover. On a chirp-like model and its parameter estimation using periodogram-type estimators. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00174-3.

Jammalamadaka:2021:FCC

[779] S. Rao Jammalamadaka, Brian Wainwright, and Qianyu Jin. Functional clustering on a circle using von Mises mixtures. Journal of Statistical Theory and Practice, 15(2):??, June 2021.

CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00173-4.

Szkutnik:2021:CAI

[780] Zbigniew Szkutnik. A comment on affine invariance and ancillarity in testing multivariate normality. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00180-5.

Sikaria:2021:BFM

[781] Shubhangi Sikaria, Rituparna Sen, and Neelesh S. Upadhye. Bayesian filtering for multi-period mean-variance portfolio selection. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00175-2.

Singh:2021:GCE

[782] G. N. Singh and M. Usman. Generalized class of estimators using dual of auxiliary variable under non-response. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00184-1.

Elshehawey:2021:GFH

[783] A. M. Elshehawey and Zhengming Qian. A gradual facilitate high-order multivariate Markov chains model with application to the changes of exchange rates in Egypt: New approach. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00179-y.

Sadeghkhani:2021:PST

[784] Abdolnasser Sadeghkhani and S. Ejaz Ahmed. Predicting the scoring time in hockey. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00186-z.

Pal:2021:BOR

[785] Ayan Pal, Sharmishtha Mitra, and Debasis Kundu. Bayesian orderrestricted inference of a Weibull multi-step step-stress model. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-020-00164-x.

Economou:2021:USB

[786] Polychronis Economou, Apostolos Batsidis, and Sonia Malefaki. Understanding the sampling bias: a case study on NBA drafts. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00167-2.

Iyengar:2021:MLE

[787] Satish Iyengar. Maximum likelihood estimation for an Ornstein– Uhlenbeck model for neural activity. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00190-3.

Babu:2021:STD

[788] G. Jogesh Babu, David Banks, and Shouyi Wang. A statistician teaches deep learning. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00193-0.

Bhushan:2021:ORT

[789] Shashi Bhushan and Abhay Pratap Pandey. Optimality of ratio-type imputation methods for estimation of population mean using higher order moment of an auxiliary variable. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00187-y.

Tanaka:2021:BMC

[790] Masahiro Tanaka. Bayesian matrix completion approach to causal inference with panel data. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00188-x.

Diawara:2021:EPU

[791] N. Diawara, B. E. Porter, and J. Kannane. Evaluation and prediction under hierarchical and bivariate copula models for seat belt use data. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00191-2.

Gao:2021:TIG

[792] Qingyi Gao and Xiao Wang. Theoretical investigation of generalization bounds for adversarial learning of deep neural networks. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00171-6.

Khalili:2021:EUG

[793] Atefeh Khalili, Farzad Eskandari, and Nader Nematollahi. Estimation of undirected graph with finite mixture of nonparanormal distribution. Journal of Statistical Theory and Practice, 15(2):??, June 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00192-1.

Sen:2021:CIB

[794] Ananda Sen and Ravindra Khattree. Characterization and inequalities based on the third moment. Journal of Statistical Theory and Practice, 15(3):??, September 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00181-4.

Schott:2021:MFP

[795] Dieter Schott. Monotone function problems and statistical selection procedures. Journal of Statistical Theory and Practice, 15 (3):??, September 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00177-0.

Bansal:2021:BDR

[796] Naveen K. Bansal, Mehdi Maadooliat, and Soroush Mahmoudiandehkordi. On Bayes decision rule for testing multiple hypotheses with nonsymmetric alternatives. Journal of Statistical Theory and Practice, 15 (3):??, September 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00182-3.

Mukhopadhyay:2021:PSP

[797] Nitis Mukhopadhyay. Purely sequential point estimation of a function of the mean in an exponential distribution. Journal of Statistical Theory and Practice, 15(3):??, September 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00183-2.

Ledbetter:2021:TPA

[798] Mark K. Ledbetter and Lucia Tabacu. Tests for patterned alternatives using logarithmic quantile estimation. Journal of Statistical Theory and Practice, 15(3):??, September 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00194-z.

SenGupta:2021:CT

[799] Ashis SenGupta and Moumita Roy. Clustering on the torus. Journal of Statistical Theory and Practice, 15(3):??, September 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00178-z.

Subrahmaniam:2021:ASQ

[800] Vignesh T. Subrahmaniam, Anup Dewanji, and Bimal K. Roy. Analysis of sequential quality improvement plans to obtain confidence bounds. Journal of Statistical Theory and Practice, 15(3):??, September 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00185-0.

Schumacher:2021:AIN

[801] Fernanda L. Schumacher, Dipak K. Dey, and Victor H. Lachos. Approximate inferences for nonlinear mixed effects models with scale mixtures of skew-normal distributions. Journal of Statistical Theory and Practice, 15(3):??, September 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00172-5.

Wywial:2021:ICP

[802] Janusz L. Wywiał and Grzegorz Sitek. The influence of clustering population on estimation accuracy of population totals vector. Journal of Statistical Theory and Practice, 15 (3):??, September 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00196-x.

Fujisawa:2021:EMC

[803] Kengo Fujisawa, Kouu Mitomi, and Kouji Tahata. Extension of marginal complementary log-log model and separations of marginal homogeneity for ordinal categorical data. Journal of Statistical Theory and Practice, 15(3):??, September 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00197-w.

Leao:2021:NBB

[804] Jeremias Leão, Marcelo Bourguignon, and Vinícius Calsavara. The negative binomial beta prime regression model with cure rate: Application with a melanoma dataset. Journal of Statistical Theory and Practice, 15(3):??, September 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00195-y.

Bhuiyan:2021:PCC

[805] Mohammad Alfrad Nobel Bhuiyan, Michael Wathen, and Marepalli Rao. Power comparisons in contingency tables. Journal of Statistical Theory and Practice, 15(3):??, September 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00199-8.

Feng:2021:IEI

[806] Xiaoshu Feng, Thomas Mathew, and Kofi Adragni. Interval estimation of the intra-class correlation in general linear mixed effects models. Journal of Statistical Theory and Practice, 15(3):??, September 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00202-2.

Heavlin:2021:EOA

[807] William D. Heavlin. On ensembles, Ioptimality, and active learning. Journal of Statistical Theory and Practice, 15(3):??, September 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00200-4.

Saulo:2021:FDL

[808] Helton Saulo, Roberto Vila, and Marcelo Bourguignon. On a family of discrete log-symmetric distributions. Journal of Statistical Theory and Practice, 15(3):??, September 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00204-0.

Jantre:2021:QRN

[809] S. R. Jantre, S. Bhattacharya, and T. Maiti. Quantile regression neural networks: a Bayesian approach. Journal of Statistical Theory and Practice, 15(3):??, September 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00189-w.

Hilow:2021:TCM

[810] Hisham Hilow and Islam Al-Maayteh. Three categories of minimum cost trend resistant 2^{n} factorial experiments derivable from the normalized Sylvester–Hadamard matrices of size 2^{n} × 2^{n}. Journal of Statistical Theory and Practice, 15 (3):??, September 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00198-9.

Leiva:2021:SSC

[811] Ricardo Leiva and Anuradha Roy. Self similar compound symmetry covariance structure. Journal of Statistical Theory and Practice, 15 (3):??, September 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00201-3.

Shanbhag:2021:GIC

[812] Damodar N. Shanbhag. Generalized integrated Cauchy functional equation with applications to probability models. Journal of Statistical Theory and Practice, 15(3):??, September 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00206-y.

Prus:2021:COD

[813] Maryna Prus, Norbert Benda, and Rainer Schwabe. Correction to: Optimal design in hierarchical random effect models for individual prediction with application in precision medicine. Journal of Statistical Theory and Practice, 15(3):??, September 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00205-z. See [698].

Liu:2021:DDA

[814] Xi Liu, Xueping Hu, and Keming Yu. A discrete density approach to Bayesian quantile and expectile regression with discrete responses. Journal of Statistical Theory and Practice, 15(3):??, September 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00203-1.

Khattree:2021:PSI

[815] Ravi Khattree, Sreenivasa Rao Jammalamadaka, and M. B. Rao. Preface to a special issue Celebrating the Birth Centenary of Professor C. R. Rao. Journal of Statistical Theory and Practice, 15(3):??, September 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00207-x.

Mathai:2021:FAR

[816] A. M. Mathai. Factor analysis revisited. Journal of Statistical Theory and Practice, 15(3):??, September 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-020-00160-1.

Lekivetz:2021:TSS

[817] Ryan Lekivetz and Joseph Morgan. On the testing of statistical software. Journal of Statistical Theory and Practice, 15(4):??, December 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00208-w.

Gomez:2021:ISF

[818] A. Nogales Gómez, L. Pronzato, and M.-J. Rendas. Incremental space-filling design based on coverings and spacings: Improving upon low discrepancy sequences. Journal of Statistical Theory and Practice, 15(4):??, December 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00210-2.

Bagchi:2021:AOP

[819] Sunanda Bagchi and Bhaskar Bagchi. Aspects of optimality of plans orthogonal through other factors and related multiway designs. Journal of Statistical Theory and Practice, 15(4):??, December 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00211-1.

Jokiel-Rokita:2021:MLE

[820] Alicja Jokiel-Rokita and Paweł Skoliński. Maximum likelihood estimation for an inhomogeneous gamma process with a log linear rate function. Journal of Statistical Theory and Practice, 15(4):??, December 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00212-0.

Zhang:2021:DBC

[821] Yumin Zhang and Arman Sabbaghi. The designed bootstrap for causal inference in big observational data. Journal of Statistical Theory and Practice, 15(4):??, December 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00213-z.

Wang:2021:CTD

[822] Yaping Wang, Fei Wang, and Qian Xiao. Connecting U-type designs before and after level permutations and expansions. Journal of Statistical Theory and Practice, 15 (4):??, December 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00209-9.

Shi:2021:MRS

[823] Chenlu Shi and Boxin Tang. Modelrobust subdata selection for big data. Journal of Statistical Theory and Practice, 15(4):??, December 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00217-9.

Li:2021:EDP

[824] Yiou Li and Xinwei Deng. On efficient design of pilot experiment for generalized linear models. Journal of Statistical Theory and Practice, 15(4):??, December 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00222-y.

Bailey:2021:OIB

[825] R. A. Bailey and Alia Sajjad. Optimal incomplete-block designs with low replication: a unified approach using graphs. Journal of Statistical Theory and Practice, 15(4):??, December 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00215-x. See cor-

rection [833].

Neumann:2021:MOC

[826] C. Neumann and J. Kunert. On MSE-optimal circular crossover designs. Journal of Statistical Theory and Practice, 15(4):??, December 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00220-0.

Beder:2021:CGS

[827] Jay H. Beder and Angela Dean. Cyclic generators for saturated orthogonal arrays. Journal of Statistical Theory and Practice, 15(4):??, December 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00214-y.

Wang:2021:OAB

[828] Zhongshen Wang and John Stufken. Orthogonal array based locally Doptimal designs for binary responses in the presence of factorial effects. Journal of Statistical Theory and Practice, 15(4):??, December 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00224-w.

Schorning:2021:ODC

[829] Kirsten Schorning and Holger Dette. Optimal designs for comparing regression curves: Dependence within and between groups. Journal of Statistical Theory and Practice, 15 (4):??, December 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00218-8.

Chai:2021:CRC

[830] Feng-Shun Chai, Rakhi Singh, and John Stufken. Connected row–column L-designs for symmetrical parallel line assays with two preparations. Journal of Statistical Theory and Practice, 15(4):??, December 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00219-7.

Zhang:2021:IBN

[831] Tianqi Zhang and Qiong Zhang. On the interface between nested designs and the multi-step interpolator. Journal of Statistical Theory and Practice, 15(4):??, December 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00216-w.

Rha:2021:BES

[832] Hyungmin Rha, Ming-Hung Kao, and Rong Pan. Bagging-enhanced sampling schedule for functional quadratic regression. Journal of Statistical Theory and Practice, 15(4):??, December 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00223-x.

Bailey:2021:COI

[833] R. A. Bailey and Alia Sajjad. Correction to: Optimal incomplete-block designs with low replication: a unified approach using graphs. Journal of Statistical Theory and Practice, 15(4):??, December 2021. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00225-9. See [825].

Idais:2021:EIO

[834] Osama Idais and Rainer Schwabe. Equivariance and invariance for optimal designs in generalized linear models exemplified by a class of gamma models. Journal of Statistical Theory and Practice, 15(4):??, December 2021. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-021-00221-z.

Zhang:2022:FSC

[835] Jin-Ting Zhang and Tianming Zhu. A further study on Chen–Qin’s test for two-sample Behrens–Fisher problems for high-dimensional data. Journal of Statistical Theory and Practice, 16(1):??, March 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00232-w.

Byukusenge:2022:IRG

[836] Béatrice Byukusenge, Dietrich von Rosen, and Martin Singull. On an important residual in the GMANOVA-MANOVA model. Journal of Statistical Theory and Practice, 16 (1):??, March 2022. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00235-7.

Ma:2022:RBC

[837] Zhuanzhuan Ma, Chanseok Park, and Min Wang. A robust bootstrap control chart for the log-logistic percentiles. Journal of Statistical Theory and Practice, 16(1):??, March 2022. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00239-3.

Tian:2022:TDG

[838] Weizhong Tian, Huihui Li, and Arjun K. Gupta. Tail dependence of generalized modified skew slash distribution. Journal of Statistical Theory and Practice, 16(1):??, March 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00242-8.

Guler:2022:PUS

[839] Nesrin Güler. Predictions under a system of linear regression models with correlated errors. Journal of Statistical Theory and Practice, 16(1):??, March 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00237-5.

Fontana:2022:CAB

[840] Roberto Fontana and Patrizia Semeraro. Computational and analytical bounds for multivariate Bernoulli distributions. Journal of Statistical Theory and Practice, 16(1):??, March 2022. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00231-x.

Beknazaryan:2022:FAD

[841] Aleksandr Beknazaryan. Function approximation by deep neural networks with parameters \{0, ±\frac{1}{2}, ±1, 2\}. Journal of Statistical Theory and Practice, 16(1):??, March 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00229-5.

Erlemann:2022:CGF

[842] Rasmus Erlemann and Bo Henry Lindqvist. Conditional goodness-of-fit tests for discrete distributions. Journal of Statistical Theory and Practice, 16(1):??, March 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00240-w. See cor-

rection [934].

Gillard:2022:SMV

[843] Jonathan Gillard, Emily O’Riordan, and Anatoly Zhigljavsky. Simplicial and minimal-variance distances in multivariate data analysis. Journal of Statistical Theory and Practice, 16(1):??, March 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00227-7.

Ohaegbulem:2022:RNR

[844] E. U. Ohaegbulem and P. E. Chigbu. Repairing non-rotatable central composite designs. Journal of Statistical Theory and Practice, 16(1):??, March 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00233-9.

Pawar:2022:DDB

[845] Somanath D. Pawar and Digambar T. Shirke. Data depth-based nonparametric tests for multivariate scales. Journal of Statistical Theory and Practice, 16(1):??, March 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00236-6.

Wang:2022:UWD

[846] H. Wang, D. Chen, and M. T. Hueman. Using weighted differences in hazards as effect sizes for survival data. Journal of Statistical Theory and Practice, 16(1):??, March 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00234-8.

Aldirawi:2022:MSD

[847] Hani Aldirawi and Jie Yang. Modeling sparse data using MLE with applications to microbiome data. Journal of Statistical Theory and Practice, 16(1):??, March 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00230-y.

Smallman:2022:LRS

[848] Luke Smallman and Andreas Artemiou. A literature review of (sparse) exponential family PCA. Journal of Statistical Theory and Practice, 16(1):??, March 2022. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00238-4.

Audu:2022:AUV

[849] Ahmed Audu, Rajesh Singh, and N. S. Dauran. Almost unbiased variance estimators under the simultaneous influence of non-response and measurement errors. Journal of Statistical Theory and Practice, 16(2):??, June 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00226-8.

Tyagi:2022:IEO

[850] Shikhar Tyagi, Arvind Pandey, and Christophe Chesneau. Identifying the effects of observed and unobserved risk factors using weighted Lindley shared regression model. Journal of Statistical Theory and Practice, 16(2):??, June 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-021-00241-9.

Benites:2022:FMB

[851] Luis Benites, Rocío Maehara, Filidor Vilca, and Fernando Marmolejo-Ramos. Finite mixture of Birnbaum– Saunders distributions using the kbumps algorithm. Journal of Statistical Theory and Practice, 16(2):??, June 2022. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic).

URL https://link.springer.com/
article/10.1007/s42519-022-00245-
z.

Vassiliou:2022:LLN

[852] P.-C. G Vassiliou. Laws of large numbers for non-homogeneous Markov systems with arbitrary transition probability matrices. Journal of Statistical Theory and Practice, 16(2):??, June 2022. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic).

URL https://link.springer.com/
article/10.1007/s42519-021-00228-
6.

Lemonte:2022:NAM

[853] Artur J. Lemonte. A note about model selection and hypothesis test procedure to discriminate Poisson and Bell models. Journal of Statistical Theory and Practice, 16(2):??, June 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00246-y.

Zhu:2022:MSN

[854] Xiaonan Zhu, Zheng Wei, and Tonghui Wang. Multivariate skew normalbased stochastic frontier models. Journal of Statistical Theory and Practice, 16(2):??, June 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00249-9.

Singh:2022:SGC

[855] G. N. Singh, M. Usman, and B. Khatoon. Some general classes of efficient estimators in case of missing data. Journal of Statistical Theory and Practice, 16(2):??, June 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00252-0.

Dey:2022:HDM

[856] Asim K. Dey and Kumer P. Das. How do mobility restrictions and social distancing during COVID-19 affect oil price? Journal of Statistical Theory and Practice, 16(2):??, June 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00247-x.

Khadher:2022:SAM

[857] Fatma Ben Khadher and Yousri Slaoui. The stochastic approximation method for recursive kernel estimation of the conditional extreme value index. Journal of Statistical Theory and Practice, 16(2):??, June 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00257-9.

Camper:2022:BRC

[858] Trevor R. Camper, Stephen W. Carden, and Russell C. Land. Bias in rank correlation under mixture models. Journal of Statistical Theory and Practice, 16(2):??, June 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00253-z.

Menezes:2022:CBR

[859] André F. B. Menezes, Marcelo Bourguignon, and Josmar Mazucheli. Complementary beta regression model for fitting bounded data. Journal of Statistical Theory and Practice, 16(2):??, June 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00256-w.

Junior:2022:PPE

[860] Paulo Cerqueira dos Santos Junior and Silvana Schneider. Power piecewise exponential model for interval-censored data. Journal of Statistical Theory and Practice, 16(2):??, June 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00254-y.

Arsham:2022:BRB

[861] Aryana Arsham, Ionut Bebu, and Thomas Mathew. A bivariate regression-based cost-effectiveness analysis. Journal of Statistical Theory and Practice, 16(2):??, June 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00255-x.

Tiwari:2022:URR

[862] Kuldeep Kumar Tiwari, Sandeep Bhougal, Sunil Kumar, and Khalid Ul Islam Rather. Using randomized response to estimate the population mean of a sensitive variable under the influence of measurement error. Journal of Statistical Theory and Practice, 16(2):??, June 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00251-1.

Yankam:2022:AOU

[863] Brenda Mbouamba Yankam and Abimibola Victoria Oladugba. Augmented orthogonal uniform composite designs for fitting third-order model. Journal of Statistical Theory and Practice, 16(2):??, June 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00260-0.

Papageorgiou:2022:BDP

[864] H. Papageorgiou and Maria Vardaki. Bivariate discrete Poisson–Lindley distributions. Journal of Statistical Theory and Practice, 16(2):??, June 2022. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00261-z.

EL-Sagheer:2022:CTH

[865] Rashad M. EL-Sagheer, Mohamed A. W. Mahmoud, and Waleed B. H. Etman. Characterizations and testing hypotheses for NBRUL-t_{0} class of life distributions. Journal of Statistical Theory and Practice, 16(2):??, June 2022. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00258-8.

Gatto:2022:BIB

[866] Riccardo Gatto and Sara Salvador. Bayesian inference on the bimodality of the generalized von Mises distribution. Journal of Statistical Theory and Practice, 16(2):??, June 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00250-2.

daSilvaBraga:2022:RER

[867] Altemir da Silva Braga, Gauss M. Cordeiro, Edwin M. M. Ortega, Giovana O. Silva, and Julio C. S. Vasconcelos. A random-effects regression model based on the odd loglogistic skew normal distribution. Journal of Statistical Theory and Practice, 16(2):??, June 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00248-w.

Qiu:2022:BSB

[868] Jin Qiu and Weining Zhu. Bandwidth selection in blocks empirical likelihood method for time series. Journal of Statistical Theory and Practice, 16(3):??, September 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00262-y.

Jayakumar:2022:SMR

[869] G. S. David Sam Jayakumar, W. Samuel, and Bejoy John Thomas. A squared Mahalanobis rank distances and detection of outliers in multivariate ordinal data. Journal of Statistical Theory and Practice, 16(3):??, September 2022.

CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00259-7.

Bansal:2022:OCD

[870] Namisha Bansal and Davinder Kumar Garg. Optimum covariate designs for three associate PBIB designs. Journal of Statistical Theory and Practice, 16(3):??, September 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00244-0.

Gomes:2022:NRF

[871] M. Ivette Gomes, Lígia Henriques-Rodrigues, and Dinis Pestana. Nonregular frameworks and the mean-of-order p extreme value index estimation. Journal of Statistical Theory and Practice, 16(3):??, September 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00264-w.

Ciuperca:2022:MCP

[872] Gabriela Ciuperca and Nicolas Dulac. Multiple change-points estimation in linear regression models via an adaptive LASSO expectile loss function. Journal of Statistical Theory and Practice, 16(3):??, September 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00265-9.

Christofides:2022:NMI

[873] Tasos C. Christofides and Charalambos Charalambous. New maximal inequalities for reverse demimartingales and reverse demisubmartingales. Journal of Statistical Theory and Practice, 16(3):??, September 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00263-x.

Singh:2022:EIM

[874] G. N. Singh and Ashok K. Jaiswal. Efficient imputation methods to handle missing data in sample surveys. Journal of Statistical Theory and Practice, 16(3):??, September 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00266-8.

Ndrin:2022:MHD

[875] Julien Apala N’drin and Ouagnina Hili. Minimum Hellinger distance estimation for discretely observed stochastic processes using recursive kernel density estimator. Journal of Statistical Theory and Practice, 16(3):??, September 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00269-5.

Dafnis:2022:BDO

[876] Spiros D. Dafnis, Markos V. Koutras, and Frosso S. Makri. Binomial distribution of order k in a modified binary sequence. Journal of Statistical Theory and Practice, 16 (3):??, September 2022. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00267-7.

Hahn:2022:TBB

[877] Eugene D. Hahn. The tilted beta-binomial distribution in overdispersed data: Maximum likelihood and Bayesian estimation. Journal of Statistical Theory and Practice, 16(3):??, September 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00271-x.

Acanal:2022:CML

[878] Nese Acanal and Tefaruk Haktanir. Computation of maximum-likelihood parameters of the generalized logistic distribution by three-step Newton– Raphson algorithm. Journal of Statistical Theory and Practice, 16 (3):??, September 2022. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00243-1.

Chowdhury:2022:SIM

[879] Faysal A. Chowdhury and K. Krishnamoorthy. Statistical intervals for Maxwell distributions. Journal of Statistical Theory and Practice, 16(3):??, September 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00270-y.

Singh:2022:EMS

[880] Neha Singh, Gajendra K. Vishwakarma, Neelesh Kumar, and Housila P. Singh. Estimation of mean of sensitive variable using multiplicative scramble variable under measurement error.

Journal of Statistical Theory and Practice, 16(3):??, September 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00272-w.

Papadatos:2022:CFD

[881] Nickos Papadatos. The characteristic function of the discrete Cauchy distribution in memory of T. Cacoullos. Journal of Statistical Theory and Practice, 16(3):??, September 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00268-6.

Pal:2022:SVE

[882] Suvra Pal, Sandip Barui, Katherine Davies, and Nutan Mishra. A stochastic version of the EM algorithm for mixture cure model with exponentiated Weibull family of lifetimes. Journal of Statistical Theory and Practice, 16(3):??, September 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00274-8.

Charalambides:2022:FDB

[883] Charalambos A. Charalambides. q-Fermi–Dirac and q-Bose–Einstein stochastic distributions (statistics). Journal of Statistical Theory and Practice, 16(3):??, September 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00277-5.

Mukhopadhyay:2022:PSM

[884] Nitis Mukhopadhyay and Jing Li. Purely sequential minimum risk point estimation (MRPE) for a survival function in an exponential distribution: Illustration with remission times for bladder cancer patients. Journal of Statistical Theory and Practice, 16(3):??, September 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00273-9.

Sapra:2022:ALT

[885] Pujita Sapra, Sadia Khalil, and Sat Gupta. Accounting for lack of trust in optional binary RRT models using a unified measure of privacy and efficiency. Journal of Statistical Theory and Practice, 16 (3):??, September 2022. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00279-3. See cor-

rection [891].

Goffard:2022:GFP

[886] P.-O. Goffard, S. Rao Jammalamadaka, and S. G. Meintanis. Goodness-of-fit procedures for compound distributions with an application to insurance. Journal of Statistical Theory and Practice, 16(3):??, September 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00276-6.

Bae:2022:TIG

[887] Taehan Bae and Andrei Volodin. Type-II generalized crack distribution with application to heavy-tailed data modeling. Journal of Statistical Theory and Practice, 16(3):??, September 2022. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00281-9.

El-Sherpieny:2022:BCD

[888] El-Sayed A. El-Sherpieny, Hiba Z. Muhammed, and Ehab M. Almetwally. Bivariate Chen distribution based on copula function: Properties and application of diabetic nephropathy. Journal of Statistical Theory and Practice, 16(3):??, September 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00275-7.

Keaton:2022:DDC

[889] Timothy J. Keaton and Arman Sabbaghi. Dismemberment and design for controlling the replication variance of regret for the multi-armed bandit. Journal of Statistical Theory and Practice, 16(4):??, December 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00280-w.

Meraou:2022:UBC

[890] M. A. Meraou, N. M. Al-Kandari, and M. Z. Raqab. Univariate and bivariate compound models based on random sum of variates with application to the insurance losses data. Journal of Statistical Theory and Practice, 16(4):??, December 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00282-8.

Sapra:2022:CAL

[891] Pujita Sapra, Sadia Khalil, and Sat Gupta. Correction to: Accounting for lack of trust in optional binary RRT models using a unified measure of privacy and efficiency. Journal of Statistical Theory and Practice, 16(4):??, December 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00285-5. See [885].

Bayramoglu:2022:RDA

[892] Ismihan Bayramoglu and Murat Ozkut. Recent developments about Marshall–Olkin bivariate distribution. Journal of Statistical Theory and Practice, 16(4):??, December 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00278-4.

Makhdum:2022:MES

[893] Mahnaz Makhdum, Aamir Sanaullah, and Muhammad Hanif. Mean estimation of a sensitive variable using optional randomized response technique in two-phase sampling when non-response and measurement error are simultaneously present. Journal of Statistical Theory and Practice, 16(4):??, December 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00283-7.

Balabdaoui:2022:ECF

[894] Fadoua Balabdaoui, Andrei Kolar, Yulia Kulagina, and Lilian Müller. Estimation of the complexity of a finite mixture distribution: From well- to less known methods. Journal of Statistical Theory and Practice, 16(4):??, December 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00289-1.

Tzoumerkas:2022:CPE

[895] G. Tzoumerkas, D. Fouskakis, and I. Ntzoufras. A comparison of power–expected-posterior priors in shrinkage regression. Journal of Statistical Theory and Practice, 16 (4):??, December 2022. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00284-6.

Nakagawa:2022:AED

[896] Tomoyuki Nakagawa and Shuntaro Ohtsuka. An asymptotic expansion for the distribution of Euclidean distance-based discriminant function in normal populations. Journal of Statistical Theory and Practice, 16 (4):??, December 2022. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00292-6.

Soyinka:2022:DCR

[897] A. T. Sóyínká and A. A. Olósundé. On discretization of continuous random variables for contingency tables:

Discrete Johnson systems of distribution as a case study with applications. Journal of Statistical Theory and Practice, 16(4):??, December 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00290-8.

Lefevre:2022:MLI

[898] Claude Lefèvre and Muhsin Tamturk. More for less insurance model: An alternative to (re)Insurance. Journal of Statistical Theory and Practice, 16(4):??, December 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00286-4.

Singh:2022:SOE

[899] Garib Nath Singh and Mahamood Usman. Some optimal estimators of finite population variance using dual of auxiliary variable in the presence of random non-response. Journal of Statistical Theory and Practice, 16(4):??, December 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00288-2.

Livingston:2022:NAB

[900] G. C. Livingston, Jr. and J. C. W. Rayner. Nonparametric analysis of balanced incomplete block rank data. Journal of Statistical Theory and Practice, 16(4):??, December 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00287-3.

Ghosh:2022:CCP

[901] Abhik Ghosh, María Jaenada, and Leandro Pardo. Classification of COVID19 patients using robust logistic regression. Journal of Statistical Theory and Practice, 16 (4):??, December 2022. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00295-3.

Ferreira:2022:BRC

[902] Jacqueline Alves Ferreira, Rosangela Helena Loschi, and Fabrizio Ruggeri. Bayesian robustness in change point analysis. Journal of Statistical Theory and Practice, 16 (4):??, December 2022. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00294-4.

El-Horbaty:2022:NCD

[903] Yahia S. El-Horbaty. A note on covariance decomposition in linear models with nested-error structure: new and alternative derivations of the F -test. Journal of Statistical Theory and Practice, 16(4):??, December 2022. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00291-7.

Marengo:2022:MPM

[904] James E. Marengo and David L. Farnsworth. The memoryless property and moments of the Gumbel distribution. Journal of Statistical Theory and Practice, 16(4):??, December 2022. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00296-2.

Azizi:2022:CBA

[905] Mohammad Javad Azizi, Sheldon M. Ross, and Zhengu Zhang. Choosing the best arm with guaranteed confidence. Journal of Statistical Theory and Practice, 16(4):??, December 2022. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00300-9.

Bansal:2023:SRE

[906] Namisha Bansal, Davinder Kumar Garg, and Ravindra Khattree. Some results on the existence or nonexistence of constant block-sum triangular and other related designs. Journal of Statistical Theory and Practice, 17(1):??, March 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00301-8.

Singh:2023:AEC

[907] Housila P. Singh, Priyanka Malviya, and Rajesh Tailor. An alternative efficient class of estimators for finite population mean using information on an auxiliary attribute in sample surveys. Journal of Statistical Theory and Practice, 17(1):??, March 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00299-z.

Zheng:2023:FCV

[908] Nan Zheng and Noel Cadigan. Frequentist conditional variance for nonlinear mixed–effects models. Journal of Statistical Theory and Practice, 17(1):??, March 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00304-5.

Cordeiro:2023:MDA

[909] Gauss M. Cordeiro, Abraão D. C. Nascimento, and Jodavid A. Ferreira. The MOGN distribution applied to medical imagery processing. Journal of Statistical Theory and Practice, 17(1):??, March 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00293-5.

Prajapati:2023:BSP

[910] Deepak Prajapati, Sharmishtha Mitra, and Debasis Kundu. Bayesian sampling plan for the exponential distribution with generalized Type-I hybrid censoring scheme. Journal of Statistical Theory and Practice, 17(1):??, March 2023. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00297-1.

Frost:2023:IPM

[911] Fay Frost, Rachel McCrea, Ruth King, Olivier Gimenez, and Elise Zipkin. Integrated population models: Achieving their potential. Journal of Statistical Theory and Practice, 17(1):??, March 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00302-7.

Ghanbari:2023:WEC

[912] Bahareh Ghanbari and Esmaeil Shirazi. Wavelet estimation of copula density via thresholding methods under censoring. Journal of Statistical Theory and Practice, 17(1):??, March 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00306-3.

Khattree:2023:CNC

[913] Ravindra Khattree. Constant and nearly constant block-sum partially balanced incomplete block designs and magic rectangles. Journal of Statistical Theory and Practice, 17(1):??, March 2023. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00307-2.

Lee:2023:BAF

[914] Dasom Lee and Sujit Ghosh. Bayesian analysis of first-order Markov models for autocorrelated binary responses. Journal of Statistical Theory and Practice, 17(1):??, March 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00305-4.

Abdallah:2023:QEM

[915] Mohamed S. Abdallah. Quantile estimation in modified ranked set sampling methods. Journal of Statistical Theory and Practice, 17(1):??, March 2023. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00298-0.

Guiatni:2023:MCT

[916] Ahlam Guiatni and Hisham Hilow. Minimum cost trend-free 2^{n−(n−k)} fractional factorial designs of resolution IV derivable from the normalized Sylvester–Hadamard matrices. Journal of Statistical Theory and Practice, 17(1):??, March 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00303-6.

Ratnasingam:2023:CPD

[917] Suthakaran Ratnasingam and Wei Ning. Change point detection in linear failure rate distribution under random censorship. Journal of Statistical Theory and Practice, 17(1):??, March 2023. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00309-0.

Sarkar:2023:FAC

[918] Pritam Sarkar, Uttam Bandyopadhyay, and Rahul Bhattacharya. On fixed accuracy confidence interval in multivariate normal distribution with order 1 autoregressive covariance structure. Journal of Statistical Theory and Practice, 17(1):??, March 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00310-7.

Zhou:2023:ICR

[919] Ying Zhou, Liang Wang, Yuhlong Lio, and Sanku Dey. Inference of competing risks model with partially observed failure causes based on minimum ranked set sampling. Journal of Statistical Theory and Practice, 17(1):??, March 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00311-6.

Pal:2023:CRB

[920] Ayan Pal, Debashis Samanta, and Debasis Kundu. Cure rate-based step-stress model. Journal of Statistical Theory and Practice, 17 (1):??, March 2023. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00313-4.

Balakrishnan:2023:PSI

[921] N. Balakrishnan, Ch. A. Charalambides, T. C. Christofides, M. V. Koutras, and S. Meintanis. Preface to a special issue in memory of Professor Theophilos Cacoullos. Journal of Statistical Theory and Practice, 17(1):??, March 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00314-3.

Modarresi:2023:SLD

[922] Navideh Modarresi, Saeid Rezakhah, and Mohammad Mohammadi. Semi-Lévy-driven CARMA process: Estimation and prediction. Journal of Statistical Theory and Practice, 17(1):??, March 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00317-0.

Morgan:2023:OSR

[923] J. P. Morgan and Sunanda Bagchi. Optimality of some row–column designs. Journal of Statistical Theory and Practice, 17(1):??, March 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00315-2. See cor-

rection [939].

Shakeel:2023:ASP

[924] Abeera Shakeel, Asifa Kamal, Getayeneh Antehunegn Tesema, and Maryam Siddiqa. Analysis of spatial patterns and associated factors of stillbirth in Pakistan, PDHS (2017–18): A spatial and multilevel analysis. Journal of Statistical Theory and Practice, 17(1):??, March 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00308-1.

Buckland:2023:WPA

[925] S. T. Buckland, D. L. Borchers, T. A. Marques, and R. M. Fewster. Wildlife population assessment: Changing priorities driven by technological advances. Journal of Statistical Theory and Practice, 17(2):??, June 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00319-6.

Ma:2023:REE

[926] Zhuanzhuan Ma, Min Wang, and Chanseok Park. Robust explicit estimation of the log-logistic distribution with applications. Journal of Statistical Theory and Practice, 17(2):??, June 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-023-00322-x.

Gross:2023:NCP

[927] Jürgen Groß and Annette Möller. A note on Cohen’s d from a partitioned linear regression model. Journal of Statistical Theory and Practice, 17(2):??, June 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-023-00323-w.

Ranjan:2023:SIP

[928] Pritam Ranjan, Joseph Resch, and Abhyuday Mandal. Solving an inverse problem for time-series-valued computer simulators via multiple contour estimation. Journal of Statistical Theory and Practice, 17(2):??, June 2023. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00312-5.

Sultana:2023:SIJ

[929] Farha Sultana, Çağatay Çetinkaya, and Debasis Kundu. Statistical inference for the jointly adaptive progressive Type-II censored Weibull distributions. Journal of Statistical Theory and Practice, 17(2):??, June 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-022-00318-z.

G:2023:NEE

[930] Rajesh G., Richu Rajesh, and S. M. Sunoj. Nonparametric entropy estimation of conditional distribution under length-biased right censored sample. Journal of Statistical Theory and Practice, 17(2):??, June 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00326-7.

Nguyen:2023:CLO

[931] Nam-Ky Nguyen, Tung-Dinh Pham, and Mai Phuong Vuong. A catalog of 2-level orthogonal minimally aliased designs with small runs. Journal of Statistical Theory and Practice, 17(2):??, June 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00321-y.

Kocyigit:2023:NSR

[932] Eda Gizem Koçyigit and Khalid Ul Islam Rather. The new sub-regression type estimator in ranked set sampling. Journal of Statistical Theory and Practice, 17(2):??, June 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00324-9.

Koley:2023:PAT

[933] Tamalika Koley, Farha Sultana, and Anup Dewanji. Parametric analysis of tampered random variable model for multiple step–stress life test. Journal of Statistical Theory and Practice, 17(2):??, June 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-022-00316-1.

Erlemann:2023:CCG

[934] Rasmus Erlemann and Bo Henry Lindqvist. Correction to: Conditional goodness-of-fit tests for discrete distributions. Journal of Statistical Theory and Practice, 17(2):??, June 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-023-00330-x. See [842].

Yanev:2023:CED

[935] George Yanev. On characterization of the exponential distribution via hypoexponential distributions. Journal of Statistical Theory and Practice, 17(2):??, June 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-023-00327-6.

Janicki:2023:SCS

[936] Ryan Janicki, Scott H. Holan, Kyle M. Irimata, James Livsey, and Andrew Raim. Spatial change of support models for differentially private decennial census counts of persons by detailed race and ethnicity. Journal of Statistical Theory and Practice, 17(2):??, June 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-023-00328-5.

Huang:2023:ANQ

[937] Mei Ling Huang, Yansan Han, and William Marshall. An algorithm of nonparametric quantile regression. Journal of Statistical Theory and Practice, 17(2):??, June 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00325-8.

Tsukada:2023:HTI

[938] Shin ichi Tsukada. Hypothesis testing for independence given a blocked compound symmetric covariance structure in a high–dimensional setting. Journal of Statistical Theory and Practice, 17(2):??, June 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00329-4.

Morgan:2023:COS

[939] J. P. Morgan and Sunanda Bagchi. Correction to: Optimality of some row– column designs. Journal of Statistical Theory and Practice, 17(2):??, June 2023. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic).

URL https://link.springer.com/
article/10.1007/s42519-023-00320-

z. See [923].

Bhushan:2023:NLT

[940] Shashi Bhushan, Anoop Kumar, and Shivam Shukla. Novel logarithmic type estimators in presence of measurement errors. Journal of Statistical Theory and Practice, 17 (3):??, September 2023. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00333-8.

Kumar:2023:CRT

[941] Pushkal Kumar and Manas Ranjan Tripathy. Classification rules for two exponential populations with common location and ordered scales under censored samples. Journal of Statistical Theory and Practice, 17 (3):??, September 2023. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00332-9.

Luhayb:2023:SBM

[942] Asamh Saleh M. Al Luhayb, Tahani Coolen-Maturi, and Frank P. A. Coolen. Smoothed bootstrap methods for bivariate data. Journal of Statistical Theory and Practice, 17 (3):??, September 2023. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00334-7.

vanDorp:2023:GTS

[943] Johan René van Dorp and Ekundayo Shittu. The generalized twosided beta distribution with applications in project risk analysis. Journal of Statistical Theory and Practice, 17(3):??, September 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00335-6.

Stoudt:2023:NIS

[944] Sara Stoudt, Perry de Valpine, and William Fithian. Nonparametric identifiability in species distribution and abundance models: Why it matters and how to diagnose a lack of it using simulation. Journal of Statistical Theory and Practice, 17 (3):??, September 2023. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00336-5.

Wywial:2023:MLE

[945] Janusz L. Wywiał. On the maximum likelihood estimation of population and domain means. Journal of Statistical Theory and Practice, 17(3):??, September 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00337-4.

Yoneyama:2023:TEE

[946] Shintaro Yoneyama and Mihoko Minami. Treatment effects estimation with missing not at random data without outcome modeling. Journal of Statistical Theory and Practice, 17(3):??, September 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00338-3.

Teng:2023:RMT

[947] Peiyuan Teng and Min Xu. Random matrix time series. Journal of Statistical Theory and Practice, 17 (3):??, September 2023. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00339-2.

Hu:2023:ROD

[948] Rui Hu, Ion Bica, and Zhichun Zhai. Robust optimal design when missing data happen at random. Journal of Statistical Theory and Practice, 17(3):??, September 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00340-9.

Yokochi:2023:RMH

[949] Clara Yokochi, Regina Bispo, Fernando Ricardo, and Ricardo Calado. Regularization methods for high-dimensional data as a tool for seafood traceability. Journal of Statistical Theory and Practice, 17(3):??, September 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00341-8.

Feldmann:2023:FMD

[950] Carlina C. Feldmann, Sina Mews, Angelica Coculla, Ralf Stanewsky, and Roland Langrock. Flexible modelling of diel and other periodic variation in hidden Markov models. Journal of Statistical Theory and Practice, 17(3):??, September 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00342-7.

Grayson:2023:SSC

[951] Matthew A. Grayson, Alain Kangabire, Can Aygen, and Kevin Considine. Stochastic simulation of continuous time random walks: Minimizing error in time-dependent rate coefficients for diffusion-limited reactions. Journal of Statistical Theory and Practice, 17(3):??, September 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00343-6.

Kvaalseth:2023:ABN

[952] Tarald O. Kvålseth. Association between nominal categorical variables: New measure formulation based on metric distances and value validity. Journal of Statistical Theory and Practice, 17(4):??, December 2023. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00344-5.

Momozaki:2024:IAE

[953] Tomotaka Momozaki, Koji Cho, Tomoyuki Nakagawa, and Sadao Tomizawa. Improving the accuracy of estimating indexes in contingency tables using Bayesian estimators. Journal of Statistical Theory and Practice, 18(1):??, March 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00353-4.

Sun:2024:TLP

[954] Mingwei Sun and Rong Zheng. A time-lagged penalized regression model and applications to economic modeling. Journal of Statistical Theory and Practice, 18(1):??, March 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00354-3.

Sakai:2024:STM

[955] Remi Sakai, Ayaka Yagi, and Takashi Seo. Simultaneous tests for mean vectors and covariance matrices with three-step monotone missing data. Journal of Statistical Theory and Practice, 18(1):??, March 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-023-00355-2.

Sapra:2024:HPE

[956] Pujita Sapra, Somya Mohanty, Sadia Khalil, and Sat Gupta. A hybrid Paillier encryption and randomized response technique model for surveying sensitive questions with binary response. Journal of Statistical Theory and Practice, 18(1):??, March 2024. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-023-00359-y.

Zhang:2024:VMC

[957] Yingjuan Zhang and Jochen Einbeck. A versatile model for clustered and highly correlated multivariate data. Journal of Statistical Theory and Practice, 18(1):??, March 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-023-00357-0.

Watanabe:2024:RPC

[958] Kodai Watanabe, Kanta Naito, and Inge Koch. Robustness of principal component analysis with Spearman’s rank matrix. Journal of Statistical Theory and Practice, 18(1):??, March 2024. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-023-00358-z.

Tsuchida:2024:CDA

[959] Jun Tsuchida and Hiroshi Yadohisa. Canonical dependency analysis using a bias-corrected χ^{2} statistics matrix. Journal of Statistical Theory and Practice, 18(1):??, March 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00360-5.

Zhan:2024:FMM

[960] Dongying Zhan and Derek S. Young. Finite mixtures of mean-parameterized Conway–Maxwell–Poisson regressions. Journal of Statistical Theory and Practice, 18(1):??, March 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00362-3.

Mukhopadhyay:2024:MSM

[961] Nitis Mukhopadhyay and Anhar S. Aloufi. Multi-stage minimum risk point estimation strategies for comparing the locations from two negative exponential models and second-order asymptotics: Illustrations with simulated data and bone marrow transplant data. Journal of Statistical Theory and Practice, 18(1):??, March 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00361-4.

Berenhaut:2024:GPL

[962] Kenneth S. Berenhaut, John D. Foley, and Liangdongsheng Lyu. Generalized partitioned local depth. Journal of Statistical Theory and Practice, 18(1):??, March 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00356-1.

Devroye:2024:RVG

[963] Luc Devroye and John P. Nolan. Random variate generation for the first hit of a ball for the symmetric stable process in R^{d}. Journal of Statistical Theory and Practice, 18(1):??, March 2024. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-023-00364-1.

Zou:2024:FBS

[964] Yixuan Zou and Derek S. Young. Fiducial-based statistical intervals for zero-inflated gamma data. Journal of Statistical Theory and Practice, 18(1):??, March 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-024-00365-8.

Alghamdi:2024:RST

[965] Fatimah M. Alghamdi, Frank P. A. Coolen, and Tahani Coolen-Maturi. Reproducibility of statistical tests based on randomised response data. Journal of Statistical Theory and Practice, 18(1):??, March 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-024-00366-7.

Moriyama:2024:RPB

[966] Taku Moriyama. Reduction of potential boundary bias in kernel cumulative distribution estimation in univariate and multivariate settings. Journal of Statistical Theory and Practice, 18(1):??, March 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00367-6.

Ghazani:2024:SES

[967] Zahra Shokooh Ghazani. Shrinkage estimators for shape parameter of Gompertz distribution. Journal of Statistical Theory and Practice, 18(1):??, March 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00369-4.

Luhayb:2024:SBM

[968] Asamh S. M. Al Luhayb, Tahani Coolen-Maturi, and Frank P. A. Coolen. Smoothed bootstrap methods for hypothesis testing. Journal of Statistical Theory and Practice, 18(1):??, March 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00370-x.

Darijani:2024:IFD

[969] Maryam Darijani, Hojatollah Zakerzadeh, and Ali Akbar Jafari. Introducing a family of distributions by using the class of normal mean-variance mixture. Journal of Statistical Theory and Practice, 18(1):??, March 2024. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00368-5.

Borhani:2024:DIO

[970] Sh. Borhani, F. Ghapani, and R. Jafaraghaie. Detecting influential observations using Liu estimator in linear mixed measurement error models. Journal of Statistical Theory and Practice, 18(2):??, June 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-023-00363-2.

Kulkarni:2024:SBT

[971] Leena Kulkarni, Sangita Kulathinal, and Isha Dewan. U-statistics based tests for marginal hazard rate orderings of two dependent variables. Journal of Statistical Theory and Practice, 18(2):??, June 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00372-9.

Susam:2024:CSF

[972] Selim Orhun Susam and Christophe Chesneau. On the construction of a semiparametric family of bivariate copulas using the opposite diagonal section of copulas. Journal of Statistical Theory and Practice, 18(2):??, June 2024. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00371-w.

He:2024:SNB

[973] Weiwei He, Janice Zgibor, and Jongphil Kim. A skew-normal Bayesian semiparametric latent trait linear mixed effect model. Journal of Statistical Theory and Practice, 18(2):??, June 2024. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00373-8.

Zhang:2024:FDT

[974] Mingjuan Zhang and Shibin Zhang. A frequency-domain test for multivariate white noise. Journal of Statistical Theory and Practice, 18(2):??, June 2024. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00374-7.

Karamikabir:2024:WSE

[975] Hamid Karamikabir. Wavelet shrinkage estimation for mean matrix of matrixvariate elliptically contoured distributions. Journal of Statistical Theory and Practice, 18(2):??, June 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00375-6.

Sui:2024:EBS

[976] Qun Sui and Sujit K. Ghosh. Entropybased subsampling methods for big data. Journal of Statistical Theory and Practice, 18(2):??, June 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00377-4.

Plante:2024:CCT

[977] Jean-François Plante and Marisa Radatz. On the capability of classification trees and random forests to estimate probabilities. Journal of Statistical Theory and Practice, 18(2):??, June 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00376-5.

Pandey:2024:EPM

[978] M. K. Pandey and G. N. Singh. Estimation of population mean using exponential ratio and product type estimators under ranked set sampling in presence of measurement error. Journal of Statistical Theory and Practice, 18(2):??, June 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00378-3.

Chatterjee:2024:DFG

[979] Rahul Chatterjee and Nabendu Pal. A distribution-free goodness of fit test for copula model: An application to Farlie–Gumbel–Morgernstern copula. Journal of Statistical Theory and Practice, 18(2):??, June 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00379-2.

Yoshida:2024:CTA

[980] Ruriko Yoshida and David Barnhill. Connecting tables by allowing negative cell counts. Journal of Statistical Theory and Practice, 18(3):??, September 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00381-8.

Yadav:2024:CBS

[981] Kaushal Kumar Yadav, Sukanta Dash, Baidya Nath Mandal, and Rajender Parsad. Construction of balanced semi-Latin rectangles in block size four: An algorithmic approach. Journal of Statistical Theory and Practice, 18(3):??, September 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00380-9.

Guesmia:2024:ECT

[982] Nour Elhouda Guesmia, Djamel Meraghni, and Louiza Soltane. Estimating the conditional tail expectation of randomly right-censored heavy-tailed data. Journal of Statistical Theory and Practice, 18(3):??, September 2024. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00383-6.

Cheng:2024:ADN

[983] Kaiming Cheng, Yueyang Shen, and Ivo D. Dinov. Applications of deep neural networks with fractal structure and attention blocks for 2D and 3D brain tumor segmentation. Journal of Statistical Theory and Practice, 18(3):??, September 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00384-5.

Murshed:2024:ICI

[984] Md. Monzur Murshed and K. Krishnamoorthy. Improved confidence intervals based on combined information in univariate calibration. Journal of Statistical Theory and Practice, 18(3):??, September 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00385-4.

Vardaki:2024:MDP

[985] Maria Vardaki and H. Papageorgiou. On multivariate discrete Poisson– Lindley distributions. Journal of Statistical Theory and Practice, 18 (3):??, September 2024. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-024-00389-0.

Yin:2024:FMH

[986] Shuang Yin, Dipak K. Dey, Emiliano A. Valdez, Guojun Gan, and Xiaomeng Li. Flexible modeling of hurdle Conway–Maxwell–Poisson distributions with application to mining injuries. Journal of Statistical Theory and Practice, 18(3): ??, September 2024. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-024-00388-1.

Bhattacharya:2024:CTE

[987] Rahul Bhattacharya, Uttam Bandyopadhyay, and Abhik Sinha. Comparing treatment effects for the AB: BA crossover design with continuous responses: An alternative nonparametric approach. Journal of Statistical Theory and Practice, 18 (3):??, September 2024. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-024-00386-3.

Shieh:2024:EAH

[988] Gwowen Shieh. The extended Anderson and Hauck tests and sample size procedures for equivalence assessment in simple linear regressions. Journal of Statistical Theory and Practice, 18(3):??, September 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00382-7.

Wang:2024:CPD

[989] Jun Wang and Wei Ning. Change point detection in length-biased Weibull distribution for random censored data based on modified information criterion. Journal of Statistical Theory and Practice, 18(3):??, September 2024. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00391-6.

Makhdoom:2024:ORS

[990] Waqas Makhdoom, Muhammad Kashif Ali Shah, Nighat Zahra, and Syed Ejaz Ahmed. Optimizing robust shape parameter: Improved methodologies for Birnbaum–Saunders distribution. Journal of Statistical Theory and Practice, 18(3):??, September 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00392-5.

Aljarrah:2024:FGS

[991] Mohammad A. Aljarrah, Carl Lee, and Felix Famoye. Family of generalized symmetric distributions: Properties and applications. Journal of Statistical Theory and Practice, 18 (3):??, September 2024. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-024-00387-2.

Meche:2024:EUV

[992] Bailey Meche, Nathan McEntire, Melissa Hall, and Scott Richter. Effect of unequal variance on statistical tests for mixed paired and two-sample designs. Journal of Statistical Theory and Practice, 18(3):??, September 2024. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-024-00394-3.

Saurabh:2024:SPR

[993] Shyam Saurabh and Kishore Sinha. A survey on partially resolvable solutions of regular group divisible designs. Journal of Statistical Theory and Practice, 18(3):??, September 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-024-00393-4.

Singh:2024:IUI

[994] Kundan Singh, Yogesh Mani Tripathi, Chandrakant Lodhi, and Liang Wang. Inference for unit inverse Weibull distribution under block progressive Type-II censoring. Journal of Statistical Theory and Practice, 18(3):??, September 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-024-00395-2.

Modak:2024:NCA

[995] Soumita Modak. A new clustering accuracy measure based on relative distances and its cross-validation using Dirichlet distribution. Journal of Statistical Theory and Practice, 18(3):??, September 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00390-7.

daSilva:2024:MLE

[996] Jackelya Araujo da Silva, Marcelo Angelo Cirillo, and Lourenço Manuel. Maximum L_{q}-likelihood estimation: A study of entropy behavior for the Pareto–exponential distribution with application. Journal of Statistical Theory and Practice, 18 (3):??, September 2024. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00396-1.

Dhameliya:2024:COD

[997] Vaibhav N. Dhameliya, Raj Kamal Maurya, and Ritwik Bhattacharya. Compound optimal design for generalized exponential distribution under progressively censored data. Journal of Statistical Theory and Practice, 18(3):??, September 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00397-0.

Joutard:2024:NCS

[998] Cyrille Joutard. Numerical comparison of some saddlepoint approximations near the endpoints of the support. Journal of Statistical Theory and Practice, 18(4):??, December 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-024-00398-z.

Shinoda:2024:GMC

[999] Satoru Shinoda, Ayumu Uehara, Ryuya Yahagi, Takuya Yoshimoto, and Kouji Tahata. Generalized marginal cumulative homogeneity model for multi-way contingency tables. Journal of Statistical Theory and Practice, 18(4):??, December 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-024-00401-7.

Miranda:2024:NC

[1000] Enrique Miranda and Marco Zaffalon. On negative conglomerability. Journal of Statistical Theory and Practice, 18(4):??, December 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-024-00400-8.

Ahmad:2024:SRL

[1001] Shakeel Ahmad, Abdul Majid, and Muhammad Aslam. On some robust Liu estimators for the linear regression model with outliers: Theory, simulation and application. Journal of Statistical Theory and Practice, 18(4):??, December 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-024-00404-4.

Backhausz:2024:EPE

[1002] Ágnes Backhausz, Edit Bognár, Villő Csiszár, Damján Tárkányi, and András Zempléni. Estimating the parameters of epidemic spread on twolayer random graphs: a classical and a neural network approach. Journal of Statistical Theory and Practice, 18(4):??, December 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00405-3.

Dalitz:2024:SBR

[1003] Christoph Dalitz, Juliane Arning, and Steffen Goebbels. A simple bias reduction for Chatterjee’s correlation. Journal of Statistical Theory and Practice, 18(4):??, December 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00399-y.

Kumar:2024:CSO

[1004] A. Anil Kumar, Baidya Nath Mandal, Rajender Parsad, Sukanta Dash, and Mukesh Kumar. On construction of sliced orthogonal Latin hypercube designs. Journal of Statistical Theory and Practice, 18(4):??, December 2024. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00402-6.

Liao:2024:SEA

[1005] Yi-Hua Liao, Frederick Kin Hing Phoa, and Dave C. Woods. The summary of effect aliasing structure for two-level supersaturated and factorial designs. Journal of Statistical Theory and Practice, 18(4):??, December 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00407-1.

Shen:2024:ETB

[1006] Pao sheng Shen. Equivalence tests before the end of follow-up under Box– Cox transformation model. Journal of Statistical Theory and Practice, 18(4):??, December 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-024-00406-2.

Evangelaras:2024:TLD

[1007] Haralambos Evangelaras and Victor Trapouzanlis. Three-level designs for definitive screening with minimum correlation among columns that correspond to pure quadratic effects. Journal of Statistical Theory and Practice, 18(4):??, December 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-024-00408-0.

Trang:2024:EUR

[1008] Bui Thuy Trang and Cao Xuan Phuong. An estimation of P(X < Y < Z) using repeated observations with unknown noise distribution. Journal of Statistical Theory and Practice, 18(4):??, December 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-024-00403-5.

Alghamdi:2024:REB

[1009] Fatimah M. Alghamdi, Frank P. A. Coolen, and Tahani Coolen-Maturi. Reproducibility of estimates based on randomised response methods. Journal of Statistical Theory and Practice, 18(4):??, December 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00409-z.

Maryaki:2024:CAI

[1010] Forouzan Jafari Maryaki and Mousa Golalizadeh. A comparative analysis of implementing adaptive lasso penalty in hierarchical data: Quantile versus mean regression. Journal of Statistical Theory and Practice, 18(4):??, December 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00411-5.

Egozcue:2024:OBP

[1011] Martín Egozcue and Luis Fuentes García. Optimal bounds and practical insights: Cantelli’s inequality revisited. Journal of Statistical Theory and Practice, 18(4):??, December 2024. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00410-6.

Smith:2024:GWG

[1012] Scott Smith and Brenda Diaz-Martinez. Gompertz–Weibull Gompertz Type I: a four-parameter generalization of the Weibull family. Journal of Statistical Theory and Practice, 18(4):??, December 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00413-3.

Kumar:2024:APM

[1013] Ajeet Kumar, Subhamoy Maitra, and Somjit Roy. Almost perfect mutually unbiased bases that are sparse. Journal of Statistical Theory and Practice, 18(4):??, December 2024. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00414-2.

Seifollahi:2025:SEZ

[1014] Solmaz Seifollahi, Hossein Bevrani, and Zakariya Yahya Algamal. Shrinkage estimators in zero-inflated Bell regression model with application. Journal of Statistical Theory and Practice, 19(1):??, March 2025. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00415-1.

Sankaran:2025:SRA

[1015] P. G. Sankaran and S. Hari. Semiparametric regression analysis of mixed recurrent event and panel count data with multiple causes of failure. Journal of Statistical Theory and Practice, 19(1):??, March 2025. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00417-z.

Ghosh:2025:REA

[1016] Suman Kalyan Ghosh and Sreenivasan Ravi. On random extrema with applications to renewal processes. Journal of Statistical Theory and Practice, 19(1):??, March 2025. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00418-y.

Ando:2025:PCO

[1017] Shuji Ando and Tatsuro Kikuchi. Parsimonious cumulative odds-symmetry model and its decomposition for square tables with ordered categories. Journal of Statistical Theory and Practice, 19(1):??, March 2025. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00419-x.

Chasiotis:2025:SOS

[1018] Vasilis Chasiotis and Dimitris Karlis.

On the selection of optimal subdata for big data regression based on leverage scores. Journal of Statistical Theory and Practice, 19(1):??, March 2025. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00420-4.

Ferreira:2025:MTS

[1019] José A. Ferreira. Methods of testing a ‘small’ or ‘moderate’ number of hypotheses simultaneously: an account focusing on the control of the probability of at least one incorrect rejection and of the false discovery rate. Journal of Statistical Theory and Practice, 19(1):??, March 2025. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00412-4. See cor-

rection [1021].

Borges:2025:CRR

[1020] Patrick Borges and Agatha Rodrigues.

Cure rate regression model for turning point analysis in unimodal hazard functions. Journal of Statistical Theory and Practice, 19(1):??, March 2025. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00421-3.

Ferreira:2025:CMT

[1021] José A. Ferreira. Correction to: Methods of testing a ‘small’ or ‘moderate’ number of hypotheses simultaneously: an account focusing on the control of the probability of at least one incorrect rejection and of the false discovery rate. Journal of Statistical Theory and Practice, 19(1):??, March 2025. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-025-00426-6.  See

[1019].

Kumar:2025:ADL

[1022] Rajeev Ranjan Kumar, S. Vishnu Shankar, Ronit Jaiswal, Mrinmoy Ray, Neeraj Budhlakoti, and K. N. Singh. Advances in deep learning for medical image analysis: a comprehensive investigation. Journal of Statistical Theory and Practice, 19(1):??, March 2025. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00422-2.

Matula:2025:ELL

[1023] Przemysław Matuła. Exact laws of large numbers for extremal statistics from the Pareto-type distributions. Journal of Statistical Theory and Practice, 19(1):??, March 2025. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-025-00428-4.

Alawady:2025:MER

[1024] M. A. Alawady, H. M. Barakat, T. S.

Taher, and I. A. Husseiny. Measures of extropy for k-record values and their concomitants based on Cambanis family. Journal of Statistical Theory and Practice, 19(1):??, March 2025. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00423-1.

Hu:2025:OSF

[1025] Xiaoru Hu, Ping Ma, Shuang Li, Wei Li, and Ming Yang. The optimization of space-filling and orthogonality for Latin hypercube design using a local search-based NSGA-II. Journal of Statistical Theory and Practice, 19(1):??, March 2025. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-025-00430-w.

Mathew:2025:GPS

[1026] Sunil Mathew, Jasbir, Sr., and Daniel B. Rowe. Generation of patient specific cardiac chamber models using generative neural networks under a Bayesian framework for electroanatomical mapping. Journal of Statistical Theory and Practice, 19(1):??, March 2025. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

//link.springer.com/article/10.
1007/s42519-024-00416-0.

Zarei:2025:FBH

[1027] R. Zarei, Shahram Yaghoobzadeh Shahrestani, and F. Fadaei. Fuzzy E-Bayesian and H-Bayesian estimation of scalar parameter in Gompertz distribution under asymmetric loss functions. Journal of Statistical Theory and Practice, 19(1):??, March 2025. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-025-00429-3.

Fathizadeh:2025:PSM

[1028] M. Fathizadeh, K. Khorshidian, and H. Mansoori. Prediction in semi-Markov repairable systems. Journal of Statistical Theory and Practice, 19(1):??, March 2025. CO-DEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s42519-025-00427-5.

Anonymous:2025:JNa

[1029] Anonymous. Journal navigation. Journal of Statistical Theory and Practice, 19(1):??, March 2025. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic).